A Complete Bibliography of Publications in Reviews in Fisheries Science & Aquaculture

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.18

Title word cross-reference

L_{max} [383, 424].

-Based [634].

19 [666, 610]. 1950s [452]. 1960s [441]. 1969 [183]. 1980s [3]. 1991/92 [88]. 1995/96 [88].

2005a [175]. 2019 [643]. 226-Year-Old [303]. 2nd [369, 692].

4th [651].

Abalone [394, 245, 392, 393, 396, 402, 346, 373, 397]. Aboriginal [262]. Absolute [526]. Abundance [409, 348, 535, 251]. Abundance-Case [409]. Acanthopagrus [403, 248]. Access [638, 559]. Acclimation [243]. Accountability [220]. Accountability-Based [220]. Achieving [695].

Acid [568, 554, 61]. Acids [635, 693, 145, 13]. Acipenseriform [200, 350, 408]. Acoustic [390]. Across [551, 469, 618]. Act [183]. Action [463]. Actions [377]. active [538]. Activities [292, 37]. Activity [104]. Adaptation [599]. Adapting [516]. Adaptive [466, 578, 627, 14]. Additional [165]. Additive [117]. Additives [590, 541]. Addressing [238, 540]. Administration [463, 199]. Adoption [190]. Adult [602]. advanced [36]. Advancements [688]. Advances [471, 477, 414, 448, 567, 85, 491, 196, 133]. Advancing [333]. Aegean [334]. aeration [21]. aerobic [73]. Aeromonas [173]. Affecting [283, 358, 196]. Africa [582, 510, 584, 533]. African [271, 638, 598]. after [388, 286, 300]. Against [477]. Agar [110]. Agar-Producing [110]. Age [425, 198, 440]. Age-Dependent [425]. Ageing [438]. Agency [89]. Agent [648]. Agent-Based [648]. Aggregation [421]. Agonist [199]. agricultural [30]. Aires [501]. al [175]. Alabama [116, 167]. Alaska [359]. albacares [545]. albus [308]. Algae [110, 7]. Algal [127]. Aliger [621]. Aligning [642]. Allele [247]. allergens [2]. Alligator [495, 495]. Alluvial [179]. Along [690, 362]. Alternative [430, 100, 397, 542, 51]. Alternatives [271]. amarus [295, 181, 182]. America [691, 341, 412]. American [690, 464, 22, 606, 674, 286, 493]. americanus [22, 59]. Americas [133]. amino [13]. Ammonia [375]. among [451]. Anaerobic [655]. Analyses [588, 94, 96, 97, 227, 550]. Analysis [622, 289, 620, 193, 58, 327, 383, 424, 447, 603, 673, 288, 606, 581, 325, 535, 564, 397, 591, 493, 518, 408, 104, 412, 490, 643, 436, 690, 276, 545, 40, 59]. Analyzing [654]. Anchovy [479]. Anesthetics [517]. Angler [156, 531]. Anglers [352, 536]. Anguilla [534]. Animal [590, 673, 698, 538]. Animals [628, 624, 61, 29]. Annotated [443, 667]. Antarctic [108]. antennatus [452]. Anthropogenic [610, 602]. Anti [235]. Anti-Predator [235]. Antioxidant [583]. Antiviral [661]. Aphanomyces [566]. Applicability [646]. Applicable [424]. Application [222, 464, 425, 463, 677, 595, 491, 332, 624, 30, 12]. Applications [422, 475, 699, 527, 604, 482, 408]. Applying [696, 632]. Approach [592, 171, 108, 596, 87, 321, 628, 415, 229, 395, 220, 114, 476, 217, 540]. Approaches [208, 480, 637, 365, 563, 223, 40, 369]. apus [303]. Aquaculture [582, 499, 586, 472, 504, 585, 355, 381, 172, 203, 668, 422, 271, 612, 691, 580, 471, 470, 623, 686, 477, 480, 603, 547, 576, 661, 443, 532, 486, 138, 551, 579, 680, 693, 673, 534, 462, 563, 638, 510, 410, 454, 100, 463, 423, 150, 209, 599, 139, 344, 384, 628, 677, 287, 588, 484, 514, 503, 372, 525, 495, 549, 670, 685, 625, 446, 665, 482, 467, 633, 562, 610, 496, 351, 687, 353, 132, 522, 285, 312, 429, 548, 672, 494, 103, 650, 196, 106, 634, 640, 557, 629, 314, 512, 72, 30, 70]. aquaculture [538, 29, 369, 692, 651]. Aquaculture-Based [384]. aquacultured [74]. Aquafeed [264, 548]. Aquafeeds [568]. Aquaponics [607, 681]. Aquatic [590, 230, 176, 662, 498, 266, 356, 127, 666, 632, 111, 186, 562, 538, 376, 672, 300, 624, 35, 61]. Aquifers [179]. Arc [166]. Archipelago [175, 170, 174]. Arctic [12]. Area [630]. Areas [411, 382]. Arenas [253]. argus [453]. Argyrosomus [229]. Arid [176, 177]. Aristeus [452]. Arowana [557]. Arrangements [280]. Art [695, 457, 588]. Artemia [79, 274]. Artificial [156, 532, 269, 100, 199, 527, 268, 683, 21]. Artisanal [407, 455]. Asia [177, 128, 258, 130, 572]. Asian [641, 308, 557]. asinina [402]. Aspects [189, 72, 13]. Assay [277, 296]. Assemblages [292, 450]. Assess [494]. Assessing [417, 245, 364, 395]. Assessment [171, 102, 409, 365, 339, 646, 367, 296, 337, 525, 328, 513, 308, 291, 387, 332, 307, 550]. Assessments [680, 51]. assisted [69]. Associated [451, 547, 223, 628, 286, 280, 27]. Associations [489, 564]. Assumptions [629]. Astacus [315, 279]. Asynchronous [596]. Atlantic [53, 615, 690, 170, 175, 129, 529, 556, 383, 424, 441, 431, 447, 526, 615, 676, 635, 283, 41, 382, 199, 326, 379, 160, 252, 201, 493, 518, 614, 618, 412, 15, 421, 349, 658, 55, 647, 690]. Attachment [142]. Attitudes [577]. Attracted [243]. aurata [345]. Australia [409, 394, 385, 346, 6, 335, 332]. Australian [403, 83, 509, 373]. Author [153]. Authors [165]. Available [383, 555]. Avoidance [539]. Awareness [124].

B2B [190]. BACI [225]. Bacillus [538]. Backcrossing [98]. Bacteria [482, 554]. Bacterial [471]. Bacteriocins [482]. Bacteriophage [471]. Bacterium [196]. Bahamas [524]. Bait [459]. Baja [266]. Balance [643]. Baltic [221]. Bank [266]. Barbel [354]. barbugede [303]. Barbus [354]. Barracuda [297]. Barrier [542]. Based [320, 171, 668, 648, 384, 231, 220, 634, 608, 446]. Basin [289, 521, 421, 185, 158, 65, 292, 169]. Basins [116, 341, 178]. Basis [210]. Bass [504, 352, 345, 660, 36]. Batch [596]. Batch-Spawning [596]. Bath [334]. batrachi [319]. batrachus [319]. Bay [195, 387, 228, 241, 87, 147, 614, 250, 217]. Be [662, 346]. Bed [292]. Before [521]. Behavior [93, 142, 402, 235, 665, 688]. Behind [330]. Beloved [660]. Beneficial [370, 608]. Benefit [352, 638, 397]. Benefits [568, 590, 202, 156, 667]. Benthic [578]. Benthos [417]. Bering [17]. Best [116]. Between [195, 539, 179, 478, 685, 674, 657, 315, 115, 564, 524]. Bibliography [443, 667]. Bibliometric [361]. Bighead [427]. Bio [563, 538]. bio-active [538]. Bio-Control [563]. Biochemistry [543]. Biocides [117]. Biocontrol [480]. Biodiversity [111, 562]. Bioeconomic [396]. Bioeconomics [378]. Biofilms [588]. Biogeochemical [356]. Biogeography [682, 613, 627]. Biological [586, 197, 411, 224, 403, 192, 227, 314]. biologist [52]. Biology [657, 470, 144, 443, 532, 213, 594, 469, 453, 340, 273, 689, 675, 514, 567, 379, 298, 207, 578, 591, 95, 621, 627, 557, 27, 26, 32]. Bioluminescent [317]. Biometric [383]. Biomineralization [589]. Bioremediation [588, 538]. Biota [168]. Biotechnology [369]. Biotelemetry [142]. Biotransformation [90]. bird [17]. Bivalve [222, 90, 294, 100, 34]. Bivalves [213, 100]. Black [81, 391, 265, 626, 479, 248]. Blacklip [394]. Bladder [570]. Blasting [89]. Bloater [278]. Blooms [127]. Blue [228, 499, 656, 612, 225, 551, 234, 241, 652, 664, 250, 242, 217]. Blue-Green [499]. Bluefin [553, 556, 383, 424, 441, 431, 447, 526, 256, 420, 199, 326, 555, 201, 55]. Bluefish [362]. Bluespotted [310]. Board [357, 406, 428, 442, 465, 488]. Body [278]. Bolivia [521]. Bones [641]. Book [692, 458, 487, 152, 651, 445]. Boosting [563]. brasiliensis [685]. Brazil [521, 486]. Brazilian [493]. Bream [504, 345, 248]. Breeding [677, 256, 329]. Bridging [688]. Brief [296]. Brine [79]. Bristol [87]. British [175, 170, 14]. Broad [690]. Broad-Scale [690]. Broodstock [256, 257]. Brooks [175, 174, 260]. Broughton [170, 175, 174]. Brünnich [303]. Buenos [501]. Buffet [540]. Buffet-Style [540]. Burbot [670]. Burning [273]. butcheri [403, 248]. Butterfly [94, 95, 96, 97]. Butyrate [568]. Butyric [568]. Bycatch [630, 683].

C. [311]. Côte [625]. caeruleostictus [310]. Cage [254, 387, 654, 640]. Cage-Conditioned [387]. Cages [243]. Calendar [520]. Calendar-Clock [520]. California [266, 122, 126, 322, 121, 99]. Callinectes [234, 217]. campechanus [275]. camtchaticus [492]. Can [370, 346]. Canada [170, 175, 676, 421, 658]. canadum [443]. Canalization [183]. Candidate [514]. Canned [497]. Cannibalism [91, 467]. Capacity [578, 250, 73]. Capsules [146]. Captive [386, 329, 199, 235]. Captive-Reared [199, 235]. Capture [281, 254, 672, 533]. Carangidae [485]. Carbohydrate [345, 151]. Caribbean [412]. Carnivorous [264, 91, 324]. Carp [427, 436, 300, 300]. carpio [436]. Carps [641]. Carrying [250]. Case [409, 509, 340, 632, 446, 455, 674, 236, 335, 544, 658, 627]. Case-study [446]. Cases [270]. Catch [197, 409, 645, 404, 536, 493, 558, 658]. Catch-and-Release [197, 558]. Catchability [307]. Catches [115, 426]. catching [28]. Catfish [116, 386, 134, 123, 167, 46]. Causal [436]. Caused [628, 270]. Causes [615]. Causing [566]. Caviar [601]. Cellular [661, 532]. Centrachidae [660]. Central [622, 310, 177, 297, 99]. Centrarchidae [660]. Centropomus [238]. Century [147, 286, 106]. Cephalopoda [335]. cephalus [43]. cerebralis [289]. Certification [172, 360, 559, 455]. Chain [635, 693]. Challenges [407, 574, 136, 460, 139, 344, 484, 597, 502, 554, 130, 132, 602, 133, 593, 540]. Change [499, 686, 283, 516, 177, 588, 571, 597, 598, 625, 618, 601]. Changes [322, 389, 360, 286]. Changing [313]. Channel [116, 134, 167, 46]. Chanos [43]. Characiformes [341]. Characteristics [403, 25]. Characterization [608]. Charles [52]. Charleston [31]. charr [12]. Checklists [311]. Chemical [520, 473]. Chemistry [620]. Chesapeake [614, 228, 234, 241, 250, 217]. Chicken [349]. Chile [202]. Chilean [455]. China [646, 254, 462, 483, 344, 581, 664, 593, 654, 640]. Chinese [255, 652, 649]. Chinook [158, 523, 161, 121]. Chionoecetes [492]. Chitala [514]. Chlorinated [117]. Choice [536]. Cholesterol [507]. Chondrichthyan [689]. chrysotaenia [297, 198]. Chum [405, 161]. Cichla [94, 95, 96, 97]. Ciliate [44]. Circadian [118]. Circular [612]. Clam [484]. Clams [219, 26]. Clarias [319]. Clark [413]. Class [282]. Climate [499, 686, 404, 313, 283, 599, 516, 177, 588, 571, 597, 598, 625, 618, 104, 644]. Clock [520]. cloning [74]. Closures [658]. Cnidae [146]. Cnidarian [146]. Co [494]. Co-Occurrence [494]. Coal [292]. Coast [625, 690, 362]. Coastal [222, 407, 215, 355, 381, 501, 411, 343, 115, 252, 23]. Coasts [335]. Coastwide [362]. Cobia [443]. Coconut [472]. Cod [489, 246, 252, 221, 316]. Coho [161]. Cohort [238]. Coldwater [273]. Collapse [441, 615, 247]. Collected [181]. Collection [669]. collectors [28]. Colorado [178]. Columbia [170, 175, 158, 65, 14, 674, 3]. Comet [277]. Comments [311]. Commercial [425, 594, 527, 525, 474, 104, 601]. commercialis [6]. Commercially [603, 501]. Common [172, 238, 365, 436, 300, 540]. Commonly [517]. Communities [486, 519, 413]. Community [94, 231, 508, 342, 17]. Community-Based [231]. Comparative [582, 356, 379, 104, 490, 4]. Comparison [146, 115, 493]. Comparisons [86, 248]. compensation [32]. Competitors [659, 62]. Complement [325]. complete [59]. Completion [521]. Complex [466, 539, 234]. Complexity [197, 688]. Compliance [679]. components [538]. Comprehensive [323]. Compromise [478]. Concentrations [497, 59]. Concept [430, 696, 588]. Concepts [271, 516]. Conceptual [149]. Concerns [123, 420]. Conch [657, 544, 524, 621]. Concise [105, 634]. Condition [391]. Conditioned [387]. conducted [12]. Conne [676]. Connectivity [657, 682, 226]. consensus [72]. Consequences [504, 375, 316]. Conservation [354, 470, 553, 176, 469, 613, 460, 192, 621, 53]. Conserving [544]. Considerations [271, 371, 633, 658]. Considering [642]. Constraints [695, 410, 562]. Construction [70]. Consumption [642, 569, 667]. Contaminants [647, 59]. Contemporary [227]. Contents [112]. Context [384]. continuing [18]. Contributing [676]. Contribution [643]. contributions [52]. Control [471, 265, 563, 91, 456, 199, 625, 275, 542]. Controlling [395]. Controversy [120]. Conventional [115]. Cooperative [559]. Coordinated [362]. Copepoda [263]. Copepodid [170, 175]. Copepods [636]. Coping [370, 571, 375]. Coral [160]. Coregonus [278]. Correction [671]. Corrections [311]. Correlations [109]. Corrigendum [305, 481]. corruscans [386]. Coryphaena [567, 263, 311, 303]. Coryphaenidae [263]. Cost [397]. Cost-Benefit [397]. Costs [200]. Council [360]. Countries [551]. country [70]. COVID [666, 610]. COVID-19 [666, 610]. Crab [228, 255, 225, 241, 223, 87, 492, 113, 250, 125, 217]. Crabs [234, 242]. Crappie [265]. Crassostrea [410]. Crayfish [315, 532, 306, 279, 584, 35]. Creatine [650]. Creek [614]. Creel [600]. Criteria [436]. Critical [87, 672, 508, 550, 8]. Criticisms [629]. Croaker [254]. crocea [254]. Crosstalk [685]. Crucial [694]. crustacea [44, 315]. Crustacean [203, 77, 76, 78, 79]. Crustaceans [626, 233, 507, 375, 467, 364, 48]. Cryopreserved [606]. Cucumaria [594, 372]. Cucumber [594, 462, 372, 351]. Cucumbers [237]. Cultivars [631]. Cultivation [592, 695]. Culture [281, 144, 400, 107, 148, 453, 331, 218, 91, 131, 254, 514, 137, 378, 130, 15, 135, 20, 30, 8, 7, 43, 38, 46]. Cultured [230, 691, 401, 248, 237, 569]. Cumulative [299]. Current [586, 102, 255, 477, 480, 574, 501, 697, 594, 439, 209, 344, 516, 492, 588, 405, 290, 549, 426, 610, 496, 595, 135, 334, 557]. Currents [170, 175]. Cutting [689]. Cutting-Edge [689]. Cycles [234]. Cycling [356]. Cyclopsetta [154]. Cynoscion [493]. Cyprinid [270, 591]. Cyprinidae [181]. Cyprinids [318]. Cyprinus [436]. Cyst [665]. Cyst-Producing [665]. Cytochrome [302].

Daily [332]. Dam [521]. Dams [188]. Data [622, 383, 425, 365, 284, 325, 493, 494, 334, 47]. Data-Limited [325]. Data-Poor [365]. Dealing [667]. Debate [509]. Decade [332, 624]. Decades [536, 385, 567, 639]. Decapod [626, 233, 375, 467]. Decapoda [315]. Decline [615, 676, 358]. Declines [260]. Deep [400, 452, 382, 257, 654]. Deep-Sea [452, 654]. Deep-Water [382]. Deepwater [365]. Defense [583]. Defining [108]. Definition [366]. Delaware [195]. Delbert [651]. Delineating [227]. Delineation [620]. Delivery [491]. Delta [358]. Demand [429]. Demands [693]. Demersal [337]. Demonstrate [636]. Density [238, 390, 11]. Density-Dependence [238]. Density-Dependent [390]. Dependence [238]. Dependent [425, 237, 390]. Depleted [394, 399, 218, 397]. Deposition [635, 272]. Depredation [459]. Depth [278]. Derived [411]. Design [620, 607, 225, 28, 336]. Designation [421]. Designs [291]. Detection [570, 69]. Determinants [656, 536, 190]. Determination [383, 206]. Determine [603, 68]. Determining [391, 247]. Developed [523]. Developing [573, 13, 70]. Development [338, 422, 271, 681, 292, 418, 532, 486, 437, 510, 100, 507, 385, 329, 137, 513, 652, 248, 496, 611, 617, 11, 640, 75]. Developmental [298]. Developmentally [314]. Developments [577, 634]. Diagnosis [639]. Dicentrarchus [345]. Dichotomy [623]. Diet [251]. Dietary [568, 590, 472, 48, 345, 119, 151]. Diets [264, 78, 100, 91]. Differ [242]. Differentiation [545, 207]. Difficult [478]. digestion [64]. Digestive [77]. Dimensions [427]. Dinoflagellate [665]. Diptera [270]. Directed [104]. Directions [407, 385, 641, 684]. Discard [114]. Discards [347]. Discharge [146]. Disciplinary [220]. Discriminant [603]. discus [245]. Discussion [139]. Disease [171, 471, 212, 173, 563, 128, 160]. Diseases [213, 294, 628, 44]. Disentangling [476]. Disparate [660, 494]. Dispersal [289, 389]. Disrupt [665]. Distribution [170, 175, 301, 635, 306, 96, 309, 251]. Ditches [179]. Dive [346]. Diverse [559]. Diversification [691, 514]. Diversity [656, 386, 631, 482, 335]. d’Ivoire [625]. Do [242, 368]. Documented [192]. Does [73, 261, 348, 242]. Dogma [678]. dolomieu [660]. Dolphin [311, 303]. Dolphin-Isopod [303]. Dolphinfish [567]. Dominant [322, 300]. Dorsal [424]. Dosidicus [444]. downriver [3]. Downstream [684, 65]. downstream-migrating [65]. drainages [23]. Drive [531]. Driven [340]. Drivers [468]. Drop [343]. Drum [398]. Due [355, 381, 31]. During [272, 410, 91, 666, 625, 639]. Dusky [244]. Dyes [552]. Dynamic [307]. Dynamics [339, 401, 510, 597, 508].

Early [225, 52, 59]. East [362, 254, 335]. Eastern [410, 198, 424, 240, 221, 658]. Ecological [466, 184, 411, 356, 393, 659, 531, 597, 600, 674, 468, 605, 395, 300, 62]. Ecological-Fishery [597]. Ecology [354, 427, 278, 553, 181, 594, 340, 377, 675, 578, 350, 614, 618, 627, 342, 550, 46]. Economic [202, 232, 340, 502, 426, 513, 286, 533, 185]. Economically [364]. Economics [230, 454]. Economies [612]. Economy [664]. Ecophys.Fish [159]. Ecosystem [592, 320, 451, 501, 695, 108, 325, 604, 578, 562, 25]. Ecosystem-Based [320]. Ecosystems [430, 699, 382, 537, 325, 436, 300, 4, 35, 75]. Ecuador [444, 85]. Edge [689]. Edition [369, 651, 692]. editor [16, 1, 5, 10, 19, 33, 45, 49, 54, 57, 60, 63, 67, 71]. Editorial [357, 406, 428, 442, 465, 488]. edwardsii [235]. Eel [534, 131, 308]. EEZ [646]. Effect [438, 570]. Effective [186]. Effectiveness [646, 240, 28]. Effects [359, 142, 170, 175, 65, 117, 292, 370, 393, 541, 105, 127, 277, 182, 23, 375, 167, 436, 342, 300, 647, 608]. Efficacy [630, 538]. Efficiencies [645]. Efficiency [203]. Effluent [116]. Effluents [117, 588]. Effort [193, 452, 478]. Egg [288, 438, 332, 59]. Egg-To-Fry [288]. Eggs [246, 13]. Egypt [582, 496]. Element [589]. Elops [340]. Emerging [691]. Emigrating [309]. Emigration [390]. Emphasis [117, 534, 503, 493, 65, 61]. Empirical [520]. Emptiness [336]. Encircling [317]. enclosed [4, 75]. Endangered [184, 182, 421, 169, 557, 244]. Endocrine [210]. Endogenous [302]. Enemies [659, 62]. Engagement [449]. Engineering [607]. England [327]. Enhance [156]. Enhancement [228, 222, 253, 215, 398, 255, 225, 243, 233, 404, 462, 392, 393, 396, 234, 331, 223, 385, 216, 321, 384, 66, 258, 500, 250, 183, 229, 231, 227, 220]. Enhancing [399, 218]. Enter [384]. Environment [230, 563, 149, 522, 647]. Environment-Friendly [563]. Environment-Habitat-Production [149]. Environmental [203, 539, 42, 271, 537, 360, 128, 588, 159, 206, 275, 591, 183, 688, 59]. Environmentally [446]. Environments [150, 177]. Enzymes [302]. EOV [357, 406, 428, 442, 465, 488]. Epinephelinae [378]. Epinephelus [43, 244]. Episodic [408]. Epizootic [566]. Epizootics [160]. Equation [424]. equiselis [311]. Era [215, 625, 618]. Eriocheir [255]. Erosion [150]. Errata [9]. Erratum [164]. Erythrinidae [341]. Escapement [328]. Escapes [504, 246]. Eschscholtz [315, 279]. Essential [693, 61]. Establishment [525]. Estimates [420]. Estimating [284, 162, 291]. estimation [3]. Estuaries [149]. Estuarine [343, 234, 299]. Estuary [322, 358, 248, 403, 356]. Ethical [197]. Ethics [505]. Ethiopia [460]. Eucheumatoid [631]. Europe [441, 695, 663, 432]. European [354, 345, 135]. EUS [566]. Evaluate [193, 596, 421]. Evaluating [604, 683]. Evaluation [225, 269, 396, 402, 456, 277, 360, 240, 236, 280, 4]. Events [665]. Evidence [259, 520, 524, 436]. Evil [347]. Evolution [680, 593, 640]. Evolutionary [468, 627]. Evolving [384]. Examining [391, 112, 211]. Example [138, 367, 296]. Examples [291, 627]. Exercise [655]. Existence [125]. Exogenous [245]. Exotic [51, 50]. Experiences [513]. Experimental [588, 436]. Experimentation [449]. Experiments [239, 544, 12]. explain [18]. Exploitation [594, 26]. Exploratory [190]. Exploring [693]. Export [546]. Exposed [296]. Exposure [296]. Expression [302]. External [270]. Extinction [259, 649]. Extracellular [624]. Extractive [580]. Extremely [614]. Exxon [58, 359, 120].

Faceted [335]. Facilities [684, 3]. facility [70]. Facing [502]. Factor [694, 547]. Factors [545, 676, 358, 375, 467, 280]. Fails [259]. Fall [441, 148]. False [623]. Family [691, 249]. FAO [643]. Farmed [205, 208, 246, 595]. Farming [116, 129, 138, 134, 448, 515, 457, 256, 420, 140, 326, 6, 461, 167, 644, 324, 56]. Farms [170, 175, 174, 535]. Fattening [326]. Fatty [635, 693, 145, 61]. Feared [660]. Feasibility [116, 232]. Featherback [514]. Fecundity [596]. Fed [686]. Federation [274]. Feed [668, 680, 541, 36, 687, 429, 643, 20]. Feed-Based [668]. Feeding [391, 144, 310, 565, 91, 295, 473, 249, 118, 20, 12]. Feeds [472, 673, 290, 611, 617]. Feedstuffs [264]. Feel [368, 678]. Feral [318]. Fertility [98]. Field [76, 559, 544]. Fifty [373]. Fillets [635]. Fin [150, 209]. Fin-Fish [209]. Fine [288]. Finfish [381, 208, 530, 563, 535, 503, 324]. fingeriings [36]. Fingerling [256]. fingerlings [61]. Firms [190]. First [424, 410, 106]. Fish [560, 592, 586, 302, 98, 230, 620, 338, 264, 281, 203, 539, 142, 422, 612, 205, 117, 696, 477, 292, 212, 561, 501, 669, 371, 322, 339, 173, 119, 439, 583, 509, 536, 272, 112, 247, 519, 166, 174, 566, 333, 606, 209, 537, 367, 552, 296, 341, 211, 571, 141, 478, 329, 473, 159, 368, 147, 160, 330, 206, 299, 207, 517, 214, 286, 438, 413, 94, 291, 415, 687, 440, 151, 285, 312, 376, 490, 569, 667, 643, 698, 390, 450, 506, 145, 650, 491, 508, 688, 53, 73, 48, 2, 27, 12, 40, 64, 74]. fish [69, 56, 13, 51, 37, 538, 61, 75, 374, 678]. Fisheries [317, 320, 430, 435, 466, 338, 407, 215, 102, 585, 573, 398, 409, 553, 441, 425, 637, 347, 521, 365, 418, 443, 682, 343, 369, 663, 437, 648, 261, 279, 536, 361, 313, 483, 519, 283, 189, 699, 459, 382, 223, 340, 516, 367, 87, 216, 384, 642, 254, 360, 531, 346, 373, 325, 505, 598, 419, 379, 600, 101, 426, 630, 685, 604, 665, 605, 122, 126, 474, 493, 518, 500, 575, 639, 558, 587, 165, 621, 653, 316, 672, 533, 231, 658, 627, 342, 363, 300, 121, 683, 540, 25, 66, 24, 47, 7, 39]. Fisheries-At [699]. Fisheries-Important [682]. Fishery [228, 657, 354, 425, 224, 92, 479, 108, 460, 190, 456, 115, 346, 675, 88, 567, 444, 597, 337, 455, 286, 104, 366, 572, 511, 658, 4, 52]. Fishes [464, 93, 485, 662, 697, 596, 613, 91, 267, 581, 277, 273, 619, 468, 180, 124, 200, 408, 118, 336, 570, 550, 23, 50]. Fishing [202, 197, 577, 193, 414, 645, 574, 411, 452, 417, 663, 646, 519, 192, 537, 348, 478, 337, 527, 630, 147, 513, 524, 285, 324, 679, 53]. Fishy [623]. fistula [301, 263]. Fix [644]. Flatfish [129, 331, 210]. Flatfishes [419]. Flavor [103]. fleet [31]. Floodplain [575]. Florida [398, 50, 94, 95, 96, 97, 308, 220, 276]. Flounder [327, 243, 565, 541, 675, 130, 231, 387, 59]. Flow [343, 188, 167, 185]. Fluctuation [404]. Fluctuations [348, 147]. Fluid [564]. Fluorochrome [552]. Focus [664, 40]. Focused [600]. Following [120]. Food [686, 322, 356, 105, 666, 95, 285, 376, 490, 643, 672, 698, 533, 2]. Foods [656, 376]. Forage [285]. Forecasting [143, 597]. Fork [413]. Formation [208]. Forms [568]. Formulated [91]. formulation [20]. Forty [346, 544]. Forward [529, 500]. Framework [200, 440]. Frameworks [516, 350]. France [77]. Francisco [386, 322, 356]. free [30, 13]. Frequency [570]. Fresh [497]. Freshwater [466, 623, 86, 148, 343, 294, 306, 279, 581, 341, 584, 457, 531, 503, 564, 440, 436, 20, 23, 47]. Friendly [563, 446]. frondosa [594, 372]. Fry [288, 291, 75]. Fukushima [231]. Function [563, 211]. Functional [208, 287, 214]. Functions [586, 302, 533, 145, 363]. Fungus [566]. Further [440]. Future [407, 355, 381, 553, 76, 480, 574, 695, 439, 313, 123, 483, 166, 581, 385, 457, 377, 140, 290, 641, 137, 122, 126, 562, 496, 684, 611, 617, 135]. Futures [693].

Gabes [310, 198]. Gadus [316]. Gain [363]. Game [677]. Gamete [208]. Ganglioneuritis [397]. Gaps [573, 470, 695, 675, 684]. Gas [292, 513]. Gastronomy [601]. Gatlin [651]. Gear [417, 630, 53]. Gear-Switching [630]. Gears [645]. Gene [211]. General [443, 139]. Generalized [395]. Genetic [265, 386, 682, 403, 369, 638, 433, 182, 206, 111, 335, 227, 121, 690, 37]. Genetics [476]. Genomic [212, 211]. Genomics [205, 208, 213, 699, 207, 214, 595]. Genotoxic [277]. Genus [485, 154, 538]. Geographic [603]. Geologic [292]. Geomorphology [180]. Germplasm [266]. Ghana [426]. Ghost [337]. Giant [484, 26]. gigas [444, 524, 621]. Gilbert [52]. Gill [278, 626]. Gillnet [317, 630]. Gills [318]. Gilthead [345]. gladius [474, 649]. Global [656, 197, 577, 355, 381, 580, 696, 686, 477, 451, 645, 551, 579, 660, 588, 420, 567, 330, 652, 664, 610, 440, 653, 285, 312, 376, 548, 643, 698, 436]. Glucocorticoid [422]. GnRHa [199]. Goals [40, 621]. Gonadal [507, 248]. Gonadogenesis [206, 207]. Gonadotropin [199]. Gonadotropin-Releasing [199]. Gonorynchiformes [323]. Gracilaria [298]. Gracilarioid [110]. Grain [673]. Grande [178, 295, 183, 185, 181, 184, 182]. gratilla [218]. Great [278, 117, 498, 509, 578, 200, 282, 350]. Green [499]. Greenlip [392, 393, 396, 346]. groundfish [41]. Grounds [646]. Groundwater [179]. Grouper [546, 378, 244, 43]. Grow [218]. Grow-Out [218]. Growout [139, 128]. Growth [612, 391, 424, 526, 551, 392, 272, 448, 507, 402, 159, 652, 248, 198, 297, 282, 118, 73, 12]. Guidelines [611, 617, 683]. Gulf [53, 269, 310, 88, 160, 198, 268, 494]. Gulf-War [88]. Gunnerus [372]. Gut [515, 473].

HAB [665]. Habitat [391, 155, 425, 489, 389, 149, 299, 614, 268, 415, 251, 390, 168, 362, 32]. Habitats [195, 571, 250, 53]. Habits [310, 249]. haematalis [319]. Hair [113]. Hake [455]. Halibut [618, 15]. Haliotis [394, 245, 402]. Hamilton [514]. Harmful [127]. Harvest [456, 408, 104]. Harvesting [528]. Hatchery [255, 245, 247, 402, 389, 405, 249, 182, 275, 236, 544, 395, 242]. Hatchery-Reared [245, 402, 389, 249, 544, 242]. Hawaii [148, 249]. Health [590, 367, 628, 641, 569, 667, 698, 608]. Healthy [515]. Helminths [341]. Hemisphere [140]. Hemoflagellates [319]. Hemogram [319]. Henry [52]. Hepatopancreatitis [196]. Herbal [576, 563]. Herbs [561]. Heritage [338]. Heterogeneity [531]. Hierarchical [430]. Highly [523]. Highways [168]. Hindcasts [355, 381]. Hippoglossus [15]. hippurus [567, 263, 311]. Histological [318]. Historic [426]. Historical [197, 162, 405, 562, 334, 476]. History [228, 109, 269, 670, 201, 200, 350, 353, 99, 542]. Hokkaido [405]. Holistic [680]. Holothuria [401]. Homarus [22]. Home [388]. Hooking [24]. Hoplias [341]. Hormonal [265, 275]. Hormone [422, 199]. Horseshoe [125]. Host [547]. Host-Associated [547]. Hosts [301]. hoyi [278]. Hudson [647]. Human [427, 579, 460, 376, 569, 679, 4, 37]. Humped [514]. Hybognathus [181, 295, 182]. Hybrid [98]. hydroacoustic [3]. hydroeiectric [3]. Hydroelectric [521]. Hydrology [180]. hydrophila [173]. Hydropower [188, 684]. Hypomesus [358]. Hypotheses [278]. Hypothesis [320, 259, 155].

Ichthyological [694]. Ide [591]. Identification [425, 423, 412, 602, 303, 121]. Identified [239]. Identifying [642]. idicus [80]. idus [591]. II [393, 207, 95]. III [396, 96, 651]. Images [646]. Immune [694, 213, 439, 563, 211]. Immunological [697, 566]. Immunology [212]. Immunomodulators [576]. Immunoreactive [541]. Immunostimulants [541]. immunostimulatory [56]. Impact [622, 394, 245, 288, 229, 395, 7]. Impacts [585, 411, 417, 519, 460, 89, 192, 537, 598, 337, 426, 513, 299, 120, 562, 610, 432, 185, 168, 53, 58, 4, 37]. Impending [615]. Implants [199]. Implement [183]. Implementation [516]. Implementing [14]. Implication [421]. Implications [655, 504, 202, 53, 528, 354, 144, 583, 401, 367, 588, 473, 299, 180, 578, 467, 562, 687, 366, 312, 533, 336, 363]. Importance [228, 580, 269, 241, 463, 226, 268, 376]. Important [603, 501, 682, 296]. Improved [271, 128, 569]. improvement [32]. Income [490]. Incorporating [339, 307]. Increase [232, 452, 672]. Increasing [264]. Indeterminate [596]. Index [153, 421]. India [338, 437, 287, 562, 575]. Indian [338, 500]. Indicator [319]. Indicators [203, 519]. Indices [251, 11]. Individual [468, 39, 608]. Indonesia [546]. Indoor [400]. Induced [460]. Induction [275]. Industrialization [654]. Industrialized [527]. Industries [492, 513]. Industry [134, 123, 344, 502, 513, 446, 679]. Ineffectiveness [183]. Infecting [485]. Infections [561]. Infectious [294]. Infective [170, 175]. Infectivity [319]. Inferred [181, 433, 690]. Infestations [174]. Inflammation [609]. Influences [558, 282, 602, 47]. Influencing [545, 375, 467]. Information [266, 555, 330, 165]. Informing [509]. Ingredient [680]. Ingredients [673]. Inhibins [560]. Inland [381, 483, 240, 558]. Innovation [637, 510, 599]. Innovations [107, 131]. Insights [560, 682, 192, 652, 496, 300, 362]. Institutional [72, 280]. Instream [185]. Insurance [525, 593]. Intake [687]. Integrated [264, 457, 531, 476]. Integrating [605]. Integration [539]. Integrative [197, 398]. Intensification [612]. Intensive [91, 20, 36]. Interactions [504, 422, 179, 674]. Interdisciplinary [264, 327]. Interest [293]. Intermediate [400]. Intermuscular [641]. International [577, 262, 55]. Internet [634]. Interpretation [602]. Intersections [87]. Introduced [388, 94, 95, 96, 97]. Introduction [432, 204]. Introductions [451, 584]. Introductory [651]. Introgression [98]. invadans [566]. Invader [660]. Invasion [370]. Invasive [498, 356, 469, 632]. Invasiveness [591]. Invertebrate [409, 346, 299]. Invertebrates [281, 662]. Investigation [358]. Investigations [333]. Involved [360]. IoT [634]. Irrigation [179]. Isopod [303]. Issues [172, 380, 107, 437, 298, 575]. IV [97]. Ivory [625]. IWGCN [76].

Japan [404, 27, 256, 405, 240, 231, 387]. Japanese [534, 240, 257, 231, 387]. japonica [255, 534, 257]. japonicus [404, 229, 82]. Jasus [235]. Jellyfish [585]. Joaquin [358]. Jones [260]. Jumbo [444]. Juvenile [238, 225, 489, 188, 174, 91, 239, 235, 236, 309, 65]. Juveniles [345, 245].

kanagurta [646]. Kattegat [316]. Keenjhar [318]. Kelp [528]. Kelt [684]. Key [547, 475]. Kill [333]. Kills [141, 27]. Kinetics [90]. King [87, 492]. Klunzinger [198, 297]. Knowledge [586, 197, 573, 470, 574, 501, 695, 594, 469, 439, 638, 675, 684, 557, 52]. Korean [190]. Krill [108]. Krkosek [175]. Kuruma [404]. Kuwait [88].

L. [383, 424, 441, 431, 447, 345, 251]. Laboratory [544, 40, 70]. labrax [345]. Lactic [554]. Ladyfish [340]. Lake [202, 318]. Lakes [156, 117, 498, 21, 278, 578]. Lambari [486]. Laminariales [528]. Lamprey [469, 542]. Lands [176]. Landscape [531, 168, 540]. Large [254, 377, 450, 75]. Largemouth [660, 36]. Larvae [144, 13, 61]. Larval [657, 222, 281, 422, 78, 79]. Last [567, 639]. Latent [511]. Latest [624]. Latin [464, 691]. Laurentian [278, 578]. Learned [219, 332]. Learning [699, 511]. Lectins [530]. Led [448]. Legislation [581]. Length [431, 447, 96]. Lepeophtheirus [170, 175]. leptodactylus [315, 279]. Lesions [318]. Less [612]. Lessepsian [198]. Lessons [219, 134, 525, 332]. Leuciscus [591]. Level [685, 312]. Lice [170, 259, 174, 260]. Life [228, 109, 269, 234, 670, 665, 201, 200, 350]. Light [527, 683]. limit [73]. Limitations [398, 11]. Limited [325]. Limits [161, 395, 227]. Limulus [125]. Lingcod [389]. Linkages [195, 149]. linked [64]. Linking [468]. Linnaeus [556, 453, 263]. Lionfish [370]. Lipases [267]. Lipid [635, 272, 209]. Lipidomics [608]. Lipids [506, 145]. Literature [293, 156, 157, 163, 187, 191, 194, 304, 311]. Litopenaeus [515, 588]. Livelihood [533]. Living [348]. liza [412]. Llanquihue [202]. Loading [322]. Loads [355, 381]. Lobatus [524]. Lobster [144, 224, 388, 257, 235, 22]. Lobsters [682, 232, 28]. Locations [239]. Loci [423, 69]. locomotion [40]. Locus [247]. Long [394, 676, 635, 693, 322, 392, 524]. Long-Chain [635, 693]. Long-Term [676, 322, 392, 524]. Longline [459]. Loss [459, 18]. Lost [343]. Lota [670]. Lotic [105]. Louse [175]. Low [685]. Low-Trophic-Level [685]. Lower [696]. Luminous [480]. Lutjanus [275]. Lysine [119].

M [651]. Machine [699, 511]. Mackerel [240]. Macro [502]. Macro-Economic [502]. Macrobrachium [20, 86, 148]. Madeira [521]. magellanicus [366]. Magnesium [520]. Magnify [348]. Magnitude [238]. Mahseers [476]. Main [691, 298]. Maine [53, 494]. Major [675]. Makers [623]. Making [520]. malabaricus [341]. Malaysia [476]. Malaysian [646]. Mammals [92, 267]. Management [622, 228, 657, 320, 430, 504, 202, 171, 102, 573, 116, 354, 409, 427, 553, 176, 425, 365, 418, 682, 224, 92, 469, 261, 173, 108, 313, 460, 546, 382, 192, 581, 516, 416, 456, 367, 87, 384, 256, 675, 405, 632, 555, 379, 600, 604, 111, 180, 500, 575, 639, 408, 268, 415, 653, 316, 262, 231, 363, 300, 540, 25, 14, 41, 39, 55]. Managers [105]. Managing [466, 216, 186]. Mangrove [223]. Mangrove-Associated [223]. Manila [219]. Manipulations [238]. Mapping [494]. marginatus [244]. Mariculture [110, 652, 664, 593, 654, 26]. Marine [222, 281, 585, 398, 380, 623, 574, 107, 626, 92, 680, 673, 339, 417, 636, 313, 189, 454, 402, 537, 416, 226, 385, 321, 384, 254, 325, 329, 140, 503, 564, 426, 604, 299, 482, 633, 639, 220, 324, 608, 276, 629, 244, 48, 66, 13, 17, 75, 360]. marinus [469]. marker [69]. marker-assisted [69]. Markers [433]. Market [529, 138]. Marketing [677]. Marking [552, 291]. Marshes [195]. Marshyhope [614]. Mass [138, 636, 608, 75]. Maternal [422]. Matter [261, 242]. Matters [376, 429, 490]. Maturation [596, 199, 42]. Maturing [134, 384]. Maturity [556]. maxima [251]. Maximum [695, 161]. Meal [118]. Means [486]. Measures [638]. Mechanism [520]. Mechanisms [146, 463, 206, 375]. Media [505]. Medicinal [561]. Mediterranean [452, 310, 198, 425, 411, 367, 567, 326, 446, 297, 351]. Mediterranean-based [446]. Mekong [696, 613]. Membership [559]. Membrane [64]. Membrane-linked [64]. Mercury [497]. merguiensis [80]. Mesocosms [236]. Meta [545, 383, 603, 288, 606, 535, 564, 591]. Meta-Analysis [383, 603, 288, 606, 535, 564, 591, 545]. Metabolic [13, 350]. Metabolism [507, 145]. Metagenomic [588]. Metamorphosis [210]. Metanephrops [257]. Metapopulation [226]. Method [332]. Methodological [620]. Methods [669, 112, 463, 456, 284, 689, 670, 513, 467, 68, 291, 450, 550]. Methylmercury [296]. Metrics [681, 680]. Mexico [184, 266, 269, 178, 160, 268]. Microbiota [515]. Microchemistry [464]. Micronucleus [296]. Microparticulate [78]. Micropterus [660]. Middle [184]. Migrant [198]. migrating [65]. Migration [523, 696, 613, 684, 602, 22, 3]. Migratory [234]. Milestones [675]. Milkfish [132, 43]. Millennia [286]. Millennium [123]. Million [125]. Mimicry [601]. Minas [421]. mineral [48]. Minerals [506]. Minnow [181, 184, 295, 182, 185]. Mississippi [46]. mississippiensis [495]. Missouri [169]. Mitigate [683]. Mitigating [89, 467]. Mitten [255]. Mixed [653]. mobile [53]. Model [531, 159, 229, 395]. Modeling [224, 339, 688]. Modelling [635]. Models [648, 325, 604, 511, 307]. Modern [173, 540]. Modulation [515]. Molecular [532, 213, 433, 207, 210, 412, 314, 244]. Mollusca [335]. Molluscan [664]. Molluscs [90, 34]. Mollusks [294]. Monitoring [528, 519, 382, 450]. monodon [81]. Monogenean [561]. Monopterus [308]. Montana [413]. mordax [632]. morhua [316]. Morphological [315]. Morphology [77, 453, 476, 570]. Mortality [655, 284, 236, 364, 24, 18]. mosquito [31]. Motility [606]. Mountain [32]. Movement [339, 284, 252, 468, 614, 688]. Movements [690]. Moving [680, 377]. MSC [455]. Mucus [669]. Mugil [412, 43]. Mullet [412, 43]. Mulloway [229]. Multi [430, 335, 220]. Multi-Disciplinary [220]. Multi-Faceted [335]. Multi-Species [430]. Multidisciplinary [217]. Multiple [416]. Multiple-Use [416]. Multisensory [539]. Mussels [294, 659, 62]. Myiasis [270]. Mystery [303]. Myths [629]. Myxobolus [289]. Myxozoan [146].

n.sp [319]. Nanoparticles [491]. Nanticoke [614]. Narragansett [147]. National [183]. Native [289, 352, 469, 581, 660, 674, 180, 413, 202]. Natural [292, 105, 659, 460, 89, 150, 62, 544]. Naturalization [432]. Nature [124]. Nautilus [335]. Navigating [679]. Nearshore [489]. Necessary [347]. Necrotizing [196]. Need [672]. Needed [637, 636]. Needs [695]. Nekton [195]. Neotropical [606, 296, 341]. Net [246, 114]. Neurobehavioral [124]. Newfoundland [676]. Nexus [499]. NHPB [196]. Nigeria [582]. niloticus [400, 399]. Nitrate [375]. Nitrite [375]. Nitrogen [355, 381]. nitrogenous [29]. Non [202, 289, 212, 536, 581, 660, 284, 674, 413, 285]. Non-catch-related [536]. Non-Food [285]. Non-Native [289, 581, 660, 674, 413, 202]. Non-Reporting [284]. Non-Salmonid [212]. Nonlethal [112]. Nordman [301]. Nordmann [263]. North [316, 615, 347, 510, 493, 690]. Northeast [382, 252]. Northern [449, 405, 372, 337, 200, 350, 441]. Northwest [53, 41, 352, 452, 143, 161, 618, 309, 46, 121]. Norway [492]. Norwegian [72]. note [1, 5, 10, 16, 19, 33, 45, 49, 54, 57, 60, 63, 67, 71]. Novel [673, 413]. Numerical [395]. Nursery [250]. Nutrient [612, 322]. Nutrigenomics [209]. Nutrition [83, 80, 76, 579, 534, 565, 290, 85, 79, 376, 601, 82]. Nutritional [81, 84, 20, 86, 100, 569, 533, 506]. Nutritive [507].

Objectives [92]. Occurrence [90, 494]. Ocean [404, 644, 615]. ocellaris [94, 95, 96, 97]. Ocellate [239]. Ochrophyta [528]. Octopus [587]. Off [103, 327, 646, 346, 332]. Off-Flavor [103]. Offering [691]. Offshore [269, 535, 513, 640]. Oil [472, 359, 513, 120, 58]. Old [303]. olivaceus [565, 541, 387]. Olive [565, 541]. Omega [693]. Omega-3 [693]. Oncorhynchus [52]. One [247, 628]. One-Locus [247]. only [34]. Ontogenetic [389]. Oocyte [596, 272, 438]. Oocytes [272]. Operation [266, 70]. Operational [681]. opilio [492]. Opportunities [691, 470, 697, 437, 642]. Optimal [391, 239, 229]. Optimizing [620]. Options [313, 287]. Oregon [289, 162]. Öresund [316]. Organisms [127, 30]. Organization [379, 17]. orientalis [254, 555]. Origin [603, 538]. Originating [170, 175]. Ornamental [281, 380, 329, 276]. Osmerus [632]. Ostariophysan [323]. Other [195, 269, 294, 200, 408, 461, 59]. Otolith [464, 620, 520, 248, 543]. Otolith-Stained [248]. Otoliths [589]. Otter [115]. Outcomes [531]. Outlets [188]. Outlook [344, 672]. Ovarian [560, 564]. Overfishing [524]. Overlap [388]. Overview [590, 498, 453, 616, 514, 503, 500, 684, 450, 646]. Oxidative [583]. Oyster [410, 549, 6].

P450 [302]. Pacific [655, 352, 526, 52, 245, 288, 143, 515, 555, 249, 455, 674, 161, 309, 15, 121]. Paddlefish [200, 282, 350, 408, 649]. Pagrus [310]. Pain [662, 368, 124, 678]. Palinuridae [682]. Pallid [169]. Pampelfisk [303]. Panarchy [632]. Pandemic [666]. Pangasius [502]. Papers [667]. Paralichthyidae [154, 518]. Paralichthys [565, 541, 518, 387]. Paralithodes [492]. Paralytic [90]. parameters [68]. Parasite [289]. Parasites [485, 659, 374, 311, 62, 44]. Parasitic [341, 263]. Parasitized [319]. Part [392, 393, 396, 206, 207]. Participation [577]. Parvalbumins [2]. Passage [684]. Passing [188]. Past [76, 483, 377, 140, 122, 126, 3]. Path [572]. Pathobiology [566, 374]. pathogen [30]. pathogen-free [30]. Pathogenicity [566]. Pathogens [477]. Pathway [276]. Pattern [317]. Patterns [278, 109, 589, 309, 362]. Peacock [94, 95, 96, 97]. Pecos [178]. Pelagic [322, 459, 285]. Penaeid [86, 196]. Penaeids [83]. Penaeus [80, 81, 84, 404, 254, 85, 82]. Peniculus [301, 263]. Pennelidae [263]. Pens [246]. People [628, 99]. Perception [605]. Perceptions [446]. Percid [697]. Perciformes [493]. Performance [203, 681, 387, 654, 640]. Persistence [120]. Perspective [694, 571, 562, 610, 639, 268, 285, 436, 41]. Perspectives [586, 197, 577, 259, 480, 574, 439, 677, 290, 496]. Peru [521]. Petroleum [269]. Petromyzon [469]. Phenological [602]. philippinarum [219]. Philippine [399]. Philippines [631, 402, 8]. Phospholipid [608]. Phosphorus [355, 381, 522]. Phycotoxin [34]. Phycotoxin-related [34].

Phylogeographic [412]. Physical [681]. Physiological [636, 375]. Physiology [560, 93, 142, 77, 473]. Pike [449]. Pilot [225, 525]. Pink [359, 259, 260, 120, 58]. Pinniped [104]. Pisces [485]. Pisciculture [338]. PL [608]. Placopecten [366]. plain [23]. Plains [200, 282, 350]. Planning [264]. Plans [280]. Plant [264]. Plants [561]. Platforms [269]. Plausible [520]. Pleuronectes [59]. Pleuronectiformes [154, 518]. poisoning [34]. Polar [146]. Policies [593, 640]. Policy [623, 663, 581, 633, 186, 183]. Policy-Makers [623]. Political [454]. Pollution [668]. Polydactylus [249]. Polynemidae [249]. Polyodon [200]. polyphemus [125]. Polyphenol [590]. Polyphenol-Rich [590]. Polyphenols [590]. Polyunsaturated [635]. Pompano [311]. pompilius [335]. Pond [107, 236, 8]. Ponds [156, 134, 21, 46]. Poor [365]. Poorly [591]. Poorly-Studied [591]. Population [545, 158, 615, 682, 339, 393, 401, 433, 348, 286, 408, 450, 307, 690]. Populations [222, 394, 104, 335, 524, 227, 53, 37]. Portugal [574]. Portuguese [478, 474]. Possibilities [554, 32]. Possible [615, 475]. Post [281, 402, 88, 236, 387]. Post-Larval [281]. Post-Release [402, 236, 387]. Potential [289, 193, 668, 686, 425, 292, 561, 661, 695, 233, 646, 192, 537, 457, 473, 372, 337, 549, 625, 591, 575, 200, 351, 494, 512]. Potentially [676, 541]. Potentials [398, 514]. Pots [113]. Poverty [625]. Powder [292]. Power [414]. Practical [331]. Practice [333, 599]. Practices [116, 446, 634, 20, 30]. Prawn [86, 148, 404, 254, 457, 20]. Pre [409]. Pre-Recruit [409]. Prebiotic [583]. Prebiotics [541, 439]. Precipitation [178]. Predator [235]. Predators [243, 659, 62]. Predatory [229, 395, 390]. Predicting [646, 364]. prediction [27]. Predictions [409]. Predictive [427]. Preface [434, 176]. Preference [295]. Preparation [199]. Presence [142, 570]. Present [76, 123, 483, 140, 122, 126]. Pressure [524]. Prevalence [467]. Prevention [128, 21]. Prince [58, 359]. Principles [692]. Prior [141]. Priorities [573]. Prioritization [440]. Probiotic [583]. Probiotics [547, 541, 463, 538]. Problems [686, 460, 140, 103]. Procedures [142]. Processes [468]. Producer [559]. Producing [110, 665]. Product [502]. Production [580, 686, 343, 636, 642, 256, 346, 373, 484, 149, 548, 672, 601, 332, 133, 36, 75]. Products [264]. Profile [639]. Program [404, 525, 169, 244, 525]. Programs [405]. Progress [253, 265, 516, 677, 298]. Project [183]. Projections [355, 381]. promising [538]. Promoting [623, 469]. Proof [588]. Propagation [182]. Properties [636, 665]. Propolis [512]. Prospect [677]. Prospective [298]. Prospects [399, 166, 209, 500, 575, 640]. Protect [188]. Protected [568]. Protection [374, 185]. Protein [673, 698]. Proteins [506]. Proteomic [208]. Provided [695]. Psephurus [649]. Psetta [251]. Pseudoplatystoma [386]. Pseudosciaena [254]. Public [577, 449]. Published [383, 288]. puerulus [28]. Puffer [239, 236]. Pugheadedness [619]. Purposes [555]. Purse [478]. puzzle [18]. Pyridoxine [586].

Qadri [319]. QTLs [423]. Quagga [659]. Qualitative [289, 325]. Quality [299, 438, 167, 569, 601, 68, 38]. Quantitative [119, 423, 325, 328, 236, 69]. Queen [657, 621, 544, 524]. Question [124]. Questions [273]. Quick [644]. quinqueradiata [461]. quotas [39].

R [651]. Race [390]. Rachycentron [443]. radiography [12]. Rain [343]. Rainbow [136, 632, 432]. Ranch [401]. Ranching [253, 219, 215, 462, 385, 254]. Ranges [388, 469]. Rastrelliger [646]. Rate [391, 73]. Rates [402, 284]. Rationale [221]. Realizing [686]. Reared [245, 402, 389, 199, 249, 235, 544, 242]. Rearing [677, 257, 182, 75]. Reasons [662]. Reassessment [475]. Rebuttal [175, 174]. Recapture [424, 402, 291]. Receptors [661]. Recommendations [89, 597, 593]. Reconsidering [420]. Reconstruction [426]. Records [476, 311]. Recovering [618]. Recovery [655, 184, 218, 284, 379, 280, 169]. Recreational [202, 197, 466, 577, 573, 398, 574, 411, 437, 536, 283, 192, 537, 531, 558, 658, 540, 24]. Recruit [409]. Recruitment [622, 657, 246, 408, 366, 362]. Recycling [612]. Red [398, 269, 452, 87, 492, 110, 275, 268, 27]. Redefining [35]. Reduce [625, 114]. Reducing [630]. Reef [160, 286]. Reefs [399, 269, 397]. Reevaluation [119]. Reference [365]. Refuting [629]. Regimes [159]. Region [129, 498, 306, 166, 99]. Regional [656, 379]. Regions [551, 503]. Regulated [684]. Regulation [116, 687, 42]. Rehabilitating [397]. Related [318, 395, 658, 536, 44, 34]. Relating [415]. relation [32]. Relationship [529, 431]. Relationships [323, 657, 315, 447, 322, 524]. Relative [102, 526, 404, 628]. Release [197, 222, 391, 234, 241, 402, 239, 558, 236, 291, 251, 229, 658, 387]. Release-Recapture [291]. Released [394, 402, 390]. Releases [395]. Releasing [230, 245, 247, 199, 237]. Remaining [675]. Remote [528, 646]. Removals [370]. Replenishment [217]. Replicated [225]. Report [264]. Reporting [284]. Reports [141]. representative [27]. Reproduction [155, 532, 507, 297, 200, 42, 59]. Reproductive [315, 596, 689, 199, 95, 621]. Requirements [81, 86, 119, 100, 20, 48, 61]. Research [83, 264, 471, 470, 697, 261, 536, 510, 166, 340, 385, 273, 505, 85, 415, 440, 621, 511, 196, 72, 41]. Reserve [402]. Reserves [416]. Reservoir [571, 500, 282]. Reservoirs [156, 413, 450]. Resident [388]. Resilience [499]. Resource [203, 668, 521, 452, 105, 638, 89, 605, 286, 55]. Resources [215, 176, 347, 212, 613, 299, 111, 186, 274, 66]. respect [59]. Response [539, 175, 694, 439, 260, 566, 275, 104]. Responses [214, 132]. Responsibility [679]. Responsible [321, 101, 312, 217]. Restocked [403]. Restocking [222, 253, 215, 394, 400, 218, 226, 385, 254, 221, 227, 244]. Restoration [184, 401, 226, 384]. Restrictions [630]. Retail [616]. Retention [116]. Return [391]. Reversal [265]. Reversed [247]. Review [568, 582, 622, 323, 504, 302, 464, 620, 156, 359, 398, 556, 526, 485, 84, 615, 480, 292, 645, 676, 521, 530, 661, 609, 669, 626, 269, 369, 534, 565, 462, 692, 541, 105, 439, 331, 536, 313, 519, 89, 154, 189, 100, 166, 433, 459, 458, 497, 340, 516, 666, 277, 296, 584, 377, 567, 329, 597, 598, 419, 555, 372, 527, 368, 625, 330, 487, 513, 604, 517, 605, 122, 126, 374, 591, 375, 467, 678, 562, 575, 291, 79, 353, 543, 651, 627, 280, 169, 445, 634, 593, 654, 624, 550, 512]. review [31, 65, 2, 59, 44, 24, 8, 36, 3]. Reviews [152, 165]. Revolution [551]. Rice [592]. Rice-Fish [592]. Rich [590]. Rights [679]. Rio [184, 181, 184, 178, 295, 182, 183, 185]. Riparian [155]. Rise [148]. Risk [289, 171, 581, 296, 276]. Risks [504, 667, 395]. Risso [452]. River [289, 158, 354, 65, 523, 292, 676, 521, 178, 377, 149, 413, 614, 169, 386, 674, 647]. Riverine [282, 688]. Rivers [684, 3]. Roadmap [264, 382]. Robert [651]. Rock [224, 232, 388, 549, 6]. Rockfish [391]. Role [622, 76, 530, 579, 589, 507, 698, 2, 35]. Roles [655, 379]. rosenbergii [20, 86, 148]. rubra [394]. rubripes [239]. Run [413]. Run-of-the-River [413]. Runs [162]. Rural [486]. Russian [274].

S. [461]. Saccostrea [6]. Sacramento [358]. Sacramento-San [358]. Safe [684, 421]. Safety [475]. sagax [332]. Sagay [402]. Saharan [598, 533]. salar [615, 349, 658]. Sales [616]. Salinity [170, 175, 23]. Salmo [615, 349, 658]. salmoides [660]. Salmon [622, 655, 202, 158, 359, 170, 175, 259, 529, 523, 615, 676, 138, 635, 288, 174, 260, 448, 143, 162, 405, 674, 120, 161, 68, 349, 658, 121, 99, 58, 65, 52]. Salmonid [212, 211, 328, 291, 38, 3]. Salmonids [188, 348, 309, 602]. salmonis [170, 175]. Salt [195, 236]. Salton [122, 126]. Samoa [286]. Sampling [189, 362, 47]. San [358, 322, 356]. São [386]. sapidus [234, 217]. Sarcophaga [270]. Sardine [478, 332]. Sardinella [685]. Sardinops [332]. Satellite [646]. Satisfying [429]. saurus [340]. scabra [401]. Scale [559, 690, 18]. Scalefish [337]. Scales [620, 699, 252]. Scallop [366, 7]. Scallops [645]. Scampi [257]. Scat [453]. Scatophagus [453]. Scenarios [579, 563, 604]. Scheme [360]. Schemes [330]. Schistosomiasis [625]. schlegelii [391]. Sciaenidae [691, 493]. Science [648, 361, 333, 384, 186, 493, 518, 627, 165]. Scientific [259]. Scientometric [493, 518]. Scoping [666]. Scotia [166]. Scrutiny [678]. Sculpin [578]. Sea [504, 253, 219, 215, 170, 259, 345, 594, 452, 479, 469, 462, 401, 174, 260, 385, 254, 372, 240, 237, 122, 126, 351, 366, 316, 349, 654, 542, 347, 452, 646, 254, 567, 326, 17]. Seabass [135]. Seabird [630, 342]. Seabream [310, 135]. seafloor [53]. Seafood [603, 475, 616, 667, 643]. Seahorse [470]. seas [4]. Season [241]. Seasonal [47]. Seasonality [642]. Seasonally [523]. Seawater [257]. Seaweed [355, 631, 695]. Seaweeds [580, 644]. Sebastes [391]. Secondary [272]. Sector [430, 680, 599]. Sectoral [446]. Security [686, 533]. Sediment [288]. Sedimentation [105]. Seine [478]. Selected [443, 667]. Selection [390, 69]. Selective [115]. Selectivity [113]. Semen [606]. semi [4]. semi-enclosed [4]. Senegalese [639]. Sensing [528, 646, 473]. Sensitivity [120]. Sentience [662]. Seriola [461]. Serranidae [378]. Services [451, 695]. Set [114, 73]. Set-Net [114]. Seto [240]. Sets [494]. Setting [275]. Settling [116]. Sex [694, 265, 261, 247, 206, 207]. Sex-Reversed [247]. sexfilis [249]. Sexual [556, 261, 665, 42]. Shad [674]. shared [55]. Sharing [638]. Sharks [109]. Sheet [643]. Shellfish [355, 90, 554, 56, 34, 538]. Shifts [602]. Shortnose [421]. Shrimp [81, 86, 480, 661, 107, 452, 463, 448, 515, 287, 128, 588, 88, 140, 562, 595, 79, 196, 8]. Siberia [274]. Significant [364]. Signs [330]. Silver [427]. Silvery [181, 184, 295, 182, 185]. Similarities [526]. Simulation [648, 159]. Since [452]. sinensis [255]. Single [284]. Single-Tag [284]. Site [536]. Sites [494]. Size [193, 278, 241, 162, 113, 237]. Size-Dependent [237]. Skeptical [662]. Small [559, 614, 291, 285]. Small-Scale [559]. Smallmouth [352, 660]. Smart [634]. Smelt [358, 632]. Snake [3]. Snapper [269, 275, 268]. Snook [238]. Snow [492]. Social [197, 466, 531, 505, 600]. Social-Ecological [466]. Social-Ecological-Systems [600]. Socio [637, 286]. Socio-Economic [286]. Socio-Technical [637]. Socioeconomic [97]. Sodium [568]. Solutions [136, 139, 140, 641, 137, 630]. Solved [303]. Some [315, 172, 77, 107]. Sound [359, 570, 58]. Sources [336, 608]. South [394, 455, 412, 582, 646, 190, 606, 341, 412]. Southeast [258, 572]. southeastern [23]. Southern [202, 388, 346, 675, 332, 166, 350]. Southwest [274]. Southwestern [180, 403]. sp [270, 43]. Space [390]. Spain [503]. Spanish [240]. Spanning [414]. Sparidae [403]. Sparus [345]. spathula [200]. Spatial [228, 339, 531, 252, 618, 335, 309, 41]. Spatio [620, 237]. Spatio-Temporal [620, 237]. Spawner [622]. Spawner-Recruitment [622]. Spawning [523, 596, 246, 275, 614]. Special [16, 365, 534]. specialization [72]. Species [430, 172, 352, 77, 691, 580, 380, 86, 603, 498, 266, 551, 693, 322, 356, 319, 234, 423, 340, 226, 296, 632, 379, 372, 270, 258, 685, 665, 578, 493, 518, 413, 618, 412, 351, 309, 421, 621, 608]. Species-Specific [309]. Specific [309, 30]. Specimens [181]. Spectrometry [608]. Spectrometry-Based [608]. Sperm [564]. Spermatozoon [532]. Sphyraena [198, 297]. Spill [359, 120, 58]. Spine [424]. Spiny [144, 682, 235, 28]. Sponges [334]. Sportfish [660]. Spot [529]. Spot-Forward [529]. Spotted [453]. spp [52, 461, 79, 476]. Squid [435, 444, 597]. Stage [170, 175, 28]. Stages [665]. Stained [248]. Stakeholder [220]. Stakeholder-Supported [220]. Standard [30]. Standardization [681]. State [695, 209, 344, 457, 588, 461, 14]. State-of-the-Art [457]. States [350, 362, 427, 454, 141, 495, 633]. Statistical [189]. Statistics [605]. Status [472, 102, 477, 521, 697, 452, 306, 437, 279, 143, 660, 492, 256, 88, 290, 419, 549, 562, 500, 575, 614, 595, 684, 440, 421, 653, 43, 334]. Steelhead [158, 161, 65]. stenolepis [15]. Steps [525]. Stewardship [360]. Stickney [651]. Stimulation [473]. Stock [228, 222, 620, 253, 215, 193, 327, 398, 409, 255, 225, 243, 339, 233, 404, 462, 392, 393, 396, 234, 331, 401, 223, 367, 385, 321, 88, 478, 597, 555, 258, 248, 201, 250, 366, 524, 221, 229, 231, 227, 220, 280, 332, 307, 11, 121, 217]. Stocking [238, 509, 241, 240]. Stocks [394, 365, 501, 246, 218, 166, 382, 537, 350, 544]. Stocks-Long-Term [394]. Stoichiometry [322, 356]. Stomach [112, 336, 550]. Story [623]. Strains [271]. Strategic [264]. Strategies [184, 661, 234, 123, 91, 416]. Strategy [622]. Stream [348, 273, 468, 167, 508, 168]. Stream-Living [348]. Streamflow [178]. Streams [179]. Strength [282]. Strengths [382]. Stress [583, 214, 364]. Stressors [610, 647]. striped [43]. Structure [156, 158, 327, 339, 433, 555, 201, 524, 316, 690]. Structures [116]. Studied [77, 591]. Studies [409, 526, 265, 288, 518, 236, 440, 12, 40, 3]. Study [464, 315, 208, 414, 213, 509, 361, 190, 340, 632, 455, 674, 658, 336, 627, 125, 446]. Studying [609]. Sturgeon [155, 614, 421, 601, 169, 690]. Sturgeons [421]. Style [540]. stylirostris [84]. Sub [400, 598, 533]. Sub-Saharan [598, 533]. Sub-Tidal [400]. Subdivision [335]. Subpopulation [316]. Subsequent [168]. substances [56]. Substitutes [601]. Success [523, 241, 237]. Successful [219, 155, 280]. Suggested [611, 617]. Suitability [251]. summary [12]. Summer [421]. Super [349]. Super-Chicken [349]. Supplementing [254]. Supply [693, 312, 376, 490, 643, 672, 698]. Support [259, 510, 652]. Supported [220]. Supporting [266, 664]. Surface [179, 188]. Survey [189, 274]. Surveys [600, 605, 524]. Survival [170, 175, 394, 392, 288, 291, 65]. Sustainability [680, 128, 588, 420, 478, 330, 180]. Sustainable [504, 612, 547, 486, 692, 563, 510, 382, 287, 346, 101, 664, 500, 575, 496, 621, 572, 611, 617, 650, 324, 106]. Sustained [373]. Swamp [308]. Swim [570]. Swimming [564]. Swimways [696]. Switching [630]. Swordfish [425, 379, 474]. SWOT [673]. Sydney [6]. Synbiotics [541, 583]. Syndrome [566]. Synergistic [610]. Synopsis [77]. Synthesis [197, 427, 365, 385, 578, 575, 524, 621, 653, 508]. System [77, 523, 213, 247]. Systematic [609, 669, 625]. Systems [466, 129, 398, 607, 686, 343, 356, 139, 131, 216, 666, 128, 600, 133].

Tag [424, 284, 239, 608]. Tag-Recapture [424]. Tagging [22]. Takifugu [239]. Targeted [370, 229]. Targeting [645]. Targets [661]. Taste [473]. Taxonomy [419, 476]. Technical [637, 131, 132]. Techniques [219, 255, 329, 328, 397, 314, 75]. Technological [418, 410, 448, 139, 135, 133]. Technologies [266, 542]. Technology [649]. Telemetry [390]. Teleost [302, 272, 689, 145]. Teleostei [263]. Teleosts [564]. Temperature [170, 175, 117, 178, 161, 658]. Temperature-Related [658]. Temporal [620, 348, 237, 309, 41]. Term [394, 676, 322, 392, 524]. Terrestrial [673]. Terrestrial-Animal [673]. Test [121]. Testable [320]. Text [651]. Their [568, 422, 322, 469, 143, 516, 675, 482, 467, 608, 2, 30, 59, 375]. Theories [377]. Theory [677, 632, 350]. Therapy [563]. Thermal [117, 273]. Things [634]. Thirty [219]. Those [248]. Thrace [306]. Threadfin [249]. Threat [352]. Threatened [386, 258]. Threats [696, 613, 641, 500]. Three [536, 270, 517, 286]. Thunnus [545, 556, 383, 424, 441, 431, 447, 199, 326, 555, 201]. thynnus [199, 556, 383, 424, 441, 431, 447, 199, 326, 201]. Thyroid [422]. Tick [520]. Tidal [400, 149]. tides [27]. Tiger [81, 236]. Tilapia [271, 451, 448, 133, 42]. Time [159, 630, 307]. Time-Area [630]. Time-Varying [159, 307].

Timescape [649]. Timing [118, 602]. Today [340]. Tomcod [647]. Tool [609, 224, 296, 542, 39]. Toolbox [205, 384]. Tools [528, 516, 211, 244]. Topics [511]. Tor [476]. Towed [645]. Toxic [375, 7]. Toxicants [318]. Toxicity [120, 375, 29, 59]. Toxins [90]. Trace [589]. Traceability [380, 475]. Trachurus [485]. Trade [281, 380, 546, 601]. Tradeoff [539]. Tradeoffs [451]. Traditional [638]. training [36]. Trait [423, 69]. Traits [315, 596]. Trajectories [564]. Transfer [90, 469, 649]. transferable [39]. Translocating [232]. Translocation [399, 224, 388]. Transmitter [142]. Transnationalism [649]. transpacificus [358]. Trap [337]. Traps [441]. Trawl [572]. Trawls [115]. Treatment [561]. Trematode [628]. Trematode-Caused [628]. Trend [588]. Trends [472, 255, 477, 173, 361, 123, 484, 286, 413, 496, 548, 672, 135, 133]. Triacylglycerol [608]. Tridacnidae [26]. trim [31]. Triploid [93, 371]. Tripneustes [218]. Trochus [400, 399]. Trophic [685, 312, 395, 550, 17]. Trophy [449]. Tropical [407, 636, 419, 549, 160, 500, 450]. Troubled [663]. Trout [202, 158, 136, 161, 432]. Trying [18]. Trypanoplasma [319]. Trypanosoma [319]. Tuna [545, 556, 383, 424, 441, 431, 447, 526, 497, 256, 420, 290, 199, 326, 555, 101, 201, 55]. Tuna-Reconsidering [420]. Tunas [553, 433]. Turbidity [105]. Turbot [251]. Turbulence [415]. Turkey [306, 279]. Turkish [315]. Twenty [106]. Twenty-First [106]. Two [319, 636, 247, 385, 567, 639, 413]. Two-Allele [247].

U.S. [92, 123, 283, 642, 23, 366]. Uganda [582]. Ulcerative [566]. Unbiased [284]. Uncertainties [177]. Uncover [511]. undecimalis [238]. Underlying [665]. Understand [390]. Understanding [197, 466, 696, 216]. Undertaking [686]. Underwater [89]. Unfounded [629]. United [350, 427, 454, 141, 495, 633, 362]. Unsustainability [455]. Unwanted [683]. Update [321, 121]. Updated [340]. Upper [282, 169]. Upstream [523]. Uptake [318]. Urbanization [168]. USA [195, 616, 122, 126, 421]. Usage [337]. Use [102, 156, 203, 668, 271, 278, 425, 521, 673, 389, 416, 257, 79, 421, 285, 227, 332, 362, 25, 32, 28, 56, 11]. Used [243, 517]. Using [409, 424, 225, 213, 340, 277, 296, 505, 236, 183, 511, 390, 476, 21, 75]. Utility [603]. Utilization [317, 264, 345, 298, 286, 151]. Utilizing [314].

Vaccine [491]. Vaccines [477]. Valdez [58, 359, 120]. Valenciennes [310]. Validation [424, 440]. Valley [99]. Value [449, 340, 506]. vannamei [515, 588, 85]. Variability [431, 404, 597, 598, 104]. Variables [539]. Variation [261, 237, 336]. Variations [178]. Various [382]. Varying [159, 307]. vectors [34]. velocity [65]. Venerupis [219]. Very [291]. Vesicles [624]. Vessel [539]. Viability [208]. Vibriosis [480]. Vietnam [378]. Vietnamese [502, 525]. View [149, 51]. Views [660]. Viral [397]. virginica [410]. Virtual [408]. Vis [582]. Vis-à-Vis [582]. Visions [140]. Vitamins [506]. Volume [153]. vs [623]. Vulnerability [417, 313].

Wakasa [387]. Wales [394]. War [88]. Warm [523]. Warming [588, 647]. Washington [14]. Waste [612, 347]. wastes [29]. Water [499, 170, 175, 117, 179, 382, 167, 38, 658, 65]. Water-Groundwater [179]. Waterborne [318]. Waters [489, 663, 478, 419]. Way [500]. Weaknesses [382]. Web [322]. Webs [356, 105]. Weight [102, 431, 447]. Welfare [371]. West [335]. Western [409, 447, 140, 160, 493, 518]. Wetland [575]. Whaling [262]. Where [343, 536]. Whirling [171]. White [155, 515]. whitefish [32]. Whiteleg [588]. Wholesale [190]. Wicked [686]. Wild [592, 528, 394, 386, 399, 403, 246, 174, 249, 235, 248, 237, 236, 242, 37]. Willamette [289]. William [58, 359]. Wind [535]. Winter [243, 59]. Winterkill [21]. Within [360, 350]. Workshop [264]. World [435, 573, 382, 587, 461]. Worldwide [691, 588]. Wr [102]. WSSV [661].

X [12]. X-radiography [12]. Xenobiotic [296]. Xenobiotics [277]. Xiphias [474].

Year [282, 303, 362]. Years [219, 414, 123, 410, 346, 373, 544, 125, 300, 663]. Yellow [254]. Yellowfin [545]. Yellowstripe [297]. Yellowtail [327, 137, 461]. Yield [370, 232]. Young [362]. Young-of-the-Year [362].

Zanzibar [605]. Zebra [659, 62]. Zebrafish [609]. Zone [627].

References

Anonymous:1993:ENa

[1] Anonymous. Editor’s note. Reviews in Fisheries Science, 1(1):1, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

     http://www.tandfonline.com/doi/abs/10.1080/10641269309388532.

Gazzaz:1993:PFT

[2] Sahl S. Gazzaz and Barbara A. Rasco. Parvalbumins in fish and their role as food allergens: a review. Reviews in Fisheries Science, 1(1):1–26, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

     http://www.tandfonline.com/doi/abs/10.1080/10641269309388533.

Thorne:1993:RHS

[3] Richard E. Thorne and Gary E. Johnson. A review of hydroacoustic studies for estimation of salmonid downriver migration past hydroeiectric facilities on the Columbia and Snake Rivers in the 1980s. Reviews in Fisheries Science, 1(1):27–56, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print),

     1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
     10.1080/10641269309388534.

Caddy:1993:TCE

[4] J. F. Caddy. Toward a comparative evaluation of human impacts on fishery ecosystems of enclosed and semi-enclosed seas. Reviews in Fisheries Science, 1(1):57–95, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269309388535.

Anonymous:1993:ENb

[5] Anonymous. Editor’s note. Reviews in Fisheries Science, 1(2):1, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269309388536.

Nell:1993:FSR

[6] John A. Nell. Farming the Sydney rock oyster (Saccostrea commercialis) in Australia. Reviews in Fisheries Science, 1(2):97–120, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269309388537.

Shumway:1993:ITA

[7] Sandra E. Shumway and Allan D. Cembella. The impact of toxic algae on scallop culture and fisheries. Reviews in Fisheries Science, 1(2): 121–150, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269309388538.

Primavera:1993:CRS

[8] J. Honculada Primavera. A critical review of shrimp pond culture in the Philippines. Reviews in Fisheries Science, 1(2):151–201, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269309388539.

Anonymous:1993:E

[9] Anonymous. Errata. Reviews in Fisheries Science, 1(2):202, 1993. CO-DEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269309388540.

Anonymous:1993:ENc

[10] Anonymous. Editor’s note. Reviews in Fisheries Science, 1(3):1, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269309388541.

Willis:1993:SDI

[11] David W. Willis, Brian R. Murphy, and Christopher S. Guy. Stock density indices: Development, use, and limitations. Reviews in Fisheries Science, 1(3):203–222, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269309388542.

Jobling:1993:AXR

[12] Malcolm Jobling, Jørgen S. Christiansen, Even H. Jørgensen, and Arne M. Arnesen. The application of X-radiography in feeding and growth studies with fish: a summary of experiments conducted on Arctic charr. Reviews in Fisheries Science, 1(3):223–237, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641269309388543.

Ronnestad:1993:MAF

[13] Ivar Rønnestad and Hans Jørgen Fyhn. Metabolic aspects of free amino acids in developing marine fish eggs and larvae. Reviews in Fisheries Science, 1(3):239–259, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269309388544.

Halbert:1993:HAA

[14] Cindy L. Halbert. How adaptive is adaptive management? Implementing adaptive management in Washington State and British Columbia. Reviews in Fisheries Science, 1(3):261–283, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641269309388545.

Stickney:1993:CAH

[15] Robert R. Stickney and Han Wu Liu. Culture of Atlantic halibut (Hippoglossus hippoglossus) and Pacific halibut (Hippoglossus stenolepis). Reviews in Fisheries Science, 1(3):285–309, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10641269309388546.

Anonymous:1993:SEN

[16] Anonymous. Special editor’s note. Reviews in Fisheries Science, 1 (4):1, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269309388547.

Schneider:1993:TOM

[17] David C. Schneider and Viacheslav P. Shuntov. The trophic organization of the marine bird community in the Bering Sea. Reviews in Fisheries Science, 1(4):311–335, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269309388548.

Smith:1993:TES

[18] Lynwood S. Smith. Trying to explain scale loss mortality: a continuing puzzle. Reviews in Fisheries Science, 1(4):337–355, 1993. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269309388549.

Anonymous:1994:EN

[19] Anonymous. Editor’s note. Reviews in Fisheries Science, 2(1):1, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269409388550.

DAbramo:1994:NRF

[20] Louis R. D’Abramo and Shyn-Shin Sheen. Nutritional requirements, feed formulation, and feeding practices for intensive culture of the freshwater prawn, Macrobrachium rosenbergii. Reviews in Fisheries Science, 2(1):1–21, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269409388551.

Fast:1994:WPL

[21] Arlo W. Fast. Winterkill prevention in lakes and ponds using artificial aeration. Reviews in Fisheries Science, 2(1):23–77, 1994. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10641269409388552.

Haakonsen:1994:TMA

[22] H. O. Haakonsen and A. O. Anoruo. Tagging and migration of the American lobster Homarus americanus. Reviews in Fisheries Science, 2(1): 79–93, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269409388553.

Peterson:1994:ESF

[23] Mark S. Peterson and Michael R. Meador. Effects of salinity on freshwater fishes in coastal plain drainages in the southeastern U.S. Reviews in Fisheries Science, 2(2):95–121, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641269409388554.

Muoneke:1994:HMR

[24] Maurice I. Muoneke and W. Michael Childress. Hooking mortality: a review for recreational fisheries. Reviews in Fisheries Science, 2(2): 123–156, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269409388555.

Apollonio:1994:UEC

[25] Spencer Apollonio. The use of ecosystem characteristics in fisheries management. Reviews in Fisheries Science, 2(2):157–180, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269409388556.

Lucas:1994:BEM

[26] John S. Lucas. The biology, exploitation, and mariculture of giant clams (Tridacnidae). Reviews in Fisheries Science, 2(3):181–223, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269409388557.

Honjo:1994:BPR

[27] Tsuneo Honjo. The biology and prediction of representative red tides associated with fish kills in Japan. Reviews in Fisheries Science, 2(3): 225–253, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269409388558.

Phillips:1994:DUE

[28] B. F. Phillips and J. D. Booth. Design, use, and effectiveness of collectors for catching the puerulus stage of spiny lobsters. Reviews in Fisheries Science, 2(3):255–289, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269409388559.

Tomasso:1994:TNW

[29] J. R. Tomasso. Toxicity of nitrogenous wastes to aquaculture animals. Reviews in Fisheries Science, 2(4):291–314, 1994. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269409388560.

Hahn:1994:SAP

[30] Kirk O. Hahn, Christopher L. Brown, and Gary D. Pruder. Standard agricultural practices for the culture of specific pathogen-free organisms, and their application to aquaculture. Reviews in Fisheries Science, 2(4):315–330, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269409388561.

Bishop:1994:WTD

[31] James M. Bishop, Glenn Ulrich, and Henrietta S. Wilson. “We are in trim to due it”: a review of Charleston’s mosquito fleet. Reviews in Fisheries Science, 2(4):331–346, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269409388562.

Northcote:1994:MWB

[32] T. G. Northcote and G. L. Ennis. Mountain whitefish biology and habitat use in relation to compensation and improvement possibilities. Reviews in Fisheries Science, 2(4):347–371, 1994. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641269409388563.

Anonymous:1995:EN

[33] Anonymous. Editor’s note. Reviews in Fisheries Science, 3(1):1, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269509388564.

Shumway:1995:PRS

[34] Sandra E. Shumway. Phycotoxin-related shellfish poisoning: Bivalve molluscs are not the only vectors. Reviews in Fisheries Science, 3(1): 1–31, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269509388565.

Momot:1995:RRC

[35] Walter T. Momot. Redefining the role of crayfish in aquatic ecosystems. Reviews in Fisheries Science, 3(1):33–63, 1995. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269509388566.

Sloane:1995:FTI

[36] Michael B. Sloane and Leonard L. Lovshin. Feed training and intensive production of advanced largemouth bass fingeriings: a review. Reviews in Fisheries Science, 3(1):65–89, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641269509388567.

Sheridan:1995:GIH

[37] A. K. Sheridan. The genetic impacts of human activities on wild fish populations. Reviews in Fisheries Science, 3(2):91–108, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269509388568.

Tarazona:1995:WQS

[38] Jose V. Tarazona and Maria J. Muñoz. Water quality in salmonid culture. Reviews in Fisheries Science, 3(2):109–139, 1995. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269509388569.

Squires:1995:ITQ

[39] Dale Squires, James Kirkley, and Clement A. Tisdell. Individual transferable quotas as a fisheries management tool. Reviews in Fisheries Science, 3(2):141–169, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269509388570.

Kotrschal:1995:GAA

[40] Kurt Kotrschal and Harald Essler. Goals and approaches in the analysis of locomotion in fish, with a focus on laboratory studies. Reviews in Fisheries Science, 3(2):171–200, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269509388571.

Langton:1995:STP

[41] Richard W. Langton, Peter J. Auster, and David C. Schneider. A spatial and temporal perspective on research and management of groundfish in the northwest Atlantic. Reviews in Fisheries Science, 3(3):201–229, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269509388572.

Brummett:1995:ERS

[42] Randall E. Brummett. Environmental regulation of sexual maturation and reproduction in tilapia. Reviews in Fisheries Science, 3(3):231–248, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269509388573.

Tamaru:1995:SCM

[43] Clyde S. Tamaru, Fuad Cholik, James Ching-Ming Kuo, and William J. FitzGerald, Jr. Status of the culture of milkfish (Chanos chanos), striped mullet (Mugil cephalus), and grouper (Epinephelus sp.). Reviews in Fisheries Science, 3(3):249–273, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/10641269509388574.

Morado:1995:CPR

[44] J. Frank Morado and Eugene B. Small. Ciliate parasites and related diseases of crustacea: a review. Reviews in Fisheries Science, 3(4): 275–354, 1995. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269509388575.

Anonymous:1996:ENa

[45] Anonymous. Editor’s note. Reviews in Fisheries Science, 4(1):1, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269609388576.

Tucker:1996:ECC

[46] Craig S. Tucker. The ecology of channel catfish culture ponds in Northwest Mississippi. Reviews in Fisheries Science, 4(1):1–55, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269609388577.

Pope:1996:SIF

[47] Kevin L. Pope and David W. Willis. Seasonal influences on freshwater fisheries sampling data. Reviews in Fisheries Science, 4(1):57–73, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269609388578.

Davis:1996:DMR

[48] D. Allen Davis and Delbert M. Gatlin III. Dietary mineral requirements of fish and marine crustaceans. Reviews in Fisheries Science, 4 (1):75–99, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269609388579.

Anonymous:1996:ENb

[49] Anonymous. Editor’s note. Reviews in Fisheries Science, 4(2):1, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269609388580.

Shafland:1996:EFF

[50] Paul L. Shafland. Exotic fishes of Florida — 1994. Reviews in Fisheries Science, 4(2):101–122, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269609388581.

Shafland:1996:EFA

[51] Paul L. Shafland. Exotic fish assessments: an alternative view. Reviews in Fisheries Science, 4(2):123–132, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641269609388582.

Dunn:1996:CHG

[52] J. Richard Dunn. Charles Henry Gilbert (1859–1928): an early fishery biologist and his contributions to knowledge of Pacific salmon (Oncorhynchus spp.). Reviews in Fisheries Science, 4(2):133–184, 1996. CO-DEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269609388583.

Auster:1996:IMF

[53] Peter J. Auster, Richard J. Malatesta, Richard W. Langton, Les Watting, Page C. Valentine, Carol Lee S. Donaldson, Elizabeth W. Langton, Andrew N. Shepard, and War G. Babb. The impacts of mobile fishing gear on seafloor habitats in the Gulf of Maine (Northwest Atlantic): Implications for conservation of fish populations. Reviews in Fisheries Science, 4(2):185–202, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269609388584.

Anonymous:1996:ENc

[54] Anonymous. Editor’s note. Reviews in Fisheries Science, 4(3):1, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269609388585.

Wagner:1996:ABT

[55] Brian Wagner. Atlantic bluefin tuna: International management of a shared resource. Reviews in Fisheries Science, 4(3):203–227, 1996. CO-DEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269609388586.

Raa:1996:UIS

[56] Jan Raa. The use of immunostimulatory substances in fish and shellfish farming. Reviews in Fisheries Science, 4(3):229–288, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269609388587.

Stickney:1996:EN

[57] Robert R. Stickney. Editor’s note. Reviews in Fisheries Science, 4 (4):1–2, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269609388588.

Brannon:1996:EVO

[58] E. L. Brannon and A. W. Maki. The Exxon Valdez oil spill: Analysis of impacts on the Prince William Sound pink salmon. Reviews in Fisheries Science, 4(4):289–337, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269609388589.

Longwell:1996:CAW

[59] A. Crosby Longwell, Sukwoo Chang, and Donald F. Gadbois. A complete analysis of winter flounder (Pleuronectes americanus) egg contaminants with respect to early reproduction, with a review of their toxicity and other environmental concentrations. Reviews in Fisheries Science, 4(4):339–386, 1996. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269609388590.

Stickney:1997:ENa

[60] Robert R. Stickney. Editor’s note. Reviews in Fisheries Science, 5 (1):1–2, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269709388591.

Takeuchi:1997:EFA

[61] Toshio Takeuchi. Essential fatty acid requirements of aquatic animals with emphasis on fish larvae and fingerlings. Reviews in Fisheries Science, 5(1):1–25, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269709388592.

Molloy:1997:NEZ

[62] Daniel P. Molloy, Alexander Y. Karatayev, Lyubov E. Burlakova, Dina P. Kurandina, and Franck Laruelle. Natural enemies of zebra mussels: Predators, parasites, and ecological competitors. Reviews in Fisheries Science, 5(1):27–97, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269709388593.

Stickney:1997:ENb

[63] Robert R. Stickney. Editor’s note. Reviews in Fisheries Science, 5 (2):1, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269709388594.

Kuzmina:1997:MLD

[64] V. V. Kuz’mina and A. G. Gelman. Membrane-linked digestion in fish. Reviews in Fisheries Science, 5(2):99–129, 1997. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269709388595.

Cada:1997:EWV

[65] Glenn F. Cada, Michael D. Deacon, Stephen V. Mitz, and Mark S. Bevelhimer. Effects of water velocity on the survival of downstream-migrating juvenile salmon and steelhead: a review with emphasis on the Columbia River Basin. Reviews in Fisheries Science, 5(2):131–183, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269709388596.

Munro:1997:EMF

[66] J. L. Munro and J. D. Bell. Enhancement of marine fisheries resources. Reviews in Fisheries Science, 5(2):185–222, 1997. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10641269709388597.

Stickney:1997:ENc

[67] Robert R. Stickney. Editor’s note. Reviews in Fisheries Science, 5 (3):1–2, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269709388598.

Sigurgisladottir:1997:SQM

[68] Sjofn Sigurgisladottir, Ole Øtorrissen, Øyvind Lie, Magny Thomassen, and Hannes Hafsteinsson. Salmon quality: Methods to determine the quality parameters. Reviews in Fisheries Science, 5(3):223–252, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269709388599.

Poompuang:1997:TDQ

[69] Supawadee Poompuang and Eric M. Hallerman. Toward detection of quantitative trait loci and marker-assisted selection in fish. Reviews in Fisheries Science, 5(3):253–277, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269709388600.

Scholey:1997:COL

[70] Vernon P. Scholey. Construction and operation of a laboratory or aquaculture facility in a developing country. Reviews in Fisheries Science, 5(3):279–302, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269709388601.

Stickney:1997:ENd

[71] Robert R. Stickney. Editor’s note. Reviews in Fisheries Science, 5 (4):1–2, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269709388602.

Aarset:1997:CSI

[72] Bernt Aarset. From consensus to specialization: Institutional aspects of Norwegian aquaculture research. Reviews in Fisheries Science, 5(4): 303–322, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269709388603.

Blier:1997:DAC

[73] P. U. Blier, D. Pelletier, and J.-D. Dutil. Does aerobic capacity set a limit on fish growth rate? Reviews in Fisheries Science, 5(4):323–340, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269709388604.

Longwell:1997:TCA

[74] A. Crosby Longwell. Toward cloning aquacultured fish. Reviews in Fisheries Science, 5(4):341–365, 1997. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/10641269709388605.

vanderMeeren:1997:DRT

[75] Terje van der Meeren and Kjell Emil Naas. Development of rearing techniques using large enclosed ecosystems in the mass production of marine fish fry. Reviews in Fisheries Science, 5(4):367–390, 1997. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10641269709388606.

Conklin:1998:IPP

[76] Douglas E. Conklin. The IWGCN: Past, present, and future role in the field of crustacean nutrition. Reviews in Fisheries Science, 6(1– 2):1–12, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269891314168.

Ceccaldi:1998:SMP

[77] Hubert J. Ceccaldi. A synopsis of the morphology and physiology of the digestive system of some crustacean species studied in France. Reviews in Fisheries Science, 6(1–2):13–39, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641269891314177.

Jones:1998:CLM

[78] D. A. Jones. Crustacean larval microparticulate diets. Reviews in Fisheries Science, 6(1–2):41–54, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/10641269891314186.

Sorgeloos:1998:UBS

[79] Patrick Sorgeloos, Peter Coutteau, Philippe Dhert, Greet Merchie, and Patrick Lavens. Use of brine shrimp, artemia spp., in larval crustacean nutrition: a review. Reviews in Fisheries Science, 6(1–2):55–68, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269891314195.

Boonyaratpalin:1998:NPM

[80] Mali Boonyaratpalin. Nutrition of Penaeus merguiensis and Penaeus idicus. Reviews in Fisheries Science, 6(1–2):69–78, 1998. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269891314203.

Chen:1998:NRB

[81] Houng-Yung Chen. Nutritional requirements of the black tiger shrimp: Penaeus monodon. Reviews in Fisheries Science, 6(1–2):79–95, 1998. CO-DEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269891314212.

Teshima:1998:NPJ

[82] Shin ichi Teshima. Nutrition of Penaeus japonicus. Reviews in Fisheries Science, 6(1–2):97–111, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269891314221.

Allan:1998:RNR

[83] Geoff L. Allan and David M. Smith. Recent nutrition research with Australian penaeids. Reviews in Fisheries Science, 6(1–2):113–127, 1998. CO-DEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269891314230.

Cuzon:1998:NRP

[84] Gérard Cuzon and Aquacop. Nutritional review of Penaeus stylirostris. Reviews in Fisheries Science, 6(1–2):129–141, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10641269891314249.

Pedrazzoli:1998:RAN

[85] Andrés Pedrazzoli, César Molina, Nelson Montoya, Soraya Townsend, Aurora León-Hing, Yela Parades, and Jorge Calderón. Recent advances on nutrition research of Penaeus vannamei in Ecuador. Reviews in Fisheries Science, 6(1–2):143–151, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269891314258.

DAbramo:1998:NRF

[86] Louis R. D’Abramo. Nutritional requirements of the freshwater prawn Macrobrachium rosenbergii: Comparisons with species of penaeid shrimp. Reviews in Fisheries Science, 6(1–2):153–163, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10641269891314267.

Loher:1998:MBB

[87] Timothy Loher, P. Scott Hill, Gretchen Harrington, and Edward Cassano. Management of Bristol Bay red king crab: a critical intersections approach to fisheries management. Reviews in Fisheries Science, 6(3):169–251, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269891314285.

Mohammed:1998:KPG

[88] Hussain M. A. Mohammed, James M. Bishop, and Yimin Ye. Kuwait’s post Gulf-war shrimp fishery and stock status from 1991/92 through 1995/96. Reviews in Fisheries Science, 6(3):253–280, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/10641269891314276.

Keevin:1998:RNR

[89] Thomas M. Keevin. A review of natural resource agency recommendations for mitigating the impacts of underwater blasting. Reviews in Fisheries Science, 6(4):281–313, 1998. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269891314302.

Bricelj:1998:PST

[90] V. Monica Bricelj and Sandra E. Shumway. Paralytic shellfish toxins in bivalve molluscs: Occurrence, transfer kinetics, and biotransformation. Reviews in Fisheries Science, 6(4):315–383, 1998. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269891314294.

Kubitza:1999:FDF

[91] Fernando Kubitza and Leonard L. Lovshin. Formulated diets, feeding strategies, and cannibalism control during intensive culture of juvenile carnivorous fishes. Reviews in Fisheries Science, 7(1):1–22, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641269991319171.

Gerber:1999:MMN

[92] L. R. Gerber, W. S. Wooster, D. P. DeMaster, and G. R. VanBlaricom. Marine mammals: New objectives in U.S. fishery management. Reviews in Fisheries Science, 7(1):23–38, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641269991319180.

Benfey:1999:PBT

[93] Tillmann J. Benfey. The physiology and behavior of triploid fishes. Reviews in Fisheries Science, 7(1):39–67, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10641269991319162.

Shafland:1999:IBPa

[94] Paul L. Shafland. The introduced butterfly peacock (Cichla ocellaris) in Florida. i. fish community analyses. Reviews in Fisheries Science, 7 (2):71–94, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269991319199.

Shafland:1999:IBPb

[95] Paul L. Shafland. The introduced butterfly peacock (Cichla ocellaris) in Florida. II. food and reproductive biology. Reviews in Fisheries Science, 7(2):95–113, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269991319207.

Shafland:1999:IBPc

[96] Paul L. Shafland. The introduced butterfly peacock (Cichla ocellaris) in Florida. III. length distribution analyses. Reviews in Fisheries Science, 7(2):115–126, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641269991319216.

Shafland:1999:IBPd

[97] Paul L. Shafland and Murray S. Stanford. The introduced butterfly peacock (Cichla ocellaris) in Florida. IV. socioeconomic analyses. Reviews in Fisheries Science, 7(2):127–135, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641269991319225.

Argue:1999:HFI

[98] Brad J. Argue and Rex A. Dunham. Hybrid fertility, introgression, and backcrossing in fish. Reviews in Fisheries Science, 7(3–4):137–195, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269908951360.

Yoshiyama:1999:HSP

[99] Ronald M. Yoshiyama. A history of salmon and people in the Central Valley Region of California. Reviews in Fisheries Science, 7(3–4): 197–239, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641269908951361.

Knauer:1999:RNR

[100] Jens Knauer and Paul C. Southgate. A review of the nutritional requirements of bivalves and the development of alternative and artificial diets for bivalve aquaculture. Reviews in Fisheries Science, 7(3–4):241–280, 1999. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269908951362.

Nishida:1999:TSR

[101] Tom Nishida and David Fluharty. Toward sustainable and responsible tuna fisheries. Reviews in Fisheries Science, 7(3–4):281–302, 1999. CO-DEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641269908951363.

Blackwell:2000:RWW

[102] Brian G. Blackwell, Michael L. Brown, and David W. Willis. Relative weight (Wr) status and current use in fisheries assessment and management. Reviews in Fisheries Science, 8(1):1–44, 2000. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641260091129161.

Tucker:2000:FPA

[103] Craig S. Tucker. Off-flavor problems in aquaculture. Reviews in Fisheries Science, 8(1):45–88, 2000. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641260091129170.

Shima:2000:RPP

[104] Michiyo Shima, Anne Babcock Hollowed, and Glenn R. VanBlaricom. Response of pinniped populations to directed harvest, climate variability, and commercial fishery activity: a comparative analysis. Reviews in Fisheries Science, 8(2):89–124, 2000. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641260091129189.

Henley:2000:EST

[105] W. F. Henley, M. A. Patterson, R. J. Neves, and A. Dennis Lemly. Effects of sedimentation and turbidity on lotic food webs: a concise review for natural resource managers. Reviews in Fisheries Science, 8(2):125–139, 2000. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641260091129198.

Wurts:2000:SAT

[106] William A. Wurts. Sustainable aquaculture in the twenty-first century. Reviews in Fisheries Science, 8(2):141–150, 2000. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/10641260091129206.

Fast:2000:SRI

[107] Arlo W. Fast and Piamsak Menasveta. Some recent issues and innovations in marine shrimp pond culture. Reviews in Fisheries Science, 8(3):151–233, 2000. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641260091129215.

Hewitt:2000:FAK

[108] Roger P. Hewitt and Elizabeth H. Linen Low. The fishery on Antarctic krill: Defining an ecosystem approach to management. Reviews in Fisheries Science, 8(3):235–298, 2000. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641260091129224.

Cortes:2000:LHP

[109] Enric Cortés. Life history patterns and correlations in sharks. Reviews in Fisheries Science, 8(4):299–344, 2000. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10408340308951115.

Oliveira:2000:MAP

[110] Eurico C. Oliveira, Krisler Alveal, and Robert J. Anderson. Mariculture of the agar-producing gracilarioid red algae. Reviews in Fisheries Science, 8(4):345–377, 2000. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10408340308951116.

Pullin:2000:MAB

[111] Roger S. V. Pullin. Management of aquatic biodiversity and genetic resources. Reviews in Fisheries Science, 8(4):379–393, 2000. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10408340308951117.

Kamler:2001:NME

[112] Jan F. Kamler and Kevin L. Pope. Nonlethal methods of examining fish stomach contents. Reviews in Fisheries Science, 9(1):1–11, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/20016491101663.

Nishiuchi:2001:SSH

[113] Shuichi Nishiuchi. Size selectivity of hair crab pots. Reviews in Fisheries Science, 9(1):13–26, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/20016491101672.

Ueno:2001:ARS

[114] Yoichiro Ueno. An approach to reduce set-net discard. Reviews in Fisheries Science, 9(1):27–31, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/20016491101681.

Matsushita:2001:CCB

[115] Yoshiki Matsushita and Masahiro Shida. Comparison of catches between selective and conventional otter trawls for a coastal fishery. Reviews in Fisheries Science, 9(1):33–42, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/20016491101690.

Boyd:2001:FRS

[116] Claude E. Boyd and Julio F. Queiroz. Feasibility of retention structures, settling basins, and best management practices in effluent regulation for Alabama channel catfish farming. Reviews in Fisheries Science, 9(2):43–67, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/20016491101708.

Cooke:2001:AEC

[117] S. J. Cooke and J. F. Schreer. Additive effects of chlorinated biocides and water temperature on fish in thermal effluents with emphasis on the great lakes. Reviews in Fisheries Science, 9(2):69–113, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/20016491101717.

Spieler:2001:CTM

[118] Richard E. Spieler. Circadian timing of meal feeding and growth in fishes. Reviews in Fisheries Science, 9(3):115–131, 2001. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/20016491101726.

Hauler:2001:RQD

[119] Rhys C. Hauler and Chris G. Carter. Reevaluation of the quantitative dietary Lysine requirements of fish. Reviews in Fisheries Science, 9(3):133–163, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/20016491101735.

Rice:2001:IPS

[120] Stanley D. Rice, Robert E. Thomas, Mark G. Carls, Ronald A. Heintz, Alex C. Wertheimer, Michael L. Murphy, Jeffrey W. Short, and Adam Moles. Impacts to pink salmon following the Exxon Valdez oil spill: Persistence, toxicity, sensitivity, and controversy. Reviews in Fisheries Science, 9(3):165–211, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/20016491101744.

Winans:2001:UGS

[121] Gary A. Winans, Dan Viele, Allen Grover, Melodie Palmer-Zwahlen, David Teel, and Donald Van Doornik. An update of genetic stock identification of Chinook salmon in the Pacific Northwest: Test fisheries in California. Reviews in Fisheries Science, 9(4):213–237, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/20016491101753.

Riedel:2001:RFS

[122] Ralf Riedel and Barry A. Costa-Pierce. Review of the fisheries of the Salton Sea, California, USA: Past, present, and future. Reviews in Fisheries Science, 9(4):239–270, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/20016491101762.

Jolly:2001:PFT

[123] Curtis Jolly, Carel Ligeon, Jerry Crews, Zach Morley, and Rex Dunham. Present and future trends in the U.S. catfish industry: Strategies and concerns for Millennium years. Reviews in Fisheries Science, 9(4):271–295, 2001. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/20016491101771.

Rose:2002:NNF

[124] James D. Rose. The neurobehavioral nature of fishes and the question of awareness and pain. Reviews in Fisheries Science, 10(1):1–38, 2002.

CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/20026491051668.

Walls:2002:HCL

[125] Elizabeth A. Walls, Jim Berkson, and Stephen A. Smith. The horseshoe crab, Limulus polyphemus: 200 million years of existence, 100 years of study. Reviews in Fisheries Science, 10(1):39–73, 2002. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/20026491051677.

Riedel:2002:RFS

[126] Ralf Riedel and Barry A. Costa-Pierce. Review of the fisheries of the Salton Sea, California, USA: Past, present, and future. Reviews in Fisheries Science, 10(1):77–112, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/20026491051686.

Landsberg:2002:EHA

[127] Jan H. Landsberg. The effects of harmful algal blooms on aquatic organisms. Reviews in Fisheries Science, 10(2):113–390, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/20026491051695.

Menasveta:2002:ISG

[128] Piamsak Menasveta. Improved shrimp growout systems for disease prevention and environmental sustainability in Asia. Reviews in Fisheries Science, 10(3–4):391–402, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/20026491051703.

Brown:2002:FFS

[129] Nick Brown. Flatfish farming systems in the Atlantic region. Reviews in Fisheries Science, 10(3–4):403–419, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/20026491051712.

Seikai:2002:FCC

[130] Tadahisa Seikai. Flounder culture and its challenges in Asia. Reviews in Fisheries Science, 10(3–4):421–432, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/20026491051721.

Liao:2002:TIE

[131] I. Chiu Liao, Ya-Ke Hsu, and Wu Chung Lee. Technical innovations in eel culture systems. Reviews in Fisheries Science, 10(3–4):433–450, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/20026491051730.

Su:2002:TRC

[132] Mao-Sen Su, Cheng-Sheng Lee, and I. Chiu Liao. Technical responses to challenges in milkfish aquaculture. Reviews in Fisheries Science, 10(3–4):451–464, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/20026491051749.

Watanabe:2002:TPS

[133] Wade O. Watanabe, Thomas M. Losordo, Kevin Fitzsimmons, and Fred Hanley. Tilapia production systems in the Americas: Technological advances, trends, and challenges. Reviews in Fisheries Science, 10(3–4): 465–498, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
20026491051758.

Hargreaves:2002:CCF

[134] John A. Hargreaves. Channel catfish farming in ponds: Lessons from a maturing industry. Reviews in Fisheries Science, 10(3–4):499–528, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/20026491051767.

Theodorou:2002:CFT

[135] John A. Theodorou. Current and future technological trends of European seabass–seabream culture. Reviews in Fisheries Science, 10(3–4): 529–543, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
20026491051776.

Fornshell:2002:RTC

[136] Gary Fornshell. Rainbow trout — challenges and solutions. Reviews in Fisheries Science, 10(3–4):545–557, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/20026491051785.

Nakada:2002:YCD

[137] Makoto Nakada. Yellowtail culture development and solutions for the future. Reviews in Fisheries Science, 10(3–4):559–575, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/20026491051794.

Forster:2002:FSE

[138] John Forster. Farming salmon: an example of aquaculture for the mass market. Reviews in Fisheries Science, 10(3–4):577–591, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/20026491051802.

Lee:2002:GDA

[139] Cheng-Sheng Lee. General discussion on “Aquaculture Growout Systems — Challenges and Technological Solutions”. Reviews in Fisheries Science, 10(3–4):593–600, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/20026491051811.

Moss:2002:MSF

[140] Shaun M. Moss. Marine shrimp farming in the Western Hemisphere: Past problems, present solutions, and future visions. Reviews in Fisheries Science, 10(3–4):601–620, 2002. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/20026491051820.

Mitchell:2003:RFK

[141] Andrew J. Mitchell. Reports of fish kills prior to 1900 in the United States. Reviews in Fisheries Science, 11(1):1–11, 2003. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/16226510390856501.

Bridger:2003:EBT

[142] C. J. Bridger and R. K. Booth. The effects of biotelemetry transmitter presence and attachment procedures on fish physiology and behavior. Reviews in Fisheries Science, 11(1):13–34, 2003. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/16226510390856510.

Lackey:2003:PNS

[143] Robert T. Lackey. Pacific Northwest salmon: Forecasting their status in 2100. Reviews in Fisheries Science, 11(1):35–88, 2003. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/16226510390856529.

Cox:2003:FBS

[144] Serena L. Cox and Danielle J. Johnston. Feeding biology of spiny lobster larvae and implications for culture. Reviews in Fisheries Science, 11(2):89–106, 2003. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/713610924.

Tocher:2003:MFL

[145] Douglas R. Tocher. Metabolism and functions of lipids and fatty acids in teleost fish. Reviews in Fisheries Science, 11(2):107–184, 2003. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/713610925.

Cannon:2003:CDM

[146] Quinn Cannon and Eric Wagner. Comparison of discharge mechanisms of cnidarian cnidae and myxozoan polar capsules. Reviews in Fisheries Science, 11(3):185–219, 2003. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10641260390244305.

Oviatt:2003:CFF

[147] Candace Oviatt, Steven Olsen, Mark Andrews, Jeremy Collie, Timothy Lynch, and Kenneth Raposa. A century of fishing and fish fluctuations in Narragansett Bay. Reviews in Fisheries Science, 11(3):221–242, 2003. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/10641260390244413.

Fast:2003:RFF

[148] Arlo W. Fast and PingSun Leung. Rise and fall of freshwater prawn (Macrobrachium rosenbergii) culture in Hawaii: 1965–1999. Reviews in Fisheries Science, 11(3):243–290, 2003. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641260390244431.

Peterson:2003:CVE

[149] Mark S. Peterson. A conceptual view of environment-habitat-production linkages in tidal river estuaries. Reviews in Fisheries Science, 11(4): 291–313, 2003. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641260390255844.

Latremouille:2003:FEA

[150] David N. Latremouille. Fin erosion in aquaculture and natural environments. Reviews in Fisheries Science, 11(4):315–335, 2003. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/10641260390255745.

Stone:2003:DCU

[151] David A. J. Stone. Dietary carbohydrate utilization by fish. Reviews in Fisheries Science, 11(4):337–369, 2003. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/10641260390260884.

Stickney:2003:BR

[152] Robert R. Stickney. Book reviews. Reviews in Fisheries Science, 11(4): 371–374, 2003. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641260390255583.

Anonymous:2003:VAI

[153] Anonymous. Volume 11 author index. Reviews in Fisheries Science, 11 (4):375, 2003. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
10641260390267067.

Khidir:2004:RGC

[154] K. Teresa Khidir, Claude B. Renaud, and François Chapleau. Review of the genus Cyclopsetta (Pleuronectiformes: Paralichthyidae). Reviews in Fisheries Science, 12(1):1–21, 2004. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260490271188.

Coutant:2004:RHH

[155] Charles C. Coutant. A riparian habitat hypothesis for successful reproduction of white sturgeon. Reviews in Fisheries Science, 12(1): 23–73, 2004. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260490273023.

Bolding:2004:UAS

[156] Bruce Bolding, Scott Bonar, and Marc Divens. Use of artificial structure to enhance angler benefits in lakes, ponds, and reservoirs: a literature review. Reviews in Fisheries Science, 12(1):75–96, 2004. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260490273050.

Stickney:2004:RL

[157] Robert R. Stickney. Recent literature. Reviews in Fisheries Science, 12 (1):97–98, 2004. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260490278424.

Brannon:2004:PSC

[158] Ernest L. Brannon, Madison S. Powell, Thomas P. Quinn, and AndrÉ Talbot. Population structure of Columbia River Basin Chinook salmon and steelhead trout. Reviews in Fisheries Science, 12(2–3): 99–232, 2004. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260490280313.

Neill:2004:EFS

[159] William H. Neill, T. Scott Brandes, Brian J. Burke, Steven R. Craig, Leonard V. Dimichele, Kerri Duchon, Randy E. Edwards, Lance P. Fontaine, Delbert M. Gatlin III, Cynthia Hutchins, John M. Miller, Bonnie J. Ponwith, Christopher J. Stahl, Joseph R. Tomasso, and Robert R. Vega. Ecophys.Fish: a simulation model of fish growth in time-varying environmental regimes. Reviews in Fisheries Science, 12(4): 233–288, 2004. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260490479818.

Panek:2005:EDC

[160] Frank M. Panek. Epizootics and disease of coral reef fish in the tropical Western Atlantic and Gulf of Mexico. Reviews in Fisheries Science, 13 (1):1–21, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260590885852.

Richter:2005:MTL

[161] Ann Richter and Steven A. Kolmes. Maximum temperature limits for Chinook, Coho, and chum salmon, and steelhead trout in the Pacific Northwest. Reviews in Fisheries Science, 13(1):23–49, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260590885861.

Meengs:2005:ESH

[162] Chad C. Meengs and Robert T. Lackey. Estimating the size of historical Oregon salmon runs. Reviews in Fisheries Science, 13(1):51– 66, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260590921509.

Stickney:2005:NL

[163] Robert R. Stickney. New literature. Reviews in Fisheries Science, 13 (1):67–69, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260590921527.

Anonymous:2005:E

[164] Anonymous. Erratum. Reviews in Fisheries Science, 13(1):71, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/10641260590921536.

Stickney:2005:RFS

[165] Robert R. Stickney. Reviews in Fisheries Science additional information for authors. Reviews in Fisheries Science, 13(2):73–74, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260590953883.

Kock:2005:FSS

[166] Karl-Hermann Kock and Christopher D. Jones. Fish stocks in the Southern Scotia Arc region — a review and prospects for future research. Reviews in Fisheries Science, 13(2):75–108, 2005. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260590953900.

Silapajarn:2005:ECC

[167] Orawan Silapajarn and Claude E. Boyd. Effects of channel catfish farming on water quality and flow in an Alabama stream. Reviews in Fisheries Science, 13(2):109–140, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260590953928.

Wheeler:2005:INH

[168] Andrew P. Wheeler, Paul L. Angermeier, and Amanda E. Rosenberger. Impacts of new highways and subsequent landscape urbanization on stream habitat and biota. Reviews in Fisheries Science, 13(3):141– 164, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260590964449.

Webb:2005:RPR

[169] Molly A. H. Webb, Jack E. Williams, and Larry R. Hildebrand. Recovery program review for endangered pallid sturgeon in the Upper Missouri River Basin. Reviews in Fisheries Science, 13(3):165–176, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260590964458.

Brooks:2005:EWT

[170] Kenneth M. Brooks. The effects of water temperature, salinity, and currents on the survival and distribution of the infective copepodid stage of sea lice (Lepeophtheirus salmonis) originating on Atlantic salmon farms in the Broughton Archipelago of British Columbia, Canada. Reviews in Fisheries Science, 13(3):177–204, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260500207109.

Bartholomew:2005:RAB

[171] Jerri L. Bartholomew, Billie L. Kerans, Ronald P. Hedrick, Stuart C. Macdiarmid, and James R. Winton. A risk assessment based approach for the management of whirling disease. Reviews in Fisheries Science, 13(4): 205–230, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260500326842.

Boyd:2005:CIS

[172] Claude E. Boyd, Aaron A. McNevin, Jason Clay, and Howard M. Johnson. Certification issues for some common aquaculture species. Reviews in Fisheries Science, 13(4):231–279, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260500326867.

Harikrishnan:2005:MTA

[173] R. Harikrishnan and C. Balasundaram. Modern trends in Aeromonas hydrophila disease management with fish. Reviews in Fisheries Science, 13 (4):281–320, 2005. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260500320845.

Krkosek:2006:FFS

[174] Martin Krkosek, Mark A. Lewis, John P. Volpe, and Alexandra Morton. Fish farms and sea lice infestations of wild juvenile salmon in the Broughton Archipelago — a rebuttal to Brooks (2005). Reviews in Fisheries Science, 14(1–2):1–11, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/10641260500433531.

Brooks:2006:EWT

[175] Kenneth M. Brooks and Dario J. Stucchi. The effects of water temperature, salinity and currents on the survival and distribution of the infective copepodid stage of the salmon louse (lepeophtheirus salmonis) originating on Atlantic salmon farms in the Broughton Archipelago of British Columbia, Canada (Brooks, 2005) — a response to the rebuttal of Krkošek et al. (2005a). Reviews in Fisheries Science, 14(1–2):13–23, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260500448893.

Cowley:2006:PCM

[176] David E. Cowley and Rossana Sallenave. Preface: Conservation and management of aquatic resources in arid lands. Reviews in Fisheries Science, 14(1–2):25–27, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260500340561.

Lioubimtseva:2006:UCC

[177] Elena Lioubimtseva and Roy Cole. Uncertainties of climate change in arid environments of central Asia. Reviews in Fisheries Science, 14(1– 2):29–49, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260500340603.

Hall:2006:RVT

[178] Alexander W. Hall, Paul H. Whitfield, and Alex J. Cannon. Recent variations in temperature, precipitation, and streamflow in the Rio Grande and Pecos River Basins of New Mexico and Colorado. Reviews in Fisheries Science, 14(1–2):51–78, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260500340835.

Fernald:2006:SWG

[179] Alexander G. Fernald and Steven J. Guldan. Surface water-groundwater interactions between irrigation ditches, alluvial aquifers, and streams. Reviews in Fisheries Science, 14(1–2):79–89, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260500341320.

Rinne:2006:HGM

[180] John N. Rinne and Dennis Miller. Hydrology, geomorphology and management: Implications for sustainability of native southwestern fishes. Reviews in Fisheries Science, 14(1–2):91–110, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260500341379.

Cowley:2006:ERG

[181] David E. Cowley, Patrick D. Shirey, and Michael D. Hatch. Ecology of the Rio Grande silvery minnow (Cyprinidae: Hybognathus amarus) inferred from specimens collected in 1874. Reviews in Fisheries Science, 14(1–2): 111–125, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260500341494.

Osborne:2006:GEH

[182] Megan J. Osborne, Melissa A. Benavides, Dominique Alò, and Thomas F. Turner. Genetic effects of hatchery propagation and rearing in the endangered Rio Grande silvery minnow, hybognathus amarus. Reviews in Fisheries Science, 14(1–2):127–138, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260500341544.

Smith:2006:IUN

[183] Laura L. Smith and Alexander G. Fernald. The ineffectiveness of using the National Environmental Policy Act of 1969 to implement environmental enhancement in the Rio Grande Canalization Project. Reviews in Fisheries Science, 14(1–2):139–167, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260500341577.

Cowley:2006:SER

[184] David E. Cowley. Strategies for ecological restoration of the Middle Rio Grande in New Mexico and recovery of the endangered Rio Grande silvery minnow. Reviews in Fisheries Science, 14(1–2):169–186, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260500341619.

Ward:2006:EII

[185] Frank A. Ward and James F. Booker. Economic impacts of instream flow protection for the Rio grande silvery minnow in the Rio grande basin. Reviews in Fisheries Science, 14(1–2):187–202, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260500341734.

Sallenave:2006:SEP

[186] Rossana Sallenave and David E. Cowley. Science and effective policy for managing aquatic resources. Reviews in Fisheries Science, 14(1–2): 203–210, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260500341783.

Stickney:2006:NLa

[187] Robert R. Stickney. New literature. Reviews in Fisheries Science, 14(1– 2):211–212, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260500385368.

Johnson:2006:SFO

[188] Gary E. Johnson and Dennis D. Dauble. Surface flow outlets to protect juvenile salmonids passing through hydropower dams. Reviews in Fisheries Science, 14(3):213–244, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260600864023.

Kimura:2006:RSA

[189] Daniel K. Kimura and David A. Somerton. Review of statistical aspects of survey sampling for marine fisheries. Reviews in Fisheries Science, 14(3): 245–283, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260600621761.

Kim:2006:DBA

[190] Jin Baek Kim. Determinants of B2B adoption in South Korean fishery wholesale firms: an exploratory study. Reviews in Fisheries Science, 14(3):285–301, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260600816510.

Stickney:2006:NLb

[191] Robert R. Stickney. New literature. Reviews in Fisheries Science, 14 (3):303, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260600580173.

Lewin:2006:DPB

[192] Wolf-Christian Lewin, Robert Arlinghaus, and Thomas Mehner. Documented and potential biological impacts of recreational fishing: Insights for management and conservation. Reviews in Fisheries Science, 14(4): 305–367, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260600886455.

Bordalo-Machado:2006:FEA

[193] Pedro Bordalo-Machado. Fishing effort analysis and its potential to evaluate stock size. Reviews in Fisheries Science, 14(4):369–393, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260600893766.

Stickney:2006:RL

[194] Robert R. Stickney. Recent literature. Reviews in Fisheries Science, 14 (4):395–398, 2006. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260600886497.

Able:2007:LBS

[195] Kenneth W. Able, John H. Balletto, Stacy M. Hagan, Paul R. Jivoff, and Kenneth Strait. Linkages between salt marshes and other nekton habitats in Delaware Bay, USA. Reviews in Fisheries Science, 15(1– 2):1–61, 2007. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260600960995.

Vincent:2007:ARN

[196] Amanda G. Vincent and Jeffrey M. Lotz. Advances in research of necrotizing hepatopancreatitis bacterium (NHPB) affecting penaeid shrimp aquaculture. Reviews in Fisheries Science, 15(1–2):63–73, 2007. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260601079902.

Arlinghaus:2007:UCC

[197] Robert Arlinghaus, Steven J. Cooke, Jon Lyman, David Policansky, Alexander Schwab, Cory Suski, Stephen G. Sutton, and Eva B. Thorstad. Understanding the complexity of catch-and-release in recreational fishing: an integrative synthesis of global knowledge from historical, ethical, social, and biological perspectives. Reviews in Fisheries Science, 15(1– 2):75–167, 2007. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260601149432.

Rim:2007:AGL

[198] Zouari-Ktari Rim, Ennejjar Samira, Bradai Mohamed-Nejmeddine, Ghorbel Mohamed, and Bouain Abderrahmen. Age and growth of the Lessepsian migrant sphyraena chrysotaenia Klunzinger, 1884 from the Gulf of Gabes (Eastern Mediterranean). Reviews in Fisheries Science, 15(3): 169–181, 2007. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701484259.

Mylonas:2007:PAG

[199] Constantinos C. Mylonas, Christopher Bridges, Hillel Gordin, Antonio Belmonte Ríos, Antonio García, Fernando De La Gándara, Christian Fauvel, Marc Suquet, Antonio Medina, Maria Papadaki, Gilad Heinisch, Gregorio De Metrio, Aldo Corriero, Robert Vassallo-Agius, Jose-María Guzmán, Evaristo Mañanos, and Yonathan Zohar. Preparation and administration of gonadotropin-releasing hormone agonist (GnRHa) implants for the artificial control of reproductive maturation in captivereared Atlantic bluefin tuna (Thunnus thynnus thynnus). Reviews in Fisheries Science, 15(3):183–210, 2007. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260701484572.

Scarnecchia:2007:LHC

[200] Dennis L. Scarnecchia, L. Fred Ryckman, Youngtaik Lim, Greg J. Power, Brad J. Schmitz, and Jon A. Firehammer. Life history and the costs of reproduction in northern great plains paddlefish (polyodon spathula) as a potential framework for other acipenseriform fishes. Reviews in Fisheries Science, 15(3):211–263, 2007. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260701486981.

Rooker:2007:LHS

[201] Jay R. Rooker, Jaime R. Alvarado Bremer, Barbara A. Block, Heidi Dewar, Gregorio de Metrio, Aldo Corriero, Richard T. Kraus, Eric D. Prince, Enrique Rodríguez-Marín, and David H. Secor. Life history and stock structure of Atlantic bluefin tuna (Thunnus thynnus). Reviews in Fisheries Science, 15(4):265–310, 2007. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260701484135.

Arismendi:2007:NNS

[202] I. Arismendi and L. Nahuelhual. Non-native salmon and trout recreational fishing in Lake Llanquihue, Southern Chile: Economic benefits and management implications. Reviews in Fisheries Science, 15(4): 311–325, 2007. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701484655.

Boyd:2007:IRU

[203] Claude E. Boyd, Craig Tucker, Aaron Mcnevin, Katherine Bostick, and Jason Clay. Indicators of resource use efficiency and environmental performance in fish and crustacean aquaculture. Reviews in Fisheries Science, 15(4):327–360, 2007. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260701624177.

Sundell:2008:I

[204] Profs. Kristina Sundell and Profs. Deborah Power. Introduction. Reviews in Fisheries Science, 16:1–2, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260802497010.

Canario:2008:GTF

[205] Adelino V. M. Canario, Luca Bargelloni, Filip Volckaert, R. D. Houston, C. Massault, and Yann Guiguen. Genomics toolbox for farmed fish. Reviews in Fisheries Science, 16:3–15, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260802319479.

Penman:2008:FGP

[206] David J. Penman and Francesc Piferrer. Fish gonadogenesis. part i: Genetic and environmental mechanisms of sex determination. Reviews in Fisheries Science, 16:16–34, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/10641260802324610.

Piferrer:2008:FGP

[207] Francesc Piferrer and Yann Guiguen. Fish gonadogenesis. part II: Molecular biology and genomics of sex differentiation. Reviews in Fisheries Science, 16:35–55, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260802324644.

Cerda:2008:FGP

[208] Joan Cerdà, Julien Bobe, Patrick J. Babin, Arie Admon, and Esther Lubzens. Functional genomics and proteomic approaches for the study of gamete formation and viability in farmed finfish. Reviews in Fisheries Science, 16:56–72, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260802324685.

Leaver:2008:TFL

[209] Michael J. Leaver, José M. Bautista, Björn Thrandur Björnsson, Elisabeth Jönsson, Grigorios Krey, Douglas R. Tocher, and Bente E. Torstensen. Towards fish lipid nutrigenomics: Current state and prospects for finfish aquaculture. Reviews in Fisheries Science, 16:73–94, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802325278.

Power:2008:MEB

[210] Deborah M. Power, Ingibjörg E. Einarsdóttir, Karin Pittman, Glen E. Sweeney, Jon Hildahl, Marco A. Campinho, Nadia Silva, Øystein Sæle, Malyka Galay-Burgos, Heiddís Smáradóttir, and Björn Thrandur Björnsson. The molecular and endocrine basis of flatfish metamorphosis. Reviews in Fisheries Science, 16:95–111, 2008. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802325377.

Martin:2008:GTE

[211] S. A. M. Martin, B. Collet, S. Mackenzie, O. Evensen, and C. J. Secombes. Genomic tools for examining immune gene function in salmonid fish. Reviews in Fisheries Science, 16:112–118, 2008. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802325476.

Dios:2008:GRI

[212] S. Dios, B. Novoa, F. Buonocore, G. Scapigliati, and A. Figueras. Genomic resources for immunology and disease of salmonid and non-salmonid fish. Reviews in Fisheries Science, 16:119–132, 2008. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802325484.

Gestal:2008:SDI

[213] Camino Gestal, Philippe Roch, Tristan Renault, Alberto Pallavicini, Christine Paillard, Beatriz Novoa, Radouane Oubella, Paola Venier, and Antonio Figueras. Study of diseases and the immune system of bivalves using molecular biology and genomics. Reviews in Fisheries Science, 16: 133–156, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802325518.

Prunet:2008:FGS

[214] Patrick Prunet, Michael T. Cairns, Svante Winberg, and Thomas G. Pottinger. Functional genomics of stress responses in fish. Reviews in Fisheries Science, 16:157–166, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260802341838.

Bell:2008:NER

[215] Johann D. Bell, Kenneth M. Leber, H. Lee Blankenship, Neil R. Loneragan, and Reiji Masuda. A new era for restocking, stock enhancement and sea ranching of coastal fisheries resources. Reviews in Fisheries Science, 16(1–3):1–9, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260701776951.

Lorenzen:2008:UME

[216] Kai Lorenzen. Understanding and managing enhancement fisheries systems. Reviews in Fisheries Science, 16(1–3):10–23, 2008. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260701790291.

Zohar:2008:CBB

[217] Yonathan Zohar, Anson H. Hines, Oded Zmora, Eric G. Johnson, Romuald N. Lipcius, Rochelle D. Seitz, David B. Eggleston, Allen R. Place, Eric J. Schott, John D. Stubblefield, and J. Sook Chung. The Chesapeake Bay blue crab (Callinectes sapidus): a multidisciplinary approach to responsible stock replenishment. Reviews in Fisheries Science, 16(1– 3):24–34, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701681623.

Juinio-Menez:2008:ERD

[218] Marie Antonette Juinio-Meñez, Helen Grage Bangi, Maria Celia Malay, and Davelyn Pastor. Enhancing the recovery of depleted Tripneustes gratilla stocks through grow-out culture and restocking. Reviews in Fisheries Science, 16(1–3):35–43, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/10641260701678116.

Becker:2008:TYS

[219] Peter Becker, Carl Barringer, and Dan C. Marelli. Thirty years of sea ranching Manila clams (Venerupis philippinarum): Successful techniques and lessons learned. Reviews in Fisheries Science, 16(1–3):

44–50, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701790259.

Tringali:2008:MSE

[220] Michael D. Tringali, Kenneth M. Leber, William G. Halstead, Robert Mcmichael, Joseph O’hop, Brent Winner, Richard Cody, Chris Young, Carole Neidig, Heather Wolfe, Ann Forstchen, and Luiz Barbieri. Marine stock enhancement in Florida: a multi-disciplinary, stakeholdersupported, accountability-based approach. Reviews in Fisheries Science, 16(1–3):51–57, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-

6553 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/10641260701776902.

Stottrup:2008:RRE

[221] J. G. Støttrup, J. L. Overton, H. Paulsen, C. Möllmann, J. Tomkiewicz, P. B. Pedersen, and P. Lauesen. Rationale for restocking the Eastern Baltic cod stock. Reviews in Fisheries Science, 16(1–3):58– 64, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701678231.

Arnold:2008:ALR

[222] William S. Arnold. Application of larval release for restocking and stock enhancement of coastal marine bivalve populations. Reviews in Fisheries Science, 16(1–3):65–71, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260701678140.

LeVay:2008:ASE

[223] Lewis Le Vay, Ma Junemie Hazel Lebata, Mark Walton, Jurgenne Primavera, Emilia Quinitio, Celia Lavilla-Pitogo, Fe Parado-Estepa, Eduard Rodriguez, Vu Ngoc Ut, Truong Trong Nghia, Patrick Sorgeloos, and Mathieu Wille. Approaches to stock enhancement in mangrove-associated crab fisheries. Reviews in Fisheries Science, 16(1–3):72–80, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260701727285.

Gardner:2008:BMT

[224] Caleb Gardner and E. Ingrid Van Putten. Biological modeling of translocation as a management tool for a rock lobster fishery. Reviews in Fisheries Science, 16(1–3):81–90, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260701696183.

Eggleston:2008:PEE

[225] David B. Eggleston, Eric G. Johnson, G. Todd Kellison, Gayle R. Plaia, and Cynthia L. Huggett. Pilot evaluation of early juvenile blue crab stock enhancement using a replicated BACI design. Reviews in Fisheries Science, 16(1–3):91–100, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260701681664.

Lipcius:2008:IMC

[226] Romuald N. Lipcius, David B. Eggleston, Sebastian J. Schreiber, Rochelle D. Seitz, Jian Shen, Mac Sisson, William T. Stockhausen, and Harry V. Wang. Importance of metapopulation connectivity to restocking and restoration of marine species. Reviews in Fisheries Science, 16(1–3): 101–110, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701812574.

Tringali:2008:LUC

[227] Michael D. Tringali, Seifu Seyoum, Elizabeth M. Wallace, Maryanne Higham, Ronald G. Taylor, Alexis A. Trotter, and James A. Whittington. Limits to the use of contemporary genetic analyses in delineating biological populations for restocking and stock enhancement. Reviews in Fisheries Science, 16(1–3):111–116, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260701686887.

Aguilar:2008:IBC

[228] Robert Aguilar, Eric G. Johnson, Anson H. Hines, Margaret A. Kramer, and Michael R. Goodison. Importance of blue crab life history for stock enhancement and spatial management of the fishery in Chesapeake Bay. Reviews in Fisheries Science, 16(1–3):117–124, 2008. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260701681599.

Taylor:2008:PIM

[229] Matthew D. Taylor and Iain M. Suthers. A predatory impact model and targeted stock enhancement approach for optimal release of mulloway (Argyrosomus japonicus). Reviews in Fisheries Science, 16(1–3): 125–134, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701727293.

Arnason:2008:ERC

[230] Ragnar Arnason. On the economics of releasing cultured fish into the aquatic environment. Reviews in Fisheries Science, 16(1–3):135– 145, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701689030.

Tomiyama:2008:CBS

[231] Takeshi Tomiyama, Masato Watanabe, and Tsuneo Fujita. Communitybased stock enhancement and fisheries management of the Japanese flounder in Fukushima, Japan. Reviews in Fisheries Science, 16(1–3): 146–153, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701681847.

Gardner:2008:EFT

[232] Caleb Gardner and E. Ingrid Van Putten. The economic feasibility of translocating rock lobsters to increase yield. Reviews in Fisheries Science, 16(1–3):154–163, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260701681789.

Hamasaki:2008:PSE

[233] Katsuyuki Hamasaki and Shuichi Kitada. Potential of stock enhancement for decapod crustaceans. Reviews in Fisheries Science, 16(1–3): 164–174, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701681722.

Hines:2008:RSE

[234] Anson H. Hines, Eric G. Johnson, Alicia C. Young, Robert Aguilar, Margaret A. Kramer, Michael Goodison, Oded Zmora, and Yonathan Zohar. Release strategies for estuarine species with complex migratory life cycles: Stock enhancement of Chesapeake blue crabs (Callinectes sapidus). Reviews in Fisheries Science, 16(1–3):175–185, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260701678090.

Oliver:2008:APB

[235] Megan D. Oliver, Alison B. Macdiarmid, Robert A. Stewart, and Caleb Gardner. Anti-predator behavior of captive-reared and wild juvenile spiny lobster (Jasus edwardsii). Reviews in Fisheries Science, 16(1–3): 186–194, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701727301.

Shimizu:2008:QEP

[236] Daisuke Shimizu, Kazutaka Sakiyama, Yoshitaka Sakakura, Tomohiro Takatani, and Yoh-Ichi Takahashi. Quantitative evaluation of postrelease mortality using salt pond mesocosms: Case studies of hatchery and wild juvenile tiger puffer. Reviews in Fisheries Science, 16(1–3): 195–203, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701681755.

Purcell:2008:STS

[237] Steven W. Purcell and Matéo Simutoga. Spatio-temporal and sizedependent variation in the success of releasing cultured sea cucumbers in the wild. Reviews in Fisheries Science, 16(1–3):204–214, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260701686895.

Brennan:2008:MSM

[238] Nathan P. Brennan, Carl J. Walters, and Kenneth M. Leber. Manipulations of stocking magnitude: Addressing density-dependence in a juvenile cohort of common snook (Centropomus undecimalis). Reviews in Fisheries Science, 16(1–3):215–227, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260701689022.

Nakajima:2008:ORL

[239] Hiroshi Nakajima, Masanobu Kai, Koji Koizumi, Toshiomi Tanaka, and Masaharu Machida. Optimal release locations of juvenile ocellate puffer Takifugu rubripes identified by tag and release experiments. Reviews in Fisheries Science, 16(1–3):228–234, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260701684106.

Obata:2008:ESE

[240] Yasuhiro Obata, Hideki Yamazaki, Akio Iwamoto, Katsuyuki Hamasaki, and Shuichi Kitada. Evaluation of stocking effectiveness of the Japanese Spanish mackerel in the Eastern Seto Inland Sea, Japan. Reviews in Fisheries Science, 16(1–3):235–242, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260701697041.

Johnson:2008:ISS

[241] Eric G. Johnson, Anson H. Hines, Margaret A. Kramer, and Alicia C. Young. Importance of season and size of release to stocking success for the blue crab in Chesapeake Bay. Reviews in Fisheries Science, 16(1–3): 243–253, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701696837.

Young:2008:DHR

[242] Alicia C. Young, Eric G. Johnson, Jana L. D. Davis, Anson H. Hines, Oded Zmora, and Yonathan Zohar. Do hatchery-reared blue crabs differ from wild crabs, and does it matter? Reviews in Fisheries Science, 16(1–3): 254–261, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701684122.

Fairchild:2008:PAA

[243] Elizabeth A. Fairchild, Nathan Rennels, and W. Huntting Howell. Predators are attracted to acclimation cages used for winter flounder stock enhancement. Reviews in Fisheries Science, 16(1–3):262–268, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260701678181.

deInnocentiis:2008:MTM

[244] Sabina de Innocentiis, Alessandro Longobardi, and Giovanna Marino. Molecular tools in a marine restocking program for the endangered dusky grouper, Epinephelus marginatus. Reviews in Fisheries Science, 16(1–3): 269–277, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701678348.

Hara:2008:AIR

[245] Moyoyiki Hara, Shizumasa Onoue, and Nobuhiko Taniguchi. Assessing the impact of releasing exogenous hatchery-reared juveniles of Pacific abalone, Haliotis discus. Reviews in Fisheries Science, 16(1–3): 278–284, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701684114.

Jorstad:2008:EEF

[246] Knut E. Jørstad, Terje van der Meeren, Ole Ingar Paulsen, Tårn Thomsen, Anders Thorsen, and Terje Svåsand. “Escapes” of eggs from farmed cod spawning in net pens: Recruitment to wild stocks. Reviews in Fisheries Science, 16(1–3):285–295, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260701678017.

Kanaiwa:2008:COL

[247] Minoru Kanaiwa and Yasushi Harada. Collapse of one-locus two-allele sex determining system by releasing sex-reversed hatchery fish. Reviews in Fisheries Science, 16(1–3):296–302, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260701678371.

Potter:2008:CGG

[248] Ian C. Potter, Daniel J. W. French, Greg I. Jenkins, S. Alex Hesp, Norman G. Hall, and Simon de Lestang. Comparisons of the growth and gonadal development of otolith-stained, cultured black bream, Acanthopagrus butcheri, in an estuary with those of its wild stock. Reviews in Fisheries Science, 16(1–3):303–316, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260701681565.

Ogawa:2008:FHW

[249] Thomas K. Ogawa, David A. Ziemann, and Scott W. S. Bloom. Feeding habits of wild and hatchery-reared Pacific threadfin, Polydactylus sexfilis (family Polynemidae), in Hawaií. Reviews in Fisheries Science, 16(1–3): 317–328, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701678397.

Seitz:2008:SEC

[250] Rochelle D. Seitz, Romuald N. Lipcius, Kathleen E. Knick, Michael S. Seebo, W. Christopher Long, Bryce J. Brylawski, and Alison Smith. Stock enhancement and carrying capacity of blue crab nursery habitats in Chesapeake Bay. Reviews in Fisheries Science, 16(1–3):329–337, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260701696985.

Sparrevohn:2008:DAD

[251] Claus R. Sparrevohn and Josianne G. Støttrup. Diet, abundance, and distribution as indices of turbot (Psetta maxima L.) release habitat suitability. Reviews in Fisheries Science, 16(1–3):338–347, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260701686846.

Pedersen:2008:SSM

[252] Torstein Pedersen, Els L. Cuveliers, and Erik Berg. Spatial scales of movement in Northeast Atlantic coastal cod. Reviews in Fisheries Science, 16(1–3):348–356, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260701678462.

Bartley:2008:RSE

[253] Devin M. Bartley and Johann D. Bell. Restocking, stock enhancement, and sea ranching: Arenas of progress. Reviews in Fisheries Science, 16(1– 3):357–365, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701678058.

Lu:2008:SMC

[254] Huaqing Lü, Junzhuo Xu, and Geraldine Vander Haegen. Supplementing marine capture fisheries in the East China Sea: Sea ranching of prawn Penaeus orientalis, restocking of large yellow croaker Pseudosciaena crocea, and cage culture. Reviews in Fisheries Science, 16(1–3): 366–376, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701678207.

Cheng:2008:CTH

[255] Yongxu Cheng, Xugan Wu, Xiaozhen Yang, and Anson H. Hines. Current trends in hatchery techniques and stock enhancement for Chinese mitten crab, Eriocheir japonica sinensis. Reviews in Fisheries Science, 16(1–3): 377–384, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701681698.

Masuma:2008:SBT

[256] Shukei Masuma, Shigeru Miyashita, Hiroshi Yamamoto, and Hidemi Kumai. Status of bluefin tuna farming, broodstock management, breeding and fingerling production in Japan. Reviews in Fisheries Science, 16(1– 3):385–390, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701484325.

Okamoto:2008:UDS

[257] Kazutoshi Okamoto. Use of deep seawater for rearing Japanese scampi lobster (Metanephrops japonica) broodstock. Reviews in Fisheries Science, 16(1–3):391–393, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260701678504.

Okuzawa:2008:SET

[258] Koichi Okuzawa, Ronald J. Maliao, Emilia T. Quinitio, Shelah Mae A. Buen-Ursua, Ma. Junemie Hazel L. Lebata, Wenresti G. Gallardo, Luis M. B. Garcia, and Jurgenne H. Primavera. Stock enhancement of threatened species in Southeast Asia. Reviews in Fisheries Science, 16(1–3): 394–402, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260701678496.

Brooks:2008:PPS

[259] Kenneth M. Brooks and Simon R. M. Jones. Perspectives on pink salmon and sea lice: Scientific evidence fails to support the extinction hypothesis. Reviews in Fisheries Science, 16(4):403–412, 2008. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260801937131.

Krkosek:2008:SLP

[260] Martin Krkosek, Jennifer S. Ford, Alexandra Morton, Subhash Lele, and Mark A. Lewis. Sea lice and pink salmon declines: a response to Brooks and Jones (2008). Reviews in Fisheries Science, 16(4):413– 420, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802013692.

Hanson:2008:SVF

[261] K. C. Hanson, M. A. Gravel, A. Graham, A. Shoji, and S. J. Cooke. Sexual variation in fisheries research and management: When does sex matter? Reviews in Fisheries Science, 16(4):421–436, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802013866.

Tillman:2008:IMA

[262] Michael F. Tillman. The international management of Aboriginal whaling. Reviews in Fisheries Science, 16(4):437–444, 2008. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802013932.

Oktener:2008:PFN

[263] Ahmet Öktener. Peniculus fistula von Nordmann, 1832 (Copepoda: Pennelidae) parasitic on Coryphaena hippurus Linnaeus, 1758 (Teleostei: Coryphaenidae). Reviews in Fisheries Science, 16(4):445– 448, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802046668.

Barrows:2008:RPP

[264] Frederic T. Barrows, Diane Bellis, Åshild Krogdahl, Jeffrey T. Silverstein, Eliot M. Herman, Wendy M. Sealey, Michael B. Rust, and Delbert M. Gatlin III. Report of the plant products in aquafeed strategic planning workshop: an integrated, interdisciplinary research roadmap for increasing utilization of plant feedstuffs in diets for carnivorous fish. Reviews in Fisheries Science, 16(4):449–455, 2008. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260802046734.

Cuevas-Uribe:2009:PSH

[265] Rafael Cuevas-Uribe, Boris Gomelsky, Steven D. Mims, and Kirk W. Pomper. Progress in studies on hormonal sex reversal and genetic sex control in black crappie. Reviews in Fisheries Science, 17(1):1–7, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802003073.

Eugenio-Gonzalez:2009:ITS

[266] Marco A. Eugenio-Gonzalez, Gerardo Padilla-Zarate, Carlos Montes De Oca, and Carmen G. Paniagua-Chavez. Information technologies supporting the operation of the germplasm bank of aquatic species of Baja California, Mexico. Reviews in Fisheries Science, 17(1):8–17, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802008031.

Kurtovic:2009:LMF

[267] Ivan Kurtovic, Susan N. Marshall, Xin Zhao, and Benjamin K. Simpson. Lipases from mammals and fishes. Reviews in Fisheries Science, 17 (1):18–40, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802031322.

Shipp:2009:PIA

[268] Robert L. Shipp and Stephen A. Bortone. A perspective of the importance of artificial habitat on the management of red snapper in the Gulf of Mexico. Reviews in Fisheries Science, 17(1):41–47, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802104244.

Gallaway:2009:LHR

[269] Benny J. Gallaway, Stephen T. Szedlmayer, and William J. Gazey. A life history review for red snapper in the Gulf of Mexico with an evaluation of the importance of offshore petroleum platforms and other artificial reefs. Reviews in Fisheries Science, 17(1):48–67, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802160717.

Oktener:2009:CEM

[270] Ahmet Öktener and Ali Alas. Cases of external myiasis caused by Diptera (Sarcophaga sp.) on three cyprinid species. Reviews in Fisheries Science, 17(1):68–69, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802178818.

Brummett:2009:CAE

[271] Randall E. Brummett and Raul W. Ponzoni. Concepts, alternatives, and environmental considerations in the development and use of improved strains of tilapia in African aquaculture. Reviews in Fisheries Science, 17(1):70–77, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802083174.

Johnson:2009:LDO

[272] Ronald B. Johnson. Lipid deposition in oocytes of teleost fish during secondary oocyte growth. Reviews in Fisheries Science, 17(1): 78–89, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802590004.

McCullough:2009:RTB

[273] Dale A. McCullough, John M. Bartholow, Henriëtte I. Jager, Robert L. Beschta, Edward F. Cheslak, Michael L. Deas, Joseph L. Ebersole, J. Scott Foott, Sherri L. Johnson, Keith R. Marine, Matthew G. Mesa, James H. Petersen, Yves Souchon, Kenneth F. Tiffan, and Wayne A. Wurtsbaugh. Research in thermal biology: Burning questions for coldwater stream fishes. Reviews in Fisheries Science, 17(1):90–115, 2009. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802590152.

VanStappen:2009:SAR

[274] Gilbert Van Stappen, Lyudmila Ilinichna Litvinenko, Aleksandr Ivanovich Litvinenko, Elena Grigorevno Boyko, Brad Marden, and Patrick Sorgeloos. A survey of Artemia resources of Southwest Siberia (Russian Federation). Reviews in Fisheries Science, 17(1):116–148, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802590095.

Phelps:2009:SRS

[275] R. P. Phelps, N. Papanikos, B. D. Bourque, F. T. Bueno, R. P. Hastey, D. L. Maus, A. Ferry, and D. A. Davis. Spawning of red snapper (Lutjanus campechanus) in response to hormonal induction or environmental control in a hatchery setting. Reviews in Fisheries Science, 17(2): 149–155, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802505689.

Zajicek:2009:FMO

[276] Paul Zajicek, Scott Hardin, and Craig Watson. A Florida marine ornamental pathway risk analysis. Reviews in Fisheries Science, 17(2): 156–169, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802536577.

MachadoDaRocha:2009:EGE

[277] Carlos Alberto Machado Da Rocha, Patrícia Danielle Lima De Lima, Raquel Alves Dos Santos, and Rommel Mario Rodríguez Burbano. Evaluation of genotoxic effects of xenobiotics in fishes using comet assay — a review. Reviews in Fisheries Science, 17(2):170–173, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802541791.

Clemens:2009:EBS

[278] Benjamin J. Clemens and Stephen S. Crawford. The ecology of body size and depth use by bloater (Coregonus hoyi gill) in the Laurentian Great Lakes: Patterns and hypotheses. Reviews in Fisheries Science, 17(2): 174–186, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802618375.

Harlioglu:2009:SFC

[279] Ayse Gül Harlioglu and Muzaffer Mustafa Harlioglu. The status of freshwater crayfish (Astacus leptodactylus Eschscholtz) fisheries in Turkey. Reviews in Fisheries Science, 17(2):187–189, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260802645311.

Wakeford:2009:RIA

[280] Robert C. Wakeford, David J. Agnew, and Christopher C. Mees. Review of institutional arrangements and evaluation of factors associated with successful stock recovery plans. Reviews in Fisheries Science, 17(2): 190–222, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802667075.

Bell:2009:CCP

[281] Johann D. Bell, Eric Clua, Cathy A. Hair, Rene Galzin, and Peter J. Doherty. The capture and culture of post-larval fish and invertebrates for the marine ornamental trade. Reviews in Fisheries Science, 17(2): 223–240, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802528541.

Scarnecchia:2009:RRI

[282] Dennis L. Scarnecchia, L. Fred Ryckman, Youngtaik Lim, Shannon E. Miller, Brad Schmitz, Greg J. Power, and Shane Shefstad. Riverine and reservoir influences on year class strength and growth of upper great plains paddlefish. Reviews in Fisheries Science, 17(2):241–266, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802621791.

Kerr:2009:CCU

[283] L. A. Kerr, W. J. Connelly, E. J. Martino, A. C. Peer, R. J. Woodland, and D. H. Secor. Climate change in the U.S. Atlantic affecting recreational fisheries. Reviews in Fisheries Science, 17(2):267–289, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260802667067.

McGarvey:2009:MEM

[284] Richard McGarvey. Methods of estimating mortality and movement rates from single-tag recovery data that are unbiased by tag non-reporting. Reviews in Fisheries Science, 17(3):291–304, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260802664841.

Tacon:2009:FAN

[285] Albert G. J. Tacon and Marc Metian. Fishing for aquaculture: Nonfood use of small pelagic forage fish — a global perspective. Reviews in Fisheries Science, 17(3):305–317, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260802677074.

Sabater:2009:TRF

[286] Marlowe G. Sabater and Benjamin P. Carroll. Trends in reef fish population and associated fishery after three millennia of resource utilization and a century of socio-economic changes in American Samoa. Reviews in Fisheries Science, 17(3):318–335, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260802702872.

Manoj:2009:FOS

[287] Valsa Remony Manoj and Namasivayam Vasudevan. Functional options for sustainable shrimp aquaculture in India. Reviews in Fisheries Science, 17(3):336–347, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260802715072.

Jensen:2009:IFS

[288] David W. Jensen, E. Ashley Steel, Aimee H. Fullerton, and George R. Pess. Impact of fine sediment on egg-to-fry survival of Pacific salmon: a meta-analysis of published studies. Reviews in Fisheries Science, 17(3): 348–359, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260902716954.

Arsan:2009:PDN

[289] E. Leyla Arsan and Jerri L. Bartholomew. Potential dispersal of the nonnative parasite Myxobolus cerebralis in the Willamette River Basin, Oregon: a qualitative analysis of risk. Reviews in Fisheries Science, 17(3): 360–372, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802013726.

Mourente:2009:TNF

[290] G. Mourente and D. R. Tocher. Tuna nutrition and feeds: Current status and future perspectives. Reviews in Fisheries Science, 17(3): 373–390, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260902752207.

Skalski:2009:RMM

[291] John R. Skalski, Rebecca A. Buchanan, and Jim Griswold. Review of marking methods and release-recapture designs for estimating the survival of very small fish: Examples from the assessment of salmonid fry survival. Reviews in Fisheries Science, 17(3):391–401, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260902752199.

Davis:2009:RPE

[292] Windy N. Davis, Robert G. Bramblett, Alexander V. Zale, and Carol L. Endicott. A review of the potential effects of coal bed natural gas development activities on fish assemblages of the Powder River Geologic Basin. Reviews in Fisheries Science, 17(3):402–422, 2009. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260902737067.

Anonymous:2009:LI

[293] Anonymous. Literature of interest. Reviews in Fisheries Science, 17 (3):423, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260802690010.

Grizzle:2009:IDF

[294] John M. Grizzle and Cindy J. Brunner. Infectious diseases of freshwater mussels and other freshwater bivalve mollusks. Reviews in Fisheries Science, 17(4):425–467, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260902879000.

Magana:2009:FPR

[295] Hugo A. Magana. Feeding preference of the Rio grande silvery minnow (Hybognathus amarus). Reviews in Fisheries Science, 17(4):468– 477, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260902985096.

MachadoDaRocha:2009:MAF

[296] Carlos Alberto Machado Da Rocha, Raquel Alves Dos Santos, Marcelo De Oliveira Bahia, Lorena Araújo Da Cunha, Helem Ferreira Ribeiro, and Rommel Mario Rodríguez Burbano. The micronucleus assay in fish species as an important tool for xenobiotic exposure risk assessment — a brief review and an example using neotropical fish exposed to methylmercury. Reviews in Fisheries Science, 17(4):478–484, 2009. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260903067852.

Rim:2009:RGY

[297] Zouari-Ktari Rim, Bradai Mohamed-Nejmeddine, and Bouain Abderrahmen. Reproduction and growth of the yellowstripe barracuda Sphyraena chrysotaenia Klunzinger, 1884, in central Mediterranean. Reviews in Fisheries Science, 17(4):485–493, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260903082471.

Peng:2009:PGB

[298] Chen Peng, Shao Hong-Bo, Xu Di, and Qin Song. Progress in Gracilaria biology and developmental utilization: Main issues and prospective. Reviews in Fisheries Science, 17(4):494–504, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641260903144586.

Peterson:2009:ICI

[299] Mark S. Peterson and Michael R. Lowe. Implications of cumulative impacts to estuarine and marine habitat quality for fish and invertebrate resources. Reviews in Fisheries Science, 17(4):505–523, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260903171803.

Weber:2009:ECC

[300] Michael J. Weber and Michael L. Brown. Effects of common carp on aquatic ecosystems 80 years after “Carp as a dominant”: Ecological insights for fisheries management. Reviews in Fisheries Science, 17(4): 524–537, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260903189243.

Bunkley-Williams:2009:HDP

[301] Lucy Bunkley-Williams and Ernest H. Williams, Jr. Hosts and distribution of Peniculus fistula von nordman, 1832. Reviews in Fisheries Science, 17(4):538–540, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260903216012.

Arellano-Aguilar:2009:EFE

[302] Omar Arellano-Aguilar, Regina Montero Montoya, and Constantino Macías Garcia. Endogenous functions and expression of cytochrome P450 enzymes in teleost fish: a review. Reviews in Fisheries Science, 17(4): 541–556, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260903243487.

Williams:2009:YOD

[303] Ernest H. Williams, Jr. and Lucy Bunkley-Williams. A 226-year-old dolphin-isopod mystery solved: Identification of barbugede pampelfisk, Coryphaena apus brünnich, 1783. Reviews in Fisheries Science, 17(4): 557–561, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260903243495.

Stickney:2009:RL

[304] Robert R. Stickney. Recent literature. Reviews in Fisheries Science, 17 (4):562–563, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260903106650.

Anonymous:2009:C

[305] Anonymous. Corrigendum. Reviews in Fisheries Science, 17(4): 564, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260903344509.

Guner:2009:SFC

[306] Utku Güner and Muzaffer Mustafa Harlioglu. Status of freshwater crayfish distribution in thrace region of Turkey. Reviews in Fisheries Science, 18(1):1–6, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260903267445.

Wilberg:2009:ITV

[307] Michael J. Wilberg, James T. Thorson, Brian C. Linton, and Jim Berkson. Incorporating time-varying catchability into population dynamic stock assessment models. Reviews in Fisheries Science, 18(1):7–24, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260903294647.

Shafland:2009:AAS

[308] Paul L. Shafland, Kelly B. Gestring, and Murray S. Stanford. An assessment of the Asian swamp eel (Monopterus albus) in Florida. Reviews in Fisheries Science, 18(1):25–39, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260903225542.

Smith:2009:SSS

[309] David L. Smith, John M. Nestler, Gary E. Johnson, and R. Andrew Goodwin. Species-specific spatial and temporal distribution patterns of emigrating juvenile salmonids in the Pacific Northwest. Reviews in Fisheries Science, 18(1):40–64, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260903304487.

Hamida:2009:FHB

[310] Nader Ben Hadj Hamida, Olfa Ben Hadj Hamida-Ben Abdallah, Mohamed Ghorbel, Othman Jarboui, and Hechmi Missaoui. The feeding habits of the bluespotted seabream, Pagrus caeruleostictus (Valenciennes, 1830), in the Gulf of Gabes (Central Mediterranean). Reviews in Fisheries Science, 18(1):65–72, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641260903318230.

Williams:2009:CPD

[311] Ernest H. Williams, Jr. and Lucy Bunkley-Williams. Checklists of the parasites of dolphin, Coryphaena hippurus, and pompano dolphin, C. equiselis with new records, corrections, and comments on the literature. Reviews in Fisheries Science, 18(1):73–93, 2009. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260903295909.

Tacon:2009:RAT

[312] Albert G. J. Tacon, Marc Metian, Giovanni M. Turchini, and Sena S. De Silva. Responsible aquaculture and trophic level implications to global fish supply. Reviews in Fisheries Science, 18(1):94–105, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260903325680.

Johnson:2009:MFM

[313] Johanna E. Johnson and David J. Welch. Marine fisheries management in a changing climate: a review of vulnerability and future options. Reviews in Fisheries Science, 18(1):106–124, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641260903434557.

Zheng:2009:DUM

[314] Jia-Yi Zheng, Wen Zhuang, Yue Tao Yi, Gang Wu, Jun Gong, and Hong-Bo Shao. Developmentally utilizing molecular biological techniques into aquaculture. Reviews in Fisheries Science, 18(1):125–130, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260903477499.

Berber:2009:SRB

[315] Selçuk Berber, Harun Yildiz, A. Suat Ates, Musa Bulut, and Mehmet Mendes. A study on the relationships between some morphological and reproductive traits of the Turkish crayfish, Astacus leptodactylus Eschscholtz, 1823 (Crustacea: Decapoda). Reviews in Fisheries Science, 18 (1):131–137, 2009. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260903491003.

Svedang:2010:CSS

[316] H. Svedäng, J. Stål, T. Sterner, and M. Cardinale. Consequences of subpopulation structure on fisheries management: Cod (Gadus morhua) in the Kattegat and Öresund (North Sea). Reviews in Fisheries Science, 18 (2):139–150, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260903511420.

Altinagac:2010:UBP

[317] Ugur Altinagaç, Ugur Özekinci, Adnan Ayaz, Ozcan Ozen, and Alkan öztekin. Utilization of a bioluminescent pattern in the encircling gillnet fisheries. Reviews in Fisheries Science, 18(2):151–156, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260903545295.

Korai:2010:HLG

[318] A. L. Korai, K. H. Lashari, G. A. Sahato, and T. G. Kazi. Histological lesions in gills of feral cyprinids, related to the uptake of waterborne toxicants from Keenjhar Lake. Reviews in Fisheries Science, 18 (2):157–176, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641260903430522.

Gupta:2010:HCB

[319] Dileep K. Gupta and Neelima Gupta. A hemogram of Clarias batrachus parasitized by two species of hemoflagellates (Trypanosoma batrachi qadri, 1962 and Trypanoplasma haematalis n.sp.) as an indicator of infectivity. Reviews in Fisheries Science, 18(2):177–182, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641260903450132.

Apollonio:2010:THE

[320] Spencer Apollonio. A testable hypothesis for ecosystem-based fisheries management. Reviews in Fisheries Science, 18(2):183–188, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2010.490885.

Lorenzen:2010:RAM

[321] Kai Lorenzen, Kenneth M. Leber, and H. Lee Blankenship. Responsible approach to marine stock enhancement: an update. Reviews in Fisheries Science, 18(2):189–210, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2010.491564.

Glibert:2010:LTC

[322] Patricia M. Glibert. Long-term changes in nutrient loading and stoichiometry and their relationships with changes in the food Web and dominant pelagic fish species in the San Francisco Estuary, California. Reviews in Fisheries Science, 18(2):211–232, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2010.492059.

Anonymous:2010:RGO

[323] Anonymous. A review of “Gonorynchiformes and ostariophysan relationships. a comprehensive review”. Reviews in Fisheries Science, 18(2): 233–234, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2010.490884.

Welch:2010:FSF

[324] A. Welch, R. Hoenig, J. Stieglitz, D. Benetti, A. Tacon, N. Sims, and B. O’Hanlon. From fishing to the sustainable farming of carnivorous marine finfish. Reviews in Fisheries Science, 18(3):235–247, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2010.504865.

Metcalf:2010:QMC

[325] Sarah J. Metcalf. Qualitative models to complement quantitative ecosystem models for the analysis of data-limited marine ecosystems and fisheries. Reviews in Fisheries Science, 18(3):248–265, 2010. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2010.505670.

Mylonas:2010:ABT

[326] Constantinos C. Mylonas, Fernando De La Gándara, Aldo Corriero, and Antonio Belmonte Ríos. Atlantic bluefin tuna (Thunnus thynnus) farming and fattening in the Mediterranean Sea. Reviews in Fisheries Science, 18(3):266–280, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2010.509520.

Cadrin:2010:IAY

[327] Steven X. Cadrin. Interdisciplinary analysis of yellowtail flounder stock structure off New England. Reviews in Fisheries Science, 18(3):281– 299, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2010.506251.

Parsons:2010:QAS

[328] Amber L. Parsons and John R. Skalski. Quantitative assessment of salmonid escapement techniques. Reviews in Fisheries Science, 18(4): 301–314, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2010.513020.

Moorhead:2010:DCB

[329] Jonathan A. Moorhead and Chaoshu Zeng. Development of captive breeding techniques for marine ornamental fish: a review. Reviews in Fisheries Science, 18(4):315–343, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2010.516035.

Parkes:2010:BSG

[330] Graeme Parkes, James A. Young, Suzannah F. Walmsley, Rigmor Abel, Jon Harman, Peter Horvat, Audun Lem, Alastair MacFarlane, Maarten Mens, and Conor Nolan. Behind the signs — a global review of fish sustainability information schemes. Reviews in Fisheries Science, 18(4): 344–356, 2010. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2010.516374.

Howell:2010:RPF

[331] W. H. Howell. A review of “Practical flatfish culture and stock enhancement”. Reviews in Fisheries Science, 18(4):357–358, 2010. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2010.519994.

Ward:2011:UDE

[332] Tim M. Ward, Paul Burch, Lachlan J. McLeay, and Alex R. Ivey. Use of the daily egg production method for stock assessment of sardine, Sardinops sagax; lessons learned over a decade of application off southern Australia. Reviews in Fisheries Science, 19(1):1–20, 2011. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2010.528711.

La:2011:ASP

[333] Van T. La and Steven J. Cooke. Advancing the science and practice of fish kill investigations. Reviews in Fisheries Science, 19(1): 21–33, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2010.531793.

Voultsiadou:2011:ABS

[334] E. Voultsiadou, T. Dailianis, C. Antoniadou, D. Vafidis, C. Dounas, and C. C. Chintiroglou. Aegean bath sponges: Historical data and current status. Reviews in Fisheries Science, 19(1):34–51, 2011. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2010.531794.

Sinclair:2011:SSG

[335] William Sinclair, Stephen J. Newman, Gabriel M. S. Vianna, Steven Williams, and William J. Aspden. Spatial subdivision and genetic diversity in populations on the East and West Coasts of Australia: The multifaceted case of Nautilus pompilius (Mollusca, Cephalopoda). Reviews in Fisheries Science, 19(1):52–61, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2010.533794.

Vinson:2011:SEF

[336] M. R. Vinson and T. R. Angradi. Stomach emptiness in fishes: Sources of variation and study design implications. Reviews in Fisheries Science, 19(2):63–73, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2010.536856.

Newman:2011:API

[337] Stephen J. Newman, Craig L. Skepper, Gabby E. A. Mitsopoulos, Corey B. Wakefield, Jessica J. Meeuwig, and Euan S. Harvey. Assessment of the potential impacts of trap usage and ghost fishing on the northern demersal scalefish fishery. Reviews in Fisheries Science, 19 (2):74–84, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2010.543961.

Bagchi:2011:FFI

[338] Anita Bagchi and Prithwiraj Jha. Fish and fisheries in Indian heritage and development of pisciculture in India. Reviews in Fisheries Science, 19 (2):85–118, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2010.535046.

Goethel:2011:ISS

[339] Daniel R. Goethel, Terrance J. Quinn II, and Steven X. Cadrin. Incorporating spatial structure in stock assessment: Movement modeling in marine fish population dynamics. Reviews in Fisheries Science, 19(2): 119–136, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2011.557451.

Levesque:2011:TFR

[340] Juan C. Levesque. Is today’s fisheries research driven by the economic value of a species? A case study using an updated review of ladyfish (Elops saurus) biology and ecology. Reviews in Fisheries Science, 19(2): 137–149, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2011.557450.

MachadodaRocha:2011:PHF

[341] Carlos Alberto Machado da Rocha. Parasitic helminths of the freshwater neotropical fish Hoplias malabaricus (Characiformes, Erythrinidae) from South America basins. Reviews in Fisheries Science, 19(2):150– 156, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2011.557752.

Wagner:2011:EFS

[342] Eric L. Wagner and P. Dee Boersma. Effects of fisheries on seabird community ecology. Reviews in Fisheries Science, 19(3):157–167, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2011.562568.

Gillson:2011:FFF

[343] Jonathan Gillson. Freshwater flow and fisheries production in estuarine and coastal systems: Where a drop of rain is not lost. Reviews in Fisheries Science, 19(3):168–186, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2011.560690.

Li:2011:AIC

[344] Xuepeng Li, Jianrong Li, Yanbo Wang, Linglin Fu, Yuying Fu, Biqing Li, and Binghua Jiao. Aquaculture industry in China: Current state, challenges, and outlook. Reviews in Fisheries Science, 19(3):187– 200, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2011.573597.

Enes:2011:DCU

[345] P. Enes, S. Panserat, S. Kaushik, and A. Oliva-Teles. Dietary carbohydrate utilization by European sea bass (Dicentrarchus labrax L.) and gilthead sea bream (Sparus aurata L.) juveniles. Reviews in Fisheries Science, 19(3):201–215, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2011.579363.

Mayfield:2011:IDF

[346] S. Mayfield, R. C. Chick, I. J. Carlson, and T. M. Ward. Invertebrate dive fisheries can be sustainable: Forty years of production from a greenlip abalone fishery off southern Australia. Reviews in Fisheries Science, 19 (3):216–230, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2011.585431.

Diamond:2011:FDN

[347] Ben Diamond and Bryce D. Beukers-Stewart. Fisheries discards in the North Sea: Waste of resources or a necessary evil? Reviews in Fisheries Science, 19(3):231–245, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2011.585432.

Lobon-Cervia:2011:WFD

[348] Javier Lobón-Cerviá. Why fishing does not magnify temporal fluctuations in the population abundance of stream-living salmonids. Reviews in Fisheries Science, 19(3):246–256, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2011.587617.

Torrissen:2011:ASS

[349] Ole Torrissen, Rolf Erik Olsen, Reidar Toresen, Gro Ingunn Hemre, Albert G. J. Tacon, Frank Asche, Ronald W. Hardy, and Santosh Lall. Atlantic salmon (Salmo salar): The “Super-Chicken” of the sea? Reviews in Fisheries Science, 19(3):257–278, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2011.597890.

Scarnecchia:2011:SNG

[350] Dennis L. Scarnecchia, Brent D. Gordon, Jason D. Schooley, L. Fred Ryckman, Brad J. Schmitz, Shannon E. Miller, and Youngtaik Lim. Southern and Northern Great Plains (United States) paddlefish stocks within frameworks of acipenseriform life history and the metabolic theory of ecology. Reviews in Fisheries Science, 19(3):279–298, 2011. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2011.598123.

Sicuro:2011:SCM

[351] Benedetto Sicuro and Jay Levine. Sea cucumber in the Mediterranean: a potential species for aquaculture in the Mediterranean. Reviews in Fisheries Science, 19(3):299–304, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2011.598249.

Carey:2011:SBP

[352] Michael P. Carey, Beth L. Sanderson, Thomas A. Friesen, Katie A. Barnas, and Julian D. Olden. Smallmouth bass in the Pacific Northwest: a threat to native species; a benefit for anglers. Reviews in Fisheries Science, 19(3):305–315, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2011.598584.

Stickney:2011:RHA

[353] Robert R. Stickney. A review of “The History of Aquaculture”. Reviews in Fisheries Science, 19(3):316–319, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2011.589541.

Britton:2011:EEB

[354] J. R. Britton and J. Pegg. Ecology of European barbel Barbus barbus: Implications for river, fishery, and conservation management. Reviews in Fisheries Science, 19(4):321–330, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2011.599886.

Bouwman:2011:GHF

[355] A. F. Bouwman, M. Pawłowski, C. Liu, A. H. W. Beusen, S. E. Shumway, P. M. Glibert, and C. C. Overbeek. Global hindcasts and future projections of coastal nitrogen and phosphorus loads due to shellfish and seaweed aquaculture. Reviews in Fisheries Science, 19(4):331–357, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2011.603849.

Glibert:2011:ESB

[356] Patricia M. Glibert, David Fullerton, Joann M. Burkholder, Jeffrey C. Cornwell, and Todd M. Kana. Ecological stoichiometry, biogeochemical cycling, invasive species, and aquatic food webs: San Francisco Estuary and comparative systems. Reviews in Fisheries Science, 19(4): 358–417, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2011.611916.

Anonymous:2011:EBE

[357] Anonymous. Editorial board EOV. Reviews in Fisheries Science, 19 (4):??, 2011. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2011.631821.

Miller:2012:IFA

[358] William J. Miller, Bryan F. J. Manly, Dennis D. Murphy, David Fullerton, and Rob Roy Ramey. An investigation of factors affecting the decline of delta smelt (Hypomesus transpacificus) in the sacramento-san joaquin estuary. Reviews in Fisheries Science, 20(1):1–19, 2012. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2011.634930.

Brannon:2012:REV

[359] E. L. Brannon, K. Collins, M. A. Cronin, L. L. Moulton, A. L. Maki, and K. R. Parker. Review of the Exxon Valdez oil spill effects on pink salmon in Prince William Sound, Alaska. Reviews in Fisheries Science, 20(1):20–60, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2011.643697.

Martin:2012:EEC

[360] Sarah M. Martin, Tracy A. Cambridge, Chris Grieve, Fiona M. Nimmo, and David J. Agnew. An evaluation of environmental changes within fisheries involved in the Marine Stewardship Council Certification Scheme. Reviews in Fisheries Science, 20(2):61–69, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2011.654287.

Jaric:2012:TFS

[361] Ivan Jarić, Gorcin Cvijanović, Jelena Knezević-Jarić, and Mirjana Lenhardt. Trends in fisheries science from 2000 to 2009: a bibliometric study. Reviews in Fisheries Science, 20(2):70–79, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/10641262.2012.659775.

Wuenschel:2012:RPH

[362] Mark J. Wuenschel, Kenneth W. Able, Jeffrey A. Buckel, James W. Morley, Thomas Lankford, A. C. Branson, David O. Conover, Damien Drisco, Adrian Jordaan, Keith Dunton, David H. Secor, Ryan J. Woodland, Francis Juanes, and David Stormer. Recruitment patterns and habitat use of young-of-the-year bluefish along the United States East Coast: Insights from coordinated coastwide sampling. Reviews in Fisheries Science, 20(2):80–102, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2012.660999.

Wang:2012:IGF

[363] You-Gan Wang and Na Wang. Implications of gain functions in fisheries management. Reviews in Fisheries Science, 20(2):103–109, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2012.663421.

Stoner:2012:ASP

[364] Allan W. Stoner. Assessing stress and predicting mortality in economically significant crustaceans. Reviews in Fisheries Science, 20(3): 111–135, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2012.689025.

Edwards:2012:FAM

[365] C. T. T. Edwards, R. M. Hillary, P. Levontin, J. L. Blanchard, and K. Lorenzen. Fisheries assessment and management: a synthesis of common approaches with special reference to deepwater and data-poor stocks. Reviews in Fisheries Science, 20(3):136–153, 2012. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2012.683210.

Stokesbury:2012:SDR

[366] Kevin D. E. Stokesbury. Stock definition and recruitment: Implications for the U.S. sea scallop (Placopecten magellanicus) fishery from 2003 to 2011. Reviews in Fisheries Science, 20(3):154–164, 2012. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2012.695816.

Lloret:2012:FHF

[367] J. Lloret, E. Faliex, G. E. Shulman, J.-A. Raga, P. Sasal, M. Muñoz, M. Casadevall, A. E. Ahuir-Baraja, F. E. Montero, A. Repullés-Albelda, M. Cardinale, H.-J. Rätz, S. Vila, and D. Ferrer. Fish health and fisheries, implications for stock assessment and management: The Mediterranean example. Reviews in Fisheries Science, 20(3):165–180, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2012.695817.

Nickum:2012:RDF

[368] John G. Nickum. A review of “Do fish feel pain?”. Reviews in Fisheries Science, 20(3):181–182, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2012.696898.

Gomez-Chiarri:2012:RAF

[369] Marta Gomez-Chiarri. A review of “Aquaculture and Fisheries Biotechnology, Genetic Approaches, 2nd Edition”. Reviews in Fisheries Science, 20(3):183, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2012.707103.

Frazer:2012:CLI

[370] Thomas K. Frazer, Charles A. Jacoby, Morgan A. Edwards, Savanna C. Barry, and Carrie M. Manfrino. Coping with the lionfish invasion: Can targeted removals yield beneficial effects? Reviews in Fisheries Science, 20(4):185–191, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2012.700655.

Fraser:2012:WCT

[371] Thomas W. K. Fraser, Per Gunnar Fjelldal, Tom Hansen, and Ian Mayer. Welfare considerations of triploid fish. Reviews in Fisheries Science, 20 (4):192–211, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2012.704598.

Nelson:2012:RNS

[372] Emily J. Nelson, Bruce A. MacDonald, and Shawn M. C. Robinson. A review of the northern sea cucumber Cucumaria frondosa (Gunnerus, 1767) as a potential aquaculture species. Reviews in Fisheries Science, 20(4): 212–219, 2012. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2012.719043.

Mayfield:2012:FYS

[373] S. Mayfield, C. Mundy, H. Gorfine, A. M. Hart, and D. Worthington. Fifty years of sustained production from the Australian abalone fisheries. Reviews in Fisheries Science, 20(4):220–250, 2012. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2012.725434.

Rohde:2012:RFP

[374] Klaus Rohde. A review of “Fish parasites. pathobiology and protection”. Reviews in Fisheries Science, 20(4):251, 2012. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2012.711194.

Romano:2013:TEA

[375] Nicholas Romano and Chaoshu Zeng. Toxic effects of ammonia, nitrite, and nitrate to decapod crustaceans: a review on factors influencing their toxicity, physiological consequences, and coping mechanisms. Reviews in Fisheries Science, 21(1):1–21, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2012.753404.

Tacon:2013:FMI

[376] Albert G. J. Tacon and Marc Metian. Fish matters: Importance of aquatic foods in human nutrition and global food supply. Reviews in Fisheries Science, 21(1):22–38, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2012.753405.

McCain:2013:MLR

[377] Kathryn N. S. McCain. Moving large river ecology from past theories to future actions: a review. Reviews in Fisheries Science, 21(1): 39–48, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2012.753867.

Petersen:2013:BGS

[378] Elizabeth H. Petersen, Dam Thi My Chinh, Ngo Thi Diu, Vu Van Phuoc, Truong Ha Phuong, Nguyen Van Dung, Nguyen Khac Dat, Pham Truong Giang, and Brett D. Glencross. Bioeconomics of grouper, Serranidae: Epinephelinae, culture in Vietnam. Reviews in Fisheries Science, 21 (1):49–57, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2012.753403.

Neilson:2013:RAS

[379] John Neilson, Freddy Arocha, Shannon Cass-Calay, Jaime Mejuto, Mauricio Ortiz, Gerry Scott, Craig Smith, Paulo Travassos, George Tserpes, and Irene Andrushchenko. The recovery of Atlantic swordfish: The comparative roles of the regional fisheries management organization and species biology. Reviews in Fisheries Science, 21(2):59–97, 2013. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2012.754842.

Cohen:2013:TIT

[380] Felipe P. A. Cohen, Wagner C. Valenti, and Ricardo Calado. Traceability issues in the trade of marine ornamental species. Reviews in Fisheries Science, 21(2):98–111, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2012.760522.

Bouwman:2013:HFP

[381] A. F. Bouwman, A. H. W. Beusen, C. C. Overbeek, D. P. Bureau, M. Pawlowski, and P. M. Glibert. Hindcasts and future projections of global inland and coastal nitrogen and phosphorus loads due to finfish aquaculture. Reviews in Fisheries Science, 21(2):112–156, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2013.790340.

Large:2013:SWM

[382] Philip A. Large, David J. Agnew, José Ángel Álvarez Pérez, Christopher Barrio Froján, Rudi Cloete, Dimitrios Damalas, Leonie Dransfeld, Charles T. T. Edwards, Stephen Feist, Ivone Figueiredo, Fernando González, Juan Gil Herrera, Andrew Kenny, Klara Jakobsdóttir, Matt Longshaw, Pascal Lorance, Paul Marchal, Chryssi Mytilineou, Benjamin Planque, and Chrissi-Yianna Politou. Strengths and weaknesses of the management and monitoring of deep-water stocks, fisheries, and ecosystems in various areas of the world — a roadmap toward sustainable deepwater fisheries in the Northeast Atlantic? Reviews in Fisheries Science, 21(2):157–180, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2013.785475.

Cort:2013:DAB

[383] José L. Cort, Simeon Deguara, Txema Galaz, Begonya Mèlich, Iñaki Artetxe, Igor Arregi, John Neilson, Irene Andrushchenko, Alex Hanke, Miguel Neves dos Santos, Vicente Estruch, Molly Lutcavage, Jessica Knapp, Guillermo Compeán-Jiménez, Rafael Solana-Sansores, Antonio Belmonte, David Martínez, Corrado Piccinetti, Ai Kimoto, Piero Addis, Marta Velasco, José M. De la Serna, Dolores Godoy, Tevfik Ceyhan, Isik Oray, Saadet Karakulak, Leif Nøttestad, Antonio López, Oriol Ribalta, Noureddine Abid, and M’Hamed Idrissi. Determination of L_{max} for Atlantic bluefin tuna, Thunnus thynnus (L.), from meta-analysis of published and available biometric data. Reviews in Fisheries Science, 21(2): 181–212, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.793284.

Lorenzen:2013:ECM

[384] Kai Lorenzen, Ann-Lisbeth Agnalt, H. Lee Blankenship, Anson H. Hines, Kenneth M. Leber, Neil R. Loneragan, and Matthew D. Taylor. Evolving context and maturing science: Aquaculture-based enhancement and restoration enter the marine fisheries management toolbox. Reviews in Fisheries Science, 21(3–4):213–221, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2013.837358.

Loneragan:2013:MSE

[385] Neil R. Loneragan, Gregory I. Jenkins, and Matthew D. Taylor. Marine stock enhancement, restocking, and sea ranching in Australia: Future directions and a synthesis of two decades of research and development. Reviews in Fisheries Science, 21(3–4):222–236, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2013.796810.

Dantas:2013:GDC

[386] Hozana Leite Dantas, Miguel Arcanjo dos Santos Neto, Karine Kelly Cavalcanti Oliveira, William Severi, Fábio Mendonça Diniz, and Maria Raquel Moura Coimbra. Genetic diversity of captive and wild threatened catfish Pseudoplatystoma corruscans in the São Francisco River.

Reviews in Fisheries Science, 21(3–4):237–246, 2013. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2013.800787.

Walsh:2013:PRP

[387] Michelle L. Walsh, Hiroshi Fujimoto, Takeo Yamamoto, Tatsuya Yamada, Yoichi Takahashi, and Yoh Yamashita. Post-release performance and assessment of cage-conditioned Japanese flounder, Paralichthys olivaceus, in Wakasa Bay, Japan. Reviews in Fisheries Science, 21(3–4): 247–257, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.800785.

Green:2013:OHR

[388] Bridget S. Green, Hugh Pederson, and Caleb Gardner. Overlap of home ranges of resident and introduced southern rock lobster after translocation. Reviews in Fisheries Science, 21(3–4):258–266, 2013. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2013.799389.

Lee:2013:OCD

[389] Jonathan S. F. Lee, Eugene P. Tezak, and Barry A. Berejikian. Ontogenetic changes in dispersal and habitat use in hatchery-reared lingcod. Reviews in Fisheries Science, 21(3–4):267–275, 2013. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2013.799391.

Taylor:2013:RSU

[390] Matthew D. Taylor, Andrew V. Fairfax, and Iain M. Suthers. The race for space: Using acoustic telemetry to understand density-dependent emigration and habitat selection in a released predatory fish. Reviews in Fisheries Science, 21(3–4):276–285, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2013.796813.

Chin:2013:DOR

[391] Byung Sun Chin, Masahiro Nakagawa, Tsutomu Noda, Toshihiro Wada, and Yoh Yamashita. Determining optimal release habitat for black rockfish, Sebastes schlegelii: Examining growth rate, feeding condition, and return rate. Reviews in Fisheries Science, 21(3–4):286–298, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2013.837364.

Hart:2013:SEGa

[392] Anthony M. Hart, Lachlan W. S. Strain, Frank Fabris, Jamin Brown, and Mark Davidson. Stock enhancement in greenlip abalone part i: Long-term growth and survival. Reviews in Fisheries Science, 21(3–4): 299–309, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.812503.

Hart:2013:SEGb

[393] Anthony M. Hart, Frank Fabris, Lachlan W. S. Strain, Mark Davidson, and Jamin Brown. Stock enhancement in greenlip abalone part II: Population and ecological effects. Reviews in Fisheries Science, 21(3–4): 310–320, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.812505.

Chick:2013:RDW

[394] Rowan C. Chick, Duncan G. Worthington, and Michael J. Kingsford. Restocking depleted wild stocks-long-term survival and impact of released blacklip abalone (Haliotis rubra) on depleted wild populations in New South Wales, Australia. Reviews in Fisheries Science, 21(3–4): 321–340, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.837278.

Taylor:2013:GPI

[395] Matthew D. Taylor, Nathan P. Brennan, Kai Lorenzen, and Kenneth M. Leber. Generalized predatory impact model: a numerical approach for assessing trophic limits to hatchery releases and controlling related ecological risks. Reviews in Fisheries Science, 21(3–4):341–353, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2013.796815.

Hart:2013:SEGc

[396] Anthony M. Hart, Lachlan W. S. Strain, and S. Alex Hesp. Stock enhancement in greenlip abalone part III: Bioeconomic evaluation. Reviews in Fisheries Science, 21(3–4):354–374, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/10641262.2013.812506.

Prince:2013:CBA

[397] Jeremy D. Prince. Cost-benefit analysis of alternative techniques for rehabilitating abalone reefs depleted by abalone viral ganglioneuritis.

Reviews in Fisheries Science, 21(3–4):375–387, 2013. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/10641262.2013.837300.

Camp:2013:PLS

[398] Edward V. Camp, Kai Lorenzen, Robert N. M. Ahrens, Luiz Barbieri, and Kenneth M. Leber. Potentials and limitations of stock enhancement in marine recreational fisheries systems: an integrative review of Florida’s red drum enhancement. Reviews in Fisheries Science, 21(3–4): 388–402, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.838075.

Dolorosa:2013:TWT

[399] Roger G. Dolorosa, Alastair Grant, and Jennifer A. Gill. Translocation of wild Trochus niloticus: Prospects for enhancing depleted Philippine reefs. Reviews in Fisheries Science, 21(3–4):403–413, 2013. CODEN RF-SCFJ. ISSN 1064-1262 (print), 1547-6553 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2013.800773.

Dolorosa:2013:IDS

[400] Roger G. Dolorosa, Alastair Grant, Jennifer A. Gill, Arlene L. Avillanosa, and Benjamin J. Gonzales. Indoor and deep sub-tidal intermediate culture of Trochus niloticus for restocking. Reviews in Fisheries Science, 21(3–4): 414–423, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.800775.

Juinio-Menez:2013:PDC

[401] Marie Antonette Juinio-Meñez, Julissah C. Evangelio, Ronald D. Olavides, Marie Antonette S. Paña, Glycinea M. De Peralta, Christine Mae A. Edullantes, Bryan Dave R. Rodriguez, and Inggat Laya N. Casilagan. Population dynamics of cultured Holothuria scabra in a sea ranch: Implications for stock restoration. Reviews in Fisheries Science, 21(3–4): 424–432, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.837282.

Lebata-Ramos:2013:EPR

[402] Ma. Junemie Hazel Lebata-Ramos, Ellen Flor C. Doyola-Solis, Jeff Bogart R. Abrogueña, Hiroshi Ogata, Joemel G. Sumbing, and Rema C. Sibonga. Evaluation of post-release behavior, recapture, and growth rates of hatchery-reared abalone Haliotis asinina released in sagay marine reserve, Philippines. Reviews in Fisheries Science, 21(3–4):433– 440, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.836445.

Gardner:2013:BGC

[403] Michelle J. Gardner, Alan Cottingham, S. Alex Hesp, Jennifer A. Chaplin, Greg I. Jenkins, Nicole M. Phillips, and Ian C. Potter. Biological and genetic characteristics of restocked and wild Acanthopagrus butcheri (Sparidae) in a Southwestern Australian Estuary. Reviews in Fisheries Science, 21(3–4):441–453, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2013.796804.

Hamasaki:2013:CFK

[404] Katsuyuki Hamasaki and Shuichi Kitada. Catch fluctuation of kuruma prawn, Penaeus japonicus in Japan relative to ocean climate variability and a stock enhancement program. Reviews in Fisheries Science, 21(3– 4):454–468, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.800781.

Miyakoshi:2013:HCH

[405] Yasuyuki Miyakoshi, Mitsuhiro Nagata, Shuichi Kitada, and Masahide Kaeriyama. Historical and current hatchery programs and management of chum salmon in Hokkaido, Northern Japan. Reviews in Fisheries Science, 21(3–4):469–479, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print),

1547-6553 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2013.836446.

Anonymous:2013:EBE

[406] Anonymous. Editorial board EOV. Reviews in Fisheries Science, 21 (3–4):??, 2013. CODEN RFSCFJ. ISSN 1064-1262 (print), 1547-6553

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.864538.

Batista:2014:TAC

[407] Vandick S. Batista, Nidia N. Fabré, Ana C. M. Malhado, and Richard J. Ladle. Tropical artisanal coastal fisheries: Challenges and future directions. Reviews in Fisheries Science & Aquaculture, 22(1):1–15, 2014. CO-

DEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/10641262.2013.822463.

Scarnecchia:2014:VPA

[408] Dennis L. Scarnecchia, Youngtaik Lim, L. Fred Ryckman, K. Michael Backes, Shannon E. Miller, R. Scott Gangl, and Brad J. Schmitz. Virtual population analysis, episodic recruitment, and harvest management of paddlefish with applications to other acipenseriform fishes. Reviews in Fisheries Science & Aquaculture, 22(1):16–35, 2014. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/10641262.2013.830592.

Caputi:2014:CPS

[409] Nick Caputi, Simon de Lestang, Anthony Hart, Mervi Kangas, Danielle Johnston, and James Penn. Catch predictions in stock assessment and management of invertebrate fisheries using pre-recruit abundance-case studies from Western Australia. Reviews in Fisheries Science & Aquaculture, 22(1):36–54, 2014. CODEN ???? ISSN 2330-8249. URL http:/

/www.tandfonline.com/doi/full/10.1080/10641262.2013.832144.

Kennedy:2014:TCD

[410] Victor S. Kennedy. Technological constraints during the first 40 years of eastern oyster Crassostrea virginica aquaculture. Reviews in Fisheries Science & Aquaculture, 22(1):55–72, 2014. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.822464.

Font:2014:BEI

[411] Toni Font and Josep Lloret. Biological and ecological impacts derived from recreational fishing in Mediterranean coastal areas. Reviews in Fisheries Science & Aquaculture, 22(1):73–85, 2014. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2013.823907.

Siccha-Ramirez:2014:MIM

[412] Raquel Siccha-Ramirez, Naércio A. Menezes, Mauro Nirchio, Fausto Foresti, and Claudio Oliveira. Molecular identification of mullet species of the Atlantic South Caribbean and South America and the phylogeographic analysis of Mugil liza. Reviews in Fisheries Science & Aquaculture, 22(1):86–96, 2014. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2013.833583.

Scarnecchia:2014:NFC

[413] Dennis L. Scarnecchia, Youngtaik Lim, Sean P. Moran, Timothy D. Tholl, Joseph M. Dos Santos, and Kenneth Breidinger. Novel fish communities: Native and non-native species trends in two run-of-the-river reservoirs, Clark Fork River, Montana. Reviews in Fisheries Science & Aquaculture, 22(1):97–111, 2014. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/10641262.2013.838937.

Damalas:2014:AFP

[414] Dimitrios Damalas, Christos D. Maravelias, and Stefanos Kavadas. Advances in fishing power: a study spanning 50 years. Reviews in Fisheries Science & Aquaculture, 22(1):112–121, 2014. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
10641262.2013.839620.

Smith:2014:RTF

[415] David L. Smith, R. Andrew Goodwin, and John M. Nestler. Relating turbulence and fish habitat: a new approach for management and research. Reviews in Fisheries Science & Aquaculture, 22(2):123–130, 2014.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/10641262.2013.803516.

Linh-Son:2014:MUM

[416] Nguyen Linh-Son, Quentin Grafton, To Dieu-Hang, and Tom Kompas. Multiple-use management strategies and marine reserves. Reviews in Fisheries Science & Aquaculture, 22(2):131–141, 2014. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/10641262.2013.842536.

Grabowski:2014:AVM

[417] Jonathan H. Grabowski, Michelle Bachman, Chad Demarest, Steve Eayrs, Bradley P. Harris, Vincent Malkoski, David Packer, and David Stevenson. Assessing the vulnerability of marine benthos to fishing gear impacts. Reviews in Fisheries Science & Aquaculture, 22(2):142–155, 2014. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/10641262.2013.846292.

Eigaard:2014:TDF

[418] Ole Ritzau Eigaard, Paul Marchal, Henrik Gislason, and Adriaan D. Rijnsdorp. Technological development and fisheries management. Reviews in Fisheries Science & Aquaculture, 22(2):156–174, 2014. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2014.899557.

Nair:2014:RFT

[419] Rekha J. Nair and A. Gopalakrishnan. A review on the fisheries, taxonomy, and status of the flatfishes in tropical waters. Reviews in Fisheries Science & Aquaculture, 22(3):175–183, 2014. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2014.902917.

Metian:2014:FBT

[420] Marc Metian, Simon Pouil, André Boustany, and Max Troell. Farming of bluefin tuna-reconsidering global estimates and sustainability concerns. Reviews in Fisheries Science & Aquaculture, 22(3):184–192, 2014. CO-

DEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2014.907771.

Stokesbury:2014:USI

[421] Kevin D. E. Stokesbury, Michael J. W. Stokesbury, Matthew T. Balazik, and Michael J. Dadswell. Use of the SAFE index to evaluate the status of a Summer aggregation of Atlantic sturgeon in minas basin, Canada, and the implication of the index for the USA endangered species designation of Atlantic and shortnose sturgeons. Reviews in Fisheries Science & Aquaculture, 22(3):193–206, 2014. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2014.913005.

Brown:2014:MTG

[422] Christopher L. Brown, Elisabeth Criscuolo Urbinati, Weimin Zhang, Shannon B. Brown, and Michelle McComb-Kobza. Maternal thyroid and glucocorticoid hormone interactions in larval fish development, and their applications in aquaculture. Reviews in Fisheries Science & Aquaculture, 22(3):207–220, 2014. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2014.918086.

Laghari:2014:IQT

[423] Muhammad Y. Laghari, Punhal Lashari, Yan Zhang, and Xiaowen Sun. Identification of quantitative trait loci (QTLs) in aquaculture species. Reviews in Fisheries Science & Aquaculture, 22(3):221–238, 2014. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2014.931172.

Cort:2014:VGE

[424] José L. Cort, Igor Arregui, Vicente D. Estruch, and Simeon Deguara. Validation of the growth equation applicable to the Eastern Atlantic bluefin tuna, Thunnus thynnus (L.), using L_{max}, tag-recapture, and first dorsal spine analysis. Reviews in Fisheries Science & Aquaculture, 22(3):239–255, 2014. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2014.931173.

Damalas:2014:ADH

[425] Dimitrios Damalas and Persefoni Megalofonou. Age-dependent habitat identification of Mediterranean swordfish: Application on commercial fishery data and potential use in fisheries management. Reviews in Fisheries Science & Aquaculture, 22(4):257–273, 2014. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2014.951987.

Nunoo:2014:MFC

[426] F. K. E. Nunoo, B. Asiedu, K. Amador, Dyhia Belhabib, Vicky Lam, Rashid Sumaila, and Daniel Pauly. Marine fisheries catches in Ghana: Historic reconstruction for 1950 to 2010 and current economic impacts. Reviews in Fisheries Science & Aquaculture, 22(4):274–283, 2014. CO-

DEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2014.962687.

Carlson:2014:SEH

[427] Andrew K. Carlson and Bruce Vondracek. Synthesis of ecology and human dimensions for predictive management of bighead and silver carp in the United States. Reviews in Fisheries Science & Aquaculture, 22(4):284–300, 2014. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2014.967747.

Anonymous:2014:EBE

[428] Anonymous. Editorial board EOV. Reviews in Fisheries Science & Aquaculture, 22(4):??, 2014. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2014.980701.

Tacon:2015:FMS

[429] Albert G. J. Tacon and Marc Metian. Feed matters: Satisfying the feed demand of aquaculture. Reviews in Fisheries Science & Aquaculture, 23(1):1–10, 2015. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2014.987209.

Apollonio:2015:HCE

[430] Spencer Apollonio. A hierarchical concept of ecosystems: an alternative to sector management for multi-species fisheries. Reviews in Fisheries Science & Aquaculture, 23(1):11–22, 2015. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2014.988867.

Cort:2015:VLW

[431] José L. Cort, Vicente D. Estruch, Miguel Neves Dos Santos, Antonio Di Natale, Noureddine Abid, and José M. de la Serna. On the variability of the length–weight relationship for Atlantic bluefin tuna, Thunnus thynnus (L.). Reviews in Fisheries Science & Aquaculture, 23(1):23–38, 2015.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2015.1008625.

Stankovic:2015:RTE

[432] David Stanković, Alain Jean Crivelli, and Ales Snoj. Rainbow trout in Europe: Introduction, naturalization, and impacts. Reviews in Fisheries Science & Aquaculture, 23(1):39–71, 2015. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1024825.

Kumar:2015:PGS

[433] Girish Kumar and Martin Kocour. Population genetic structure of tunas inferred from molecular markers: a review. Reviews in Fisheries Science & Aquaculture, 23(1):72–89, 2015. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1024826.

Arkhipkin:2015:P

[434] Alexander Arkhipkin. Preface. Reviews in Fisheries Science & Aquaculture, 23(2):91, 2015. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2015.1026228.

Arkhipkin:2015:WSF

[435] Alexander I. Arkhipkin, Paul G. K. Rodhouse, Graham J. Pierce, Warwick Sauer, Mitsuo Sakai, Louise Allcock, Juan Arguelles, John R. Bower, Gladis Castillo, Luca Ceriola, Chih-Shin Chen, Xinjun Chen, Mariana Diaz-Santana, Nicola Downey, Angel F. González, Jasmin Granados Amores, Corey P. Green, Angel Guerra, Lisa C. Hendrickson, Christian Ibáñez, Kingo Ito, Patrizia Jereb, Yoshiki Kato, Oleg N. Katugin, Mitsuhisa Kawano, Hideaki Kidokoro, Vladimir V. Kulik, Vladimir V. Laptikhovsky, Marek R. Lipinski, Bilin Liu, Luis Mariátegui, Wilbert Marin, Ana Medina, Katsuhiro Miki, Kazutaka Miyahara, Natalie Moltschaniwskyj, Hassan Moustahfid, Jaruwat Nabhitabhata, Nobuaki Nanjo, Chingis M. Nigmatullin, Tetsuya Ohtani, Gretta Pecl, J. Angel A. Perez, Uwe Piatkowski, Pirochana Saikliang, Cesar A. Salinas-Zavala, Michael Steer, Yongjun Tian, Yukio Ueta, Dharmamony Vijai, Toshie Wakabayashi, Tadanori Yamaguchi, Carmen Yamashiro, Norio Yamashita, and Louis D. Zeidberg. World squid fisheries. Reviews in Fisheries Science & Aquaculture, 23(2):92–252, 2015. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2015.1026226.

Vilizzi:2015:EEC

[436] L. Vilizzi, A. S. Tarkan, and G. H. Copp. Experimental evidence from causal criteria analysis for the effects of common carp Cyprinus carpio on freshwater ecosystems: a global perspective. Reviews in Fisheries Science & Aquaculture, 23(3):253–290, 2015. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1051214.

Gupta:2015:SRF

[437] Nishikant Gupta, Shannon D. Bower, Rajeev Raghavan, Andy J. Danylchuk, and Steven J. Cooke. Status of recreational fisheries in India: Development, issues, and opportunities. Reviews in Fisheries Science & Aquaculture, 23(3):291–301, 2015. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1052366.

Samarin:2015:FOA

[438] Azin Mohagheghi Samarin, Tomas Policar, and Franz Lahnsteiner. Fish oocyte ageing and its effect on egg quality. Reviews in Fisheries Science & Aquaculture, 23(3):302–314, 2015. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1053560.

Hoseinifar:2015:PFI

[439] Seyed Hossein Hoseinifar, Maria Ángeles Esteban, Alberto Cuesta, and Yun-Zhang Sun. Prebiotics and fish immune response: a review of current knowledge and future perspectives. Reviews in Fisheries Science & Aquaculture, 23(4):315–328, 2015. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2015.1052365.

Spurgeon:2015:GSF

[440] Jonathan J. Spurgeon, Martin J. Hamel, Kevin L. Pope, and Mark A. Pegg. The global status of freshwater fish age validation studies and a prioritization framework for further research. Reviews in Fisheries Science & Aquaculture, 23(4):329–345, 2015. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1068737.

Cort:2015:FTT

[441] José L. Cort and Pablo Abaunza. The fall of the tuna traps and the collapse of the Atlantic bluefin tuna, Thunnus thynnus (L.), fisheries of Northern Europe from the 1960s. Reviews in Fisheries Science & Aquaculture, 23(4):346–373, 2015. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2015.1079166.

Anonymous:2015:EBE

[442] Anonymous. Editorial board EOV. Reviews in Fisheries Science & Aquaculture, 23(4):??, 2015. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2015.1093842.

Estrada:2016:CRC

[443] Uriel Rodriguez Estrada, Fanny Ayumi Yasumaru, Albert G. J. Tacon, and Daniel Lemos. Cobia (Rachycentron canadum): a selected annotated bibliography on aquaculture, general biology and fisheries 1967–2015. Reviews in Fisheries Science & Aquaculture, 24(1):1–97, 2016. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2015.1088821.

Morales-Bojorquez:2016:JSD

[444] Enrique Morales-Bojórquez and José Luis Pacheco-Bedoya. Jumbo squid Dosidicus gigas: a new fishery in Ecuador. Reviews in Fisheries Science & Aquaculture, 24(1):98–110, 2016. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1102862.

White:2016:BR

[445] J. Wilson White. Book review. Reviews in Fisheries Science & Aquaculture, 24(1):111–112, 2016. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2015.1103603.

Perdikaris:2016:EFP

[446] Constantinos Perdikaris, Anna Chrysafi, and Kostas Ganias. Environmentally friendly practices and perceptions in aquaculture: a sectoral case-study from a Mediterranean-based industry. Reviews in Fisheries Science & Aquaculture, 24(2):113–125, 2016. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1112358.

Cort:2016:ALW

[447] José L. Cort and Vicente D. Estruch. Analysis of the length–weight relationships for the Western Atlantic bluefin tuna, Thunnus thynnus (L.).

Reviews in Fisheries Science & Aquaculture, 24(2):126–135, 2016. CO-

DEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2015.1112359.

Kumar:2016:TAL

[448] Ganesh Kumar and Carole R. Engle. Technological advances that led to growth of shrimp, salmon, and tilapia farming. Reviews in Fisheries Science & Aquaculture, 24(2):136–152, 2016. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1112357.

Carlson:2016:TNP

[449] Andrew K. Carlson. Trophy northern pike: The value of experimentation and public engagement. Reviews in Fisheries Science & Aquaculture, 24 (2):153–159, 2016. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2015.1112765.

Tessier:2016:FAL

[450] Anne Tessier, Stephane Descloux, Raymond Lae, Maud Cottet, Pierre Guedant, and Jean Guillard. Fish assemblages in large tropical reservoirs: Overview of fish population monitoring methods. Reviews in Fisheries Science & Aquaculture, 24(2):160–177, 2016. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2015.1112766.

Deines:2016:TAE

[451] Andrew M. Deines, Marion E. Wittmann, Jillian M. Deines, and David M. Lodge. Tradeoffs among ecosystem services associated with global tilapia introductions. Reviews in Fisheries Science & Aquaculture, 24(2):178–191, 2016. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2015.1115466.

Gorelli:2016:FEI

[452] Giulia Gorelli, Francisco Sardà, and Joan B. Company. Fishing effort increase and resource status of the deep-sea red shrimp Aristeus antennatus (Risso 1816) in the Northwest Mediterranean Sea since the 1950s. Reviews in Fisheries Science & Aquaculture, 24(2):192–202, 2016. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2015.1119799.

Gupta:2016:OMB

[453] Sandipan Gupta. An overview on morphology, biology, and culture of spotted scat Scatophagus argus (Linnaeus 1766). Reviews in Fisheries Science & Aquaculture, 24(2):203–212, 2016. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2015.1119800.

Knapp:2016:PEM

[454] Gunnar Knapp and Michael C. Rubino. The political economics of marine aquaculture in the United States. Reviews in Fisheries Science & Aquaculture, 24(3):213–229, 2016. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2015.1121202.

Plotnek:2016:UMC

[455] Emma Plotnek, Felipe Paredes, Mauricio Galvez, and Mónica Pérez-Ramírez. From unsustainability to MSC certification: a case study of the artisanal Chilean South Pacific hake fishery. Reviews in Fisheries Science & Aquaculture, 24(3):230–243, 2016. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2016.1161003.

Liu:2016:EHC

[456] Owen R. Liu, Lennon R. Thomas, Michaela Clemence, Rod Fujita, Jacob P. Kritzer, Gavin McDonald, and Cody Szuwalski. An evaluation of harvest control methods for fishery management. Reviews in Fisheries Science & Aquaculture, 24(3):244–263, 2016. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2016.1161002.

Marques:2016:IFP

[457] Helcio L. A. Marques, Michael B. New, Marcello Villar Boock, Helenice Pereira Barros, Margarete Mallasen, and Wagner C. Valenti. Integrated freshwater prawn farming: State-of-the-art and future potential. Reviews in Fisheries Science & Aquaculture, 24(3):264–293, 2016. CO-

DEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2016.1169245.

Kumar:2016:BR

[458] Ganesh Kumar. Book review. Reviews in Fisheries Science & Aquaculture, 24(3):294, 2016. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2016.1142304.

Kumar:2016:BBL

[459] K. V. Aneesh Kumar, P. Pravin, and B. Meenakumari. Bait, bait loss, and depredation in pelagic longline fisheries — a review. Reviews in Fisheries Science & Aquaculture, 24(4):295–304, 2016. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2016.1162134.

Kebede:2016:FMP

[460] Mathewos Temesgen Kebede and Abebe Getahun Gubale. Fishery management problems in Ethiopia: Natural and human induced impacts and the conservation challenges. Reviews in Fisheries Science & Aquaculture, 24(4):305–313, 2016. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2016.1187582.

Sicuro:2016:SSS

[461] Benedetto Sicuro and Umberto Luzzana. The state of Seriola spp. other than yellowtail (S. quinqueradiata) farming in the world. Reviews in Fisheries Science & Aquaculture, 24(4):314–325, 2016. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2016.1187583.

Han:2016:RSC

[462] Qingxi Han, John K. Keesing, and Dongyan Liu. A review of sea cucumber aquaculture, ranching, and stock enhancement in China. Reviews in Fisheries Science & Aquaculture, 24(4):326–341, 2016. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2016.1193472.

Kumar:2016:APS

[463] Vikash Kumar, Suvra Roy, Dharmendra Kumar Meena, and Uttam Kumar Sarkar. Application of probiotics in shrimp aquaculture: Importance, mechanisms of action, and methods of administration. Reviews in Fisheries Science & Aquaculture, 24(4):342–368, 2016. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2016.1193841.

Avigliano:2016:RAO

[464] Esteban Avigliano and Alejandra Vanina Volpedo. A review of the application of otolith microchemistry toward the study of Latin American fishes. Reviews in Fisheries Science & Aquaculture, 24(4):369–384, 2016.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2016.1202189.

Anonymous:2016:EBE

[465] Anonymous. Editorial board EOV. Reviews in Fisheries Science & Aquaculture, 24(4):??, 2016. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2016.1213992.

Arlinghaus:2017:UMF

[466] R. Arlinghaus, J. Alós, B. Beardmore, K. Daedlow, M. Dorow, M. Fujitani, D. Hühn, W. Haider, L. M. Hunt, B. M. Johnson, F. Johnston, T. Klefoth, S. Matsumura, C. Monk, T. Pagel, J. R. Post, T. Rapp, C. Riepe, H. Ward, and C. Wolter. Understanding and managing freshwater recreational fisheries as complex adaptive social-ecological systems. Reviews in Fisheries Science & Aquaculture, 25(1):1–41, 2017. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2016.1209160.

Romano:2017:CDC

[467] Nicholas Romano and Chaoshu Zeng. Cannibalism of decapod crustaceans and implications for their aquaculture: a review of its prevalence, influencing factors, and mitigating methods. Reviews in Fisheries Science & Aquaculture, 25(1):42–69, 2017. CODEN ???? ISSN 2330-8249. URL http:/

/www.tandfonline.com/doi/full/10.1080/23308249.2016.1221379.

Rasmussen:2017:IMS

[468] Josh E. Rasmussen and Mark C. Belk. Individual movement of stream fishes: Linking ecological drivers with evolutionary processes. Reviews in Fisheries Science & Aquaculture, 25(1):70–83, 2017. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2016.1232697.

Guo:2017:SLP

[469] Z. Guo, D. Andreou, and J. R. Britton. Sea lamprey Petromyzon marinus biology and management across their native and invasive ranges: Promoting conservation by knowledge transfer. Reviews in Fisheries Science & Aquaculture, 25(1):84–99, 2017. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2016.1233166.

Cohen:2017:SAB

[470] Felipe P. A. Cohen, Wagner C. Valenti, Miquel Planas, and Ricardo Calado. Seahorse aquaculture, biology and conservation: Knowledge gaps and research opportunities. Reviews in Fisheries Science & Aquaculture, 25(1):100–111, 2017. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2016.1237469.

Choudhury:2017:ABR

[471] Tanmoy Gon Choudhury, Vinay Tharabenahalli Nagaraju, Samchetshabam Gita, Anutosh Paria, and Janmejay Parhi. Advances in bacteriophage research for bacterial disease control in aquaculture. Reviews in Fisheries Science & Aquaculture, 25(2):113–125, 2017. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2016.1241977.

Apraku:2017:TSD

[472] Andrews Apraku, Liping Liu, and Christian Larbi Ayisi. Trends and status of dietary coconut oil in aquaculture feeds. Reviews in Fisheries Science & Aquaculture, 25(2):126–132, 2017. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2016.1245275.

Morais:2017:PTF

[473] Sofia Morais. The physiology of taste in fish: Potential implications for feeding stimulation and gut chemical sensing. Reviews in Fisheries Science & Aquaculture, 25(2):133–149, 2017. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2016.1249279.

Roxo:2017:PCF

[474] André Roxo, Susana Mendes, and João Correia. Portuguese commercial fisheries of swordfish, Xiphias gladius. Reviews in Fisheries Science & Aquaculture, 25(2):150–157, 2017. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2016.1251879.

ElSheikha:2017:TKS

[475] Aly Farag El Sheikha and Jianping (Jp) Xu. Traceability as a key of seafood safety: Reassessment and possible applications. Reviews in Fisheries Science & Aquaculture, 25(2):158–170, 2017. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2016.1254158.

Walton:2017:DTM

[476] S. E. Walton, H. M. Gan, R. Raghavan, A. C. Pinder, and A. Ahmad. Disentangling the taxonomy of the mahseers (Tor spp.) of Malaysia: an integrated approach using morphology, genetics and historical records. Reviews in Fisheries Science & Aquaculture, 25(3):171–183, 2017. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2016.1251391.

Dadar:2017:AAV

[477] Maryam Dadar, Kuldeep Dhama, Vikram N. Vakharia, Seyed Hossein Hoseinifar, Kumaragurubaran Karthik, Ruchi Tiwari, Rekha Khandia, Ashok Munjal, Celene Salgado-Miranda, and Sunil K. Joshi. Advances in aquaculture vaccines against fish pathogens: Global status and current trends. Reviews in Fisheries Science & Aquaculture, 25(3):184–217, 2017.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2016.1261277.

Monteiro:2017:PSF

[478] Pedro Valadas Monteiro. The purse Seine fishing of sardine in Portuguese waters: a difficult compromise between fish stock sustainability and fishing effort. Reviews in Fisheries Science & Aquaculture, 25(3):218–229, 2017.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2016.1269720.

Gucu:2017:BSA

[479] Ali Cemal Gücü, Yasar Genç, Murat Dagtekin, Serdar Sakinan, Orhan Ak, Meltem Ok, and Ilhan Aydin. On black sea anchovy and its fishery. Reviews in Fisheries Science & Aquaculture, 25(3):230–244, 2017.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2016.1276152.

Dash:2017:BLV

[480] Pragyan Dash, Satheesha Avunje, Ritesh S. Tandel, Sandeep K. P., and Akshaya Panigrahi. Biocontrol of luminous vibriosis in shrimp aquaculture: a review of current approaches and future perspectives. Reviews in Fisheries Science & Aquaculture, 25(3):245–255, 2017. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2016.1277973.

Anonymous:2017:C

[481] Anonymous. Corrigendum. Reviews in Fisheries Science & Aquaculture, 25(3):256, 2017. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2017.1286874.

Rather:2017:DMB

[482] Irfan A. Rather, Richelle Galope, Vivek K. Bajpai, Jeongheui Lim, Woon Kee Paek, and Yong-Ha Park. Diversity of marine bacteria and their bacteriocins: Applications in aquaculture. Reviews in Fisheries Science & Aquaculture, 25(4):257–269, 2017. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2017.1282417.

Kang:2017:IFC

[483] Bin Kang, Xiaoxia Huang, Jun Li, Min Liu, Longgen Guo, and Chiao-Chuan Han. Inland fisheries in China: Past, present, and future. Reviews in Fisheries Science & Aquaculture, 25(4):270–285, 2017. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2017.1285863.

Mies:2017:PGC

[484] M. Mies, P. Dor, A. Z. Güth, and P. Y. G. Sumida. Production in giant clam aquaculture: Trends and challenges. Reviews in Fisheries Science & Aquaculture, 25(4):286–296, 2017. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2017.1285864.

Costa:2017:RPI

[485] Graça Costa, Ken MacKenzie, and Marcelo E. Oliva. A review of the parasites infecting fishes of the genus Trachurus (Pisces: Carangidae). Reviews in Fisheries Science & Aquaculture, 25(4):297–315, 2017. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2017.1293607.

Fonseca:2017:LAM

[486] Tamara Fonseca, Barry A. Costa-Pierce, and Wagner C. Valenti. Lambari aquaculture as a means for the sustainable development of rural communities in Brazil. Reviews in Fisheries Science & Aquaculture, 25(4):316–330, 2017. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2017.1320647.

Parkyn:2017:BR

[487] Daryl Parkyn. Book review. Reviews in Fisheries Science & Aquaculture, 25(4):331, 2017. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2017.1327283.

Anonymous:2017:EBE

[488] Anonymous. Editorial board EOV. Reviews in Fisheries Science & Aquaculture, 25(4):??, 2017. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2017.1337437.

Grabowski:2018:HAJ

[489] Jonathan H. Grabowski, Christian W. Conroy, Rachel K. Gittman, Joseph T. Kelley, Sally Sherman, Graham D. Sherwood, and Gail Wippelhauser. Habitat associations of juvenile cod in nearshore waters. Reviews in Fisheries Science & Aquaculture, 26(1):1–14, 2018. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2017.1328660.

Tacon:2018:FMF

[490] Albert G. J. Tacon and Marc Metian. Food matters: Fish, income, and food supply — a comparative analysis. Reviews in Fisheries Science & Aquaculture, 26(1):15–28, 2018. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2017.1328659.

Vinay:2018:RAA

[491] Tharabenahalli Nagaraju Vinay, Shreelatha Bhat, Tanmoy Gon Choudhury, Anutosh Paria, Myung-Hwa Jung, Girisha Shivani Kallappa, and Sung-Ju Jung. Recent advances in application of nanoparticles in fish vaccine delivery. Reviews in Fisheries Science & Aquaculture, 26(1):29–41, 2018. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2017.1334625.

Lorentzen:2018:CSR

[492] Grete Lorentzen, Gøril Voldnes, Ragnhild D. Whitaker, Ingrid Kvalvik, Birthe Vang, Runar Gjerp Solstad, Marte R. Thomassen, and Sten I. Siikavuopio. Current status of the red King crab (Paralithodes camtchaticus) and snow crab (Chionoecetes opilio) industries in Norway. Reviews in Fisheries Science & Aquaculture, 26(1):42–54, 2018. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2017.1335284.

Santos:2018:SAFa

[493] Sérgio Ricardo Santos and Marcelo Vianna. Scientometric analysis of the fisheries science for the species of Cynoscion (Sciaenidae: Perciformes) from the Western Atlantic, with emphasis in the comparison of the North American and Brazilian fisheries catch data. Reviews in Fisheries Science & Aquaculture, 26(1):55–69, 2018. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2017.1337078.

Tlusty:2018:COM

[494] Michael F. Tlusty, Brooke Wikgren, Kerry Lagueux, Hauke Kite-Powell, Di Jin, Porter Hoagland, Robert D. Kenney, and Scott D. Kraus. Cooccurrence mapping of disparate data sets to assess potential aquaculture sites in the Gulf of Maine. Reviews in Fisheries Science & Aquaculture, 26(1):70–85, 2018. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2017.1343798.

Nickum:2018:AAM

[495] Mary J. Nickum, Michael Masser, Robert Reigh, and John G. Nickum. Alligator (Alligator mississippiensis) aquaculture in the United States. Reviews in Fisheries Science & Aquaculture, 26(1):86–98, 2018. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2017.1355350.

Shaalan:2018:AEI

[496] Mohamed Shaalan, Magdy El-Mahdy, Mona Saleh, and Mansour El-Matbouli. Aquaculture in Egypt: Insights on the current trends and future perspectives for sustainable development. Reviews in Fisheries Science & Aquaculture, 26(1):99–110, 2018. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2017.1358696.

Kumar:2018:MCF

[497] Girish Kumar. Mercury concentrations in fresh and canned tuna: a review. Reviews in Fisheries Science & Aquaculture, 26(1):111–120, 2018.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2017.1362370.

Escobar:2018:AIS

[498] Luis E. Escobar, Sophie Mallez, Michael McCartney, Christine Lee, Daniel P. Zielinski, Ratna Ghosal, Przemyslaw G. Bajer, Carli Wagner, Becca Nash, Megan Tomamichel, Paul Venturelli, Prince P. Mathai, Adam Kokotovich, Joaquin Escobar-Dodero, and Nicholas B. D. Phelps. Aquatic invasive species in the great lakes region: an overview. Reviews in Fisheries Science & Aquaculture, 26(1):121–138, 2018. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2017.1363715.

Ahmed:2018:BGW

[499] Nesar Ahmed, James D. Ward, Shirley Thompson, Christopher P. Saint, and James S. Diana. Blue-green water nexus in aquaculture for resilience to climate change. Reviews in Fisheries Science & Aquaculture, 26(2): 139–154, 2018. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2017.1373743.

Sarkar:2018:SPT

[500] U. K. Sarkar, K. M. Sandhya, P. Mishal, G. Karnatak, Lianthuamluaia, S. Kumari, P. Panikkar, R. Palaniswamy, M. Karthikeyan, S. Sibina Mol, T. T. Paul, V. L. Ramya, D. S. K. Rao, M. Feroz Khan, D. Panda, and B. K. Das. Status, prospects, threats, and the way forward for sustainable management and enhancement of the tropical Indian reservoir fisheries: an overview. Reviews in Fisheries Science & Aquaculture, 26(2):155–175, 2018. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2017.1373744.

Fainburg:2018:CKC

[501] Leandro A. Fainburg and Pedro J. Fernández Iriarte. Current knowledge on commercially important fish stocks of the buenos aires coastal ecosystem. Reviews in Fisheries Science & Aquaculture, 26(2):176–182, 2018.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2017.1374343.

Nguyen:2018:MEP

[502] Tram Anh Thi Nguyen and Curtis M. Jolly. Macro-economic and product challenges facing Vietnamese the Pangasius industry. Reviews in Fisheries Science & Aquaculture, 26(2):183–194, 2018. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2017.1379948.

Munoz-Lechuga:2018:OFM

[503] Rubén Muñoz-Lechuga, Víctor Sanz-Fernández, and Remedios Cabrera-Castro. An overview of freshwater and marine finfish aquaculture in Spain: Emphasis on regions. Reviews in Fisheries Science & Aquaculture, 26 (2):195–213, 2018. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2017.1381832.

Arechavala-Lopez:2018:ISB

[504] P. Arechavala-Lopez, K. Toledo-Guedes, D. Izquierdo-Gomez, T. Segvić-Bubić, and P. Sanchez-Jerez. Implications of sea bream and sea bass escapes for sustainable aquaculture management: a review of interactions, risks and consequences. Reviews in Fisheries Science & Aquaculture, 26 (2):214–234, 2018. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2017.1384789.

Monkman:2018:EUS

[505] Graham George Monkman, Michel Kaiser, and Kieran Hyder. The ethics of using social media in fisheries research. Reviews in Fisheries Science & Aquaculture, 26(2):235–242, 2018. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2017.1389854.

Tilami:2018:NVF

[506] Sarvenaz Khalili Tilami and Sabine Sampels. Nutritional value of fish: Lipids, proteins, vitamins, and minerals. Reviews in Fisheries Science & Aquaculture, 26(2):243–253, 2018. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2017.1399104.

Kumar:2018:MNR

[507] Vikas Kumar, Amit Kumar Sinha, Nicholas Romano, Kristy M. Allen, Benjamin A. Bowman, Kenneth R. Thompson, and James H. Tidwell. Metabolism and nutritive role of cholesterol in the growth, gonadal development, and reproduction of crustaceans. Reviews in Fisheries Science & Aquaculture, 26(2):254–273, 2018. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2018.1429384.

Vokoun:2018:SFC

[508] Jason C. Vokoun. Stream fish community dynamics — a critical synthesis. Reviews in Fisheries Science & Aquaculture, 26(2):274, 2018. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2018.1430925.

Hunt:2018:IGF

[509] Taylor L. Hunt and Paul Jones. Informing the great fish stocking debate: an Australian case study. Reviews in Fisheries Science & Aquaculture, 26 (3):275–308, 2018. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2017.1407916.

Kara:2018:DRA

[510] Mohamed Hichem Kara, Denis Lacroix, Hélène Rey-Valette, Syndhia Mathé, and Jean Paul Blancheton. Dynamics of research in aquaculture in North Africa and support for sustainable development and innovation. Reviews in Fisheries Science & Aquaculture, 26(3):309–318, 2018. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2017.1410521.

Syed:2018:UML

[511] Shaheen Syed and Charlotte Teresa Weber. Using machine learning to uncover latent research topics in fishery models. Reviews in Fisheries Science & Aquaculture, 26(3):319–336, 2018. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2017.1416331.

delaCruz-Cervantes:2018:PAR

[512] Jesús Alberto de la Cruz-Cervantes, Flaviano Benavides-González, Jesús Genaro Sánchez-Martínez, María de la Luz Vázquez-Sauceda, and Antonio Joel Ruiz-Uribe. Propolis in aquaculture: a review of its potential. Reviews in Fisheries Science & Aquaculture, 26(3):337–349, 2018.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2018.1424798.

Pascoe:2018:EID

[513] Sean Pascoe and James Pembroke Innes. Economic impacts of the development of an offshore oil and gas industry on fishing industries: a review of experiences and assessment methods. Reviews in Fisheries Science & Aquaculture, 26(3):350–370, 2018. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2018.1436521.

Mitra:2018:OBC

[514] Anisa Mitra, Pratap Kumar Mukhopadhyay, and Sumit Homechaudhuri. An overview of biology and culture potentials of humped featherback Chitala chitala (Hamilton, 1822) — a new candidate for aquaculture diversification. Reviews in Fisheries Science & Aquaculture, 26(3):371–380, 2018.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2018.1437118.

Li:2018:GMM

[515] Erchao Li, Chang Xu, Xiaodan Wang, Shifeng Wang, Qun Zhao, Meiling Zhang, Jian G. Qin, and Liqiao Chen. Gut microbiota and its modulation for healthy farming of Pacific white shrimp Litopenaeus vannamei. Reviews in Fisheries Science & Aquaculture, 26(3):381–399, 2018. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2018.1440530.

Lindegren:2018:AFT

[516] Martin Lindegren and Keith Brander. Adapting fisheries and their management to climate change: a review of concepts, tools, frameworks, and current progress toward implementation. Reviews in Fisheries Science & Aquaculture, 26(3):400–415, 2018. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2018.1445980.

Priborsky:2018:RTC

[517] Josef Priborsky and Josef Velisek. A review of three commonly used fish anesthetics. Reviews in Fisheries Science & Aquaculture, 26(4):417–442, 2018. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2018.1442812.

Santos:2018:SAFb

[518] Sérgio Ricardo Santos and Marcelo Vianna. Scientometric analysis of fisheries science studies of Western Atlantic species of Paralichthys (Paralichthyidae: Pleuronectiformes). Reviews in Fisheries Science & Aquaculture, 26(4):443–459, 2018. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2018.1452896.

Kantoussan:2018:RFI

[519] Justin Kantoussan, Raymond Laë, and Mbaye Tine. Review of the fisheries indicators for monitoring the impacts of fishing on fish communities. Reviews in Fisheries Science & Aquaculture, 26(4):460–478, 2018. CO-

DEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2018.1458282.

Limburg:2018:MOM

[520] Karin E. Limburg, Mark J. Wuenschel, Karin Hüssy, Yvette Heimbrand, and Melvin Samson. Making the otolith magnesium chemical calendar-clock tick: Plausible mechanism and empirical evidence. Reviews in Fisheries Science & Aquaculture, 26(4):479–493, 2018. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2018.1458817.

Doria:2018:RFR

[521] Carolina R. C. Doria, Fabrice Duponchelle, Maria Alice L. Lima, Aurea Garcia, Fernando M. Carvajal-Vallejos, Claudia Coca Méndez, Michael Fabiano Catarino, Carlos Edwar de Carvalho Freitas, Blanca Vega, Guido Miranda-Chumacero, and Paul A. Van Damme. Review of fisheries resource use and status in the Madeira River Basin (Brazil, Bolivia, and Peru) before hydroelectric dam completion. Reviews in Fisheries Science & Aquaculture, 26(4):494–514, 2018. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2018.1463511.

Sugiura:2018:PAE

[522] Shozo H. Sugiura. Phosphorus, aquaculture, and the environment. Reviews in Fisheries Science & Aquaculture, 26(4):515–521, 2018. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2018.1471040.

Connor:2019:UMS

[523] William P. Connor, Kenneth F. Tiffan, James A. Chandler, Dennis W. Rondorf, Billy D. Arnsberg, and Kelvin C. Anderson. Upstream migration and spawning success of Chinook salmon in a highly developed, seasonally warm river system. Reviews in Fisheries Science & Aquaculture, 27(1):1–50, 2019. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2018.1477736.

Stoner:2019:RBF

[524] Allan W. Stoner, Martha H. Davis, and Andrew S. Kough. Relationships between fishing pressure and stock structure in Queen conch (Lobatus gigas) populations: Synthesis of long-term surveys and evidence for overfishing in the Bahamas. Reviews in Fisheries Science & Aquaculture, 27(1):51–71, 2019. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2018.1480008.

Nguyen:2019:STE

[525] Kim Anh T. Nguyen and Curtis M. Jolly. Steps toward the establishment of a commercial aquaculture insurance program: Lessons from an assessment of the Vietnamese Pilot Insurance Program. Reviews in Fisheries Science & Aquaculture, 27(1):72–87, 2019. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2018.1481363.

Cort:2019:RRS

[526] José L. Cort and Vicente D. Estruch. Review of recent studies on the absolute and relative growth of Atlantic bluefin tuna: Similarities with the Pacific bluefin tuna. Reviews in Fisheries Science & Aquaculture, 27 (1):88–105, 2019. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2018.1488817.

Nguyen:2019:ALC

[527] Khanh Q. Nguyen and Paul D. Winger. Artificial light in commercial industrialized fishing applications: a review. Reviews in Fisheries Science & Aquaculture, 27(1):106–126, 2019. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2018.1496065.

Bennion:2019:RSK

[528] Matthew Bennion, Jessica Fisher, Chris Yesson, and Juliet Brodie. Remote sensing of kelp (laminariales, ochrophyta): Monitoring tools and implications for wild harvesting. Reviews in Fisheries Science & Aquaculture, 27(2):127–141, 2019. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2018.1509056.

Chen:2019:SFR

[529] Xing Chen and Bert Scholtens. The spot-forward relationship in the Atlantic salmon market. Reviews in Fisheries Science & Aquaculture, 27 (2):142–151, 2019. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2018.1519523.

Elumalai:2019:RLF

[530] Preetham Elumalai, Abdul Salam Rubeena, Jesu Arockiaraj, Ratree Wongpanya, Matteo Cammarata, Einar Ringø, and Baskaralingam Vaseeharan. The role of lectins in finfish: a review. Reviews in Fisheries Science & Aquaculture, 27(2):152–169, 2019. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2018.1520191.

Matsumura:2019:EAS

[531] Shuichi Matsumura, Ben Beardmore, Wolfgang Haider, Ulf Dieckmann, and Robert Arlinghaus. Ecological, angler, and spatial heterogeneity drive social and ecological outcomes in an integrated landscape model of freshwater recreational fisheries. Reviews in Fisheries Science & Aquaculture, 27(2):170–197, 2019. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2018.1540549.

Farhadi:2019:MCB

[532] Ardavan Farhadi and Ayse Gül Harlioglu. Molecular and cellular biology of the crayfish spermatozoon: Toward development of artificial reproduction in aquaculture. Reviews in Fisheries Science & Aquaculture, 27 (2):198–214, 2019. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2018.1552246.

Temesgen:2019:LFC

[533] Mathewos Temesgen, Abebe Getahun, and Brook Lemma. Livelihood functions of capture fisheries in Sub-Saharan Africa: Food security, nutritional, and economic implications. Reviews in Fisheries Science & Aquaculture, 27(2):215–225, 2019. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2019.1565754.

Hamidoghli:2019:RJE

[534] Ali Hamidoghli, Jinho Bae, Seonghun Won, Seunghan Lee, Dae-Jung Kim, and Sungchul C. Bai. A review on Japanese eel (Anguilla japonica) aquaculture, with special emphasis on nutrition. Reviews in Fisheries Science & Aquaculture, 27(2):226–241, 2019. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2019.1583165.

Methratta:2019:MAF

[535] Elizabeth T. Methratta and William R. Dardick. Meta-analysis of finfish abundance at offshore wind farms. Reviews in Fisheries Science & Aquaculture, 27(2):242–260, 2019. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2019.1584601.

Hunt:2019:CNC

[536] Len M. Hunt, Ed Camp, Brett van Poorten, and Robert Arlinghaus. Catch and non-catch-related determinants of where anglers fish: a review of three decades of site choice research in recreational fisheries. Reviews in Fisheries Science & Aquaculture, 27(3):261–286, 2019. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2019.1583166.

Lewin:2019:PEI

[537] Wolf-Christian Lewin, Marc Simon Weltersbach, Keno Ferter, Kieran Hyder, Estanis Mugerza, Raúl Prellezo, Zachary Radford, Lucia Zarauz, and Harry Vincent Strehlow. Potential environmental impacts of recreational fishing on marine fish stocks and ecosystems. Reviews in Fisheries Science & Aquaculture, 27(3):287–330, 2019. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2019.1586829.

Soltani:2019:GBP

[538] Mehdi Soltani, Koushik Ghosh, Seyed Hossein Hoseinifar, Vikash Kumar, Alan J. Lymbery, Suvra Roy, and Einar Ringø. Genus Bacillus, promising probiotics in aquaculture: Aquatic animal origin, bioactive components, bioremediation and efficacy in fish and shellfish. Reviews in Fisheries Science & Aquaculture, 27(3):331–379, 2019. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2019.1597010.

Brehmer:2019:VAR

[539] Patrice Brehmer, Abdoulaye Sarré, Yvon Guennégan, and Jean Guillard. Vessel avoidance response: a complex tradeoff between fish multisensory integration and environmental variables. Reviews in Fisheries Science & Aquaculture, 27(3):380–391, 2019. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2019.1601157.

vanPoorten:2019:ACC

[540] Brett T. van Poorten and Edward V. Camp. Addressing challenges common to modern recreational fisheries with a buffet-style landscape management approach. Reviews in Fisheries Science & Aquaculture, 27(4): 393–416, 2019. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2019.1619071.

Hasan:2019:EIP

[541] Md Tawheed Hasan, Won Je Jang, Jong Min Lee, Bong-Joo Lee, Sang Woo Hur, Sang Gu Lim, Kang Woong Kim, Hyon-Sob Han, and In-Soo Kong. Effects of immunostimulants, prebiotics, probiotics, synbiotics, and potentially immunoreactive feed additives on olive flounder (Paralichthys olivaceus): a review. Reviews in Fisheries Science & Aquaculture, 27(4):417–437, 2019. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2019.1622510.

Zielinski:2019:ASL

[542] Daniel P. Zielinski, Robert McLaughlin, Theodore Castro-Santos, Bhuwani Paudel, Pete Hrodey, and Andrew Muir. Alternative sea lamprey barrier technologies: History as a control tool. Reviews in Fisheries Science & Aquaculture, 27(4):438–457, 2019. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2019.1625300.

Thomas:2019:OBR

[543] Oliver R. B. Thomas and Stephen E. Swearer. Otolith biochemistry — a review. Reviews in Fisheries Science & Aquaculture, 27(4):458–489, 2019.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2019.1627285.

Stoner:2019:FYF

[544] Allan W. Stoner. Forty years of field and laboratory experiments with hatchery-reared Queen conch: The case for conserving natural stocks. Reviews in Fisheries Science & Aquaculture, 27(4):490–516, 2019. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2019.1628705.

Anderson:2019:MAF

[545] Giulia Anderson, Elisabeth Holland, and Ciro Rico. Meta-analysis of factors influencing population differentiation in yellowfin tuna (Thunnus albacares). Reviews in Fisheries Science & Aquaculture, 27(4):517–532, 2019. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2019.1633270.

Khasanah:2020:MGE

[546] Miftakhul Khasanah, Nadiarti Nurdin, Yvonne Sadovy de Mitcheson, and Jamaluddin Jompa. Management of the grouper export trade in Indonesia. Reviews in Fisheries Science & Aquaculture, 28(1):1–15, 2020. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2018.1542420.

Doan:2020:HAP

[547] Hien Van Doan, Seyed Hossein Hoseinifar, Einar Ringø, Maria Ángeles Esteban, Maryam Dadar, Mahmoud A. O. Dawood, and Caterina Faggio. Host-associated probiotics: a key factor in sustainable aquaculture. Reviews in Fisheries Science & Aquaculture, 28(1):16–42, 2020. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2019.1643288.

Tacon:2020:TGA

[548] Albert G. J. Tacon. Trends in global aquaculture and aquafeed production: 2000–2017. Reviews in Fisheries Science & Aquaculture, 28(1):43–56, 2020. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2019.1649634.

Nowland:2020:CSP

[549] Samantha J. Nowland, Wayne A. O’Connor, Matthew W. J. Osborne, and Paul C. Southgate. Current status and potential of tropical rock oyster aquaculture. Reviews in Fisheries Science & Aquaculture, 28(1):57–70, 2020. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2019.1670134.

daSilveira:2020:MTE

[550] Estevan Luiz da Silveira, Nabil Semmar, Joan E. Cartes, Victor M. Tuset, Antoni Lombarte, Eduardo Luis Cupertino Ballester, and André Martins Vaz dos Santos. Methods for trophic ecology assessment in fishes: a critical review of stomach analyses. Reviews in Fisheries Science & Aquaculture, 28(1):71–106, 2020. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2019.1678013.

Garlock:2020:GBR

[551] Taryn Garlock, Frank Asche, James Anderson, Trond Bjørndal, Ganesh Kumar, Kai Lorenzen, Andrew Ropicki, Martin D. Smith, and Ragnar Tveterås. A global blue revolution: Aquaculture growth across regions, species, and countries. Reviews in Fisheries Science & Aquaculture, 28 (1):107–116, 2020. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2019.1678111.

Lu:2020:MFF

[552] Hongjian Lü, Mei Fu, Zhixin Zhang, Shengqi Su, and Weizhi Yao. Marking fish with fluorochrome dyes. Reviews in Fisheries Science & Aquaculture, 28(1):117–135, 2020. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2019.1681358.

Collette:2020:FBT

[553] Bruce B. Collette. The future of bluefin tunas: Ecology, fisheries management, and conservation. Reviews in Fisheries Science & Aquaculture, 28(1):136–137, 2020. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2019.1665237.

Ringo:2020:LAB

[554] Einar Ringø, Hien Van Doan, Soonho Lee, and Seong Kyu Song. Lactic acid bacteria in shellfish: Possibilities and challenges. Reviews in Fisheries Science & Aquaculture, 28(2):139–169, 2020. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2019.1683151.

Nakatsuka:2020:SSP

[555] Shuya Nakatsuka. Stock structure of Pacific bluefin tuna (Thunnus orientalis) for management purposes — a review of available information. Reviews in Fisheries Science & Aquaculture, 28(2):170–181, 2020. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2019.1686455.

Corriero:2020:RSM

[556] Aldo Corriero, Gilad Heinisch, Hanna Rosenfeld, Ivan Katavić, Letizia Passantino, Rosa Zupa, Leon Grubisić, and Molly E. Lutcavage. Review of sexual maturity in Atlantic bluefin tuna, Thunnus thynnus (Linnaeus, 1758). Reviews in Fisheries Science & Aquaculture, 28(2):182–192, 2020.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2019.1685456.

Yue:2020:CKB

[557] Gen Hua Yue, Alex Chang, A. Yuzer, and A. Suwanto. Current knowledge on the biology and aquaculture of the endangered Asian arowana. Reviews in Fisheries Science & Aquaculture, 28(2):193–210, 2020. CO-

DEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2019.1697641.

Sass:2020:CRI

[558] Greg G. Sass and Stephanie L. Shaw. Catch-and-release influences on inland recreational fisheries. Reviews in Fisheries Science & Aquaculture, 28(2):211–227, 2020. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2019.1701407.

Nguyen:2020:DCF

[559] Tram Anh Thi Nguyen and Curtis M. Jolly. Diverse cooperative field membership and small-scale producers’ access to certification. Reviews in Fisheries Science & Aquaculture, 28(2):228–246, 2020. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2019.1701984.

Ahmad:2020:NII

[560] Irshad Ahmad, Irfan Ahmad Bhat, Dhanashree D. Jagtap, C. Selvaa Kumar, S. Ferosekhan, and Luis O. B. Afonso. New insights of inhibins in ovarian physiology of fish. Reviews in Fisheries Science & Aquaculture, 28(2):247–259, 2020. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2020.1711511.

Doan:2020:MHP

[561] Hien Van Doan, Elahe Soltani, Jack Ingelbrecht, and Mehdi Soltani. Medicinal herbs and plants: Potential treatment of monogenean infections in fish. Reviews in Fisheries Science & Aquaculture, 28(2):260–282, 2020. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2020.1712325.

Salunke:2020:RSA

[562] Mahesh Salunke, Amol Kalyankar, Chandraprakash D. Khedkar, Mahesh Shingare, and Gulab D. Khedkar. A review on shrimp aquaculture in India: Historical perspective, constraints, status and future implications for impacts on aquatic ecosystem and biodiversity. Reviews in Fisheries Science & Aquaculture, 28(3):283–302, 2020. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2020.1723058.

Hoseinifar:2020:BIF

[563] Seyed Hossein Hoseinifar, Yun-Zhang Sun, Zhigzhang Zhou, Hien Van Doan, Simon J. Davies, and R. Harikrishnan. Boosting immune function and disease bio-control through environment-friendly and sustainable approaches in finfish aquaculture: Herbal therapy scenarios. Reviews in Fisheries Science & Aquaculture, 28(3):303–321, 2020. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2020.1731420.

Myers:2020:ABO

[564] Jaelen Nicole Myers, Alistair Senior, Vahid Zadmajid, Sune Riis Sørensen, and Ian Anthony Ernest Butts. Associations between ovarian fluid and sperm swimming trajectories in marine and freshwater teleosts: a metaanalysis. Reviews in Fisheries Science & Aquaculture, 28(3):322–339, 2020. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2020.1739623.

Hamidoghli:2020:NFO

[565] Ali Hamidoghli, Seonghun Won, Seunghyung Lee, Seunghan Lee, Nathaniel W. Farris, and Sungchul C. Bai. Nutrition and feeding of olive flounder Paralichthys olivaceus: a review. Reviews in Fisheries Science & Aquaculture, 28(3):340–357, 2020. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2020.1740166.

Kumar:2020:PPE

[566] Praveen Kumar I, Purabi Sarkar, Stefi Raju V, Manikandan V, Ajay Guru, Aziz Arshad, Preetham Elumalai, and Jesu Arockiaraj. Pathogenicity and pathobiology of epizootic ulcerative syndrome (EUS) causing fungus Aphanomyces invadans and its immunological response in fish. Reviews in Fisheries Science & Aquaculture, 28(3):358–375, 2020. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2020.1753167.

Molto:2020:GRB

[567] Vicenç Moltó, Pilar Hernández, Mauro Sinopoli, Amina Besbes-Benseddik, Raouf Besbes, Adriano Mariani, Miriam Gambin, Francisco Alemany, Beatriz Morales-Nin, Antoni María Grau, Juan Antonio Camiñas, José Carlos Báez, Marcelo Vasconcellos, Luca Ceriola, and Ignacio A. Catalán. A global review on the biology of the dolphinfish (Coryphaena hippurus) and its fishery in the Mediterranean Sea: Advances in the last two decades. Reviews in Fisheries Science & Aquaculture, 28(3):376–420, 2020. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2020.1757618.

Abdel-Latif:2020:BDB

[568] Hany M. R. Abdel-Latif, Mohsen Abdel-Tawwab, Mahmoud A. O. Dawood, Simon Menanteau-Ledouble, and Mansour El-Matbouli. Benefits of dietary butyric acid, sodium butyrate, and their protected forms in aquafeeds: a review. Reviews in Fisheries Science & Aquaculture, 28 (4):421–448, 2020. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2020.1758899.

Tacon:2020:FHI

[569] Albert G. J. Tacon, Daniel Lemos, and Marc Metian. Fish for health: Improved nutritional quality of cultured fish for human consumption. Reviews in Fisheries Science & Aquaculture, 28(4):449–458, 2020. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2020.1762163.

Wiernicki:2020:ESB

[570] Caroline J. Wiernicki, Dong Liang, Helen Bailey, and David H. Secor. The effect of swim bladder presence and morphology on sound frequency detection for fishes. Reviews in Fisheries Science & Aquaculture, 28(4):459–477, 2020. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2020.1762536.

Miranda:2020:RFH

[571] L. E. Miranda, Giancarlo Coppola, and Jeff Boxrucker. Reservoir fish habitats: a perspective on coping with climate change. Reviews in Fisheries Science & Aquaculture, 28(4):478–498, 2020. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2020.1767035.

Suuronen:2020:PST

[572] Petri Suuronen, C. Roland Pitcher, Robert A. McConnaughey, Michel J. Kaiser, Jan G. Hiddink, and Ray Hilborn. A path to a sustainable trawl fishery in Southeast Asia. Reviews in Fisheries Science & Aquaculture, 28(4):499–517, 2020. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2020.1767036.

Bower:2020:KGM

[573] Shannon D. Bower, Øystein Aas, Robert Arlinghaus, T. Douglas Beard, Ian G. Cowx, Andy J. Danylchuk, Kátia M. F. Freire, Warren M. Potts, Stephen G. Sutton, and Steven J. Cooke. Knowledge gaps and management priorities for recreational fisheries in the developing world. Reviews in Fisheries Science & Aquaculture, 28(4):518–535, 2020. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2020.1770689.

Diogo:2020:MRF

[574] Hugo Diogo, Pedro Veiga, Cristina Pita, Alina Sousa, David Lima, João Gil Pereira, Jorge M. S. Gonçalves, Karim Erzini, and Mafalda Rangel. Marine recreational fishing in Portugal: Current knowledge, challenges, and future perspectives. Reviews in Fisheries Science & Aquaculture, 28(4):536–560, 2020. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2020.1777083.

Sarkar:2021:SPP

[575] Uttam Kumar Sarkar, P. Mishal, Simanku Borah, Gunjan Karnatak, Ganesh Chandra, Suman Kumari, D. K. Meena, Dipesh Debnath, Sona Yengkokpam, Pronob Das, Piyashi DebRoy, Anil K. Yadav, Md. Aftabuddin, Pranab Gogoi, Arun Pandit, Biren K. Bhattacharjya, Tasso Tayung, L. Lianthuamluaia, and B. K. Das. Status, potential, prospects, and issues of floodplain wetland fisheries in India: Synthesis and review for sustainable management. Reviews in Fisheries Science & Aquaculture, 29(1):1–32, 2021. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2020.1779650.

Elumalai:2021:HIA

[576] Preetham Elumalai, Amitha Kurian, Sreeja Lakshmi, Caterina Faggio, Maria Angeles Esteban, and Einar Ringø. Herbal immunomodulators in aquaculture. Reviews in Fisheries Science & Aquaculture, 29(1):33–57, 2021. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2020.1779651.

Arlinghaus:2021:GPP

[577] Robert Arlinghaus, Øystein Aas, Josep Alós, Ivan Arismendi, Shannon Bower, Steven Carle, Tomasz Czarkowski, Kátia M. F. Freire, John Hu, Len M. Hunt, Roman Lyach, Andrzej Kapusta, Pekka Salmi, Alexander Schwab, Jun ichi Tsuboi, Marek Trella, Daryl McPhee, Warren Potts, Arkadiusz Wołos, and Zi-Jiang Yang. Global participation in and public attitudes toward recreational fishing: International perspectives and developments. Reviews in Fisheries Science & Aquaculture, 29(1):58–95, 2021. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2020.1782340.

Robinson:2021:SBE

[578] Kelly F. Robinson, Charles R. Bronte, David B. Bunnell, Peter T. Euclide, Darryl W. Hondorp, John A. Janssen, Matthew S. Kornis, Derek H. Ogle, Will Otte, Stephen C. Riley, Mark R. Vinson, Shea L. Volkel, and Brian C. Weidel. A synthesis of the biology and ecology of sculpin species in the Laurentian Great Lakes and implications for the adaptive capacity of the benthic ecosystem. Reviews in Fisheries Science & Aquaculture, 29(1):96–121, 2021. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2020.1782341.

Gephart:2021:SGA

[579] Jessica A. Gephart, Christopher D. Golden, Frank Asche, Ben Belton, Cecile Brugere, Halley E. Froehlich, Jillian P. Fry, Benjamin S. Halpern, Christina C. Hicks, Robert C. Jones, Dane H. Klinger, David C. Little, Douglas J. McCauley, Shakuntala H. Thilsted, Max Troell, and Edward H. Allison. Scenarios for global aquaculture and its role in human nutrition. Reviews in Fisheries Science & Aquaculture, 29(1):122–138, 2021. CO-

DEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2020.1782342.

Chopin:2021:ISE

[580] Thierry Chopin and Albert G. J. Tacon. Importance of seaweeds and extractive species in global aquaculture production. Reviews in Fisheries Science & Aquaculture, 29(2):139–148, 2021. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2020.1810626.

Li:2021:FLP

[581] Shan Li, Hui Wei, Lorenzo Vilizzi, Aibin Zhan, Julian D. Olden, Daniel L. Preston, Stacey A. Clarke, Becky Cudmore, Gareth D. Davies, Xiaoming Wang, and Gordon H. Copp. The future of legislation, policy, risk analysis, and management of non-native freshwater fishes in China. Reviews in Fisheries Science & Aquaculture, 29(2):149–166, 2021. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2020.1782830.

Adeleke:2021:AAC

[582] Babatunde Adeleke, Deborah Robertson-Andersson, Gan Moodley, and Simon Taylor. Aquaculture in Africa: a comparative review of Egypt, Nigeria, and Uganda vis-à-vis South Africa. Reviews in Fisheries Science & Aquaculture, 29(2):167–197, 2021. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2020.1795615.

Hoseinifar:2021:OSA

[583] Seyed Hossein Hoseinifar, Samira Yousefi, Hien Van Doan, Ghasem Ashouri, Giorgia Gioacchini, Francesca Maradonna, and Oliana Carnevali. Oxidative stress and antioxidant defense in fish: The implications of probiotic, prebiotic, and synbiotics. Reviews in Fisheries Science & Aquaculture, 29(2):198–217, 2021. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2020.1795616.

Madzivanzira:2021:RFC

[584] Takudzwa C. Madzivanzira, Josie South, Louisa E. Wood, Ana L. Nunes, and Olaf L. F. Weyl. A review of freshwater crayfish introductions in Africa. Reviews in Fisheries Science & Aquaculture, 29(2):218–241, 2021.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2020.1802405.

Bosch-Belmar:2021:JIM

[585] Mar Bosch-Belmar, Giacomo Milisenda, Lorena Basso, Thomas K. Doyle, Antonella Leone, and Stefano Piraino. Jellyfish impacts on marine aquaculture and fisheries. Reviews in Fisheries Science & Aquaculture, 29 (2):242–259, 2021. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2020.1806201.

Akhtar:2021:PBF

[586] M. S. Akhtar and A. Ciji. Pyridoxine and its biological functions in fish: Current knowledge and perspectives in aquaculture. Reviews in Fisheries Science & Aquaculture, 29(2):260–278, 2021. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2020.1813081.

Sauer:2021:WOF

[587] Warwick H. H. Sauer, Ian G. Gleadall, Nicola Downey-Breedt, Zöe Doubleday, Graham Gillespie, Manuel Haimovici, Christian M. Ibáñez, Oleg N. Katugin, Stephen Leporati, Marek R. Lipinski, Unai Markaida, Jorge E. Ramos, Rui Rosa, Roger Villanueva, Juan Arguelles, Felipe A. Briceño, Sergio A. Carrasco, Leo J. Che, Chih-Shin Chen, Rosario Cisneros, Elizabeth Conners, Augusto C. Crespi-Abril, Vladimir V. Kulik, Evgenyi N. Drobyazin, Timothy Emery, Fernando A. Fernández-Álvarez, Hidetaka Furuya, Leo W. González, Charlie Gough, P. Krishnan, Biju Kumar, Tatiana Leite, Chung-Cheng Lu, Kolliyil S. Mohamed, Jaruwat Nabhitabhata, Kyosei Noro, Jinda Petchkamnerd, Delta Putra, Steve Rocliffe, K. K. Sajikumar, Hideo Sakaguchi, Deepak Samuel, Geetha Sasikumar, Toshifumi Wada, Xiaodong Zheng, Yongjun Tian, Yumeng Pang, Anyanee Yamrungrueng, and Gretta Pecl. World octopus fisheries. Reviews in Fisheries Science & Aquaculture, 29(3):279–429, 2021. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2019.1680603.

Mendoza:2021:MAB

[588] Oscar Mendoza, Krizia Pretell, Beder Ramírez, John Sandoval, José Luis Caballero, and Gabriel Dorado. Metagenomic analyses of biofilms on whiteleg shrimp (Litopenaeus vannamei) effluents: Implications for worldwide aquaculture bioremediation and environmental sustainability in the current trend of climate change and global warming — state of the art and experimental proof of concept. Reviews in Fisheries Science & Aquaculture, 29(4):431–444, 2021. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2020.1820720.

Hussy:2021:TEP

[589] Karin Hüssy, Karin E. Limburg, Hélène de Pontual, Oliver R. B. Thomas, Philip K. Cook, Yvette Heimbrand, Martina Blass, and Anna M. Sturrock. Trace element patterns in otoliths: The role of biomineralization. Reviews in Fisheries Science & Aquaculture, 29(4):445–477, 2021. CO-

DEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2020.1760204.

Ahmadifar:2021:BDP

[590] Ehsan Ahmadifar, Morteza Yousefi, Msoumeh Karimi, Raha Fadaei Raieni, Maryam Dadar, Sevdan Yilmaz, Mahmoud A. O. Dawood, and Hany M. R. Abdel-Latif. Benefits of dietary polyphenols and polyphenolrich additives to aquatic animal health: an overview. Reviews in Fisheries Science & Aquaculture, 29(4):478–511, 2021. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2020.1818689.

Rohtla:2021:RMA

[591] Mehis Rohtla, Lorenzo Vilizzi, Vladimír Kovác, David Almeida, Bernice Brewster, J. Robert Britton, Lukasz Głowacki, Michael J. Godard, Ruth Kirk, Sarah Nienhuis, Karin H. Olsson, Jan Simonsen, Michał E. Skóra, Saulius Stakenas, Ali Serhan Tarkan, Nildeniz Top, Hugo Verreycken, Grzegorz Zieba, and Gordon H. Copp. Review and meta-analysis of the environmental biology and potential invasiveness of a poorly-studied cyprinid, the ide Leuciscus idus. Reviews in Fisheries Science & Aquaculture, 29(4):512–548, 2021. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2020.1822280.

Ahmed:2021:EAW

[592] Nesar Ahmed, Shirley Thompson, Bruce Hardy, and Giovanni M. Turchini. An ecosystem approach to wild Rice-fish cultivation. Reviews in Fisheries Science & Aquaculture, 29(4):549–565, 2021. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2020.1833833.

Yu:2021:EMI

[593] Jinkai Yu and Jing Yu. Evolution of mariculture insurance policies in China: Review, challenges, and recommendations. Reviews in Fisheries Science & Aquaculture, 29(4):566–581, 2021. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2020.1837067.

Gianasi:2021:CKB

[594] Bruno L. Gianasi, Jean-François Hamel, Emaline M. Montgomery, Jiamin Sun, and Annie Mercier. Current knowledge on the biology, ecology, and commercial exploitation of the sea cucumber Cucumaria frondosa. Reviews in Fisheries Science & Aquaculture, 29(4):582–653, 2021. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2020.1839015.

Shekhar:2021:GFS

[595] Mudagandur Shashi Shekhar, Vinaya Kumar Katneni, Ashok Kumar Jangam, Karthic Krishnan, Nimisha Kaikkolante, and Koyadan Kizhakedath Vijayan. The genomics of the farmed shrimp: Current status and application. Reviews in Fisheries Science & Aquaculture, 29(4):654–665, 2021. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2020.1858271.

Horkova-Zitnanova:2021:NAE

[596] Kristína Hôrková-Zitnanová, Kristína Svolíková, Danka Harustiaková, and Vladimír Kovác. A new approach to evaluate reproductive traits in batch-spawning fishes of indeterminate fecundity and asynchronous oocyte maturation. Reviews in Fisheries Science & Aquaculture, 29(4):666–681, 2021. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2020.1863331.

Moustahfid:2021:EFF

[597] Hassan Moustahfid, Lisa C. Hendrickson, Alexander Arkhipkin, Graham J. Pierce, Avijit Gangopadhyay, Hideaki Kidokoro, Unai Markaida, Chingiz Nigmatullin, Warwick H. Sauer, Patrizia Jereb, Gretta Pecl, Thibaut de la Chesnais, Luca Ceriola, Najih Lazar, Christopher J. Firmin, and Vladimir Laptikhovsky. Ecological-fishery forecasting of squid stock dynamics under climate variability and change: Review, challenges, and recommendations. Reviews in Fisheries Science & Aquaculture, 29(4): 682–705, 2021. CODEN ???? ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2020.1864720.

Muringai:2021:CCV

[598] Rodney. T. Muringai, Paramu. L. Mafongoya, and Romano Lottering. Climate change and variability impacts on Sub-Saharan African fisheries: a review. Reviews in Fisheries Science & Aquaculture, 29(4):706–720, 2021. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2020.1867057.

Lebel:2021:IPA

[599] Louis Lebel, Hap Navy, Tuantong Jutagate, Michael Joseph Akester, Lenore Sturm, Phimphakan Lebel, and Boripat Lebel. Innovation, practice, and adaptation to climate in the aquaculture sector. Reviews in Fisheries Science & Aquaculture, 29(4):721–738, 2021. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2020.1869695.

Nieman:2021:CSS

[600] Chelsey L. Nieman, Carolyn Iwicki, Abigail J. Lynch, Greg G. Sass, Christopher T. Solomon, Ashley Trudeau, and Brett van Poorten. Creel surveys for social-ecological-systems focused fisheries management. Reviews in Fisheries Science & Aquaculture, 29(4):739–752, 2021. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2020.1869696.

Tavakoli:2021:SCC

[601] Samad Tavakoli, Yongkang Luo, Joe M. Regenstein, Ehsan Daneshvar, Amit Bhatnagar, Yuqing Tan, and Hui Hong. Sturgeon, caviar, and caviar substitutes: From production, gastronomy, nutrition, and quality change to trade and commercial mimicry. Reviews in Fisheries Science & Aquaculture, 29(4):753–768, 2021. CODEN ???? ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2021.1873244.

Tillotson:2021:CII

[602] M. D. Tillotson, M. C. Arostegui, C. S. Austin, A. E. Lincoln, W. Matsubu, K. N. McElroy, and T. P. Quinn. Challenges in the identification and interpretation of phenological shifts: Anthropogenic influences on adult migration timing in salmonids. Reviews in Fisheries Science & Aquaculture, 29(4):769–790, 2021. CODEN ???? ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2021.1874292.

Davis:2021:UDA

[603] Robert P. Davis, Claude E. Boyd, and Donald Allen Davis. The utility of discriminant analysis to determine the geographic origin of commercially important seafood and aquaculture species: a meta-analysis. Reviews in Fisheries Science & Aquaculture, 29(4):791–799, 2021. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2021.1882935.

Perryman:2021:RAE

[604] H. A. Perryman, C. Hansen, D. Howell, and E. Olsen. A review of applications evaluating fisheries management scenarios through marine ecosystem models. Reviews in Fisheries Science & Aquaculture, 29(4):800–835, 2021. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2021.1884642.

Rehren:2022:IRP

[605] Jennifer Rehren, Melita Samoilys, Hauke Reuter, Narriman Jiddawi, and Matthias Wolff. Integrating resource perception, ecological surveys, and fisheries statistics: a review of the fisheries in Zanzibar. Reviews in Fisheries Science & Aquaculture, 30(1):1–18, 2022. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2020.1802404.

Lassen:2022:CSM

[606] Paula Graziela Lassen, Maria Eugênia Andrighetto Canozzi, Bruna Bitencourt da Costa, Júlio Otávio Jardim Barcellos, and Danilo Pedro Streit, Jr. Cryopreserved semen motility of South American neotropical fish: a meta-analysis. Reviews in Fisheries Science & Aquaculture, 30(1):19–32, 2022. CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2021.1884643.

Colt:2022:EDA

[607] John Colt, Anthonie M. Schuur, Dallas Weaver, and Kenneth Semmens. Engineering design of aquaponics systems. Reviews in Fisheries Science & Aquaculture, 30(1):33–80, 2022. CODEN ???? ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2021.1886240.

Yeo:2022:MSB

[608] JuDong Yeo and Christopher C. Parrish. Mass spectrometry-based lipidomics in the characterization of individual triacylglycerol (TAG) and phospholipid (PL) species from marine sources and their beneficial health effects. Reviews in Fisheries Science & Aquaculture, 30(1):81–100, 2022.

CODEN ???? ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2021.1897968.

Falcao:2022:ZTS

[609] M. A. P. Falcão, M. C. dos Santos Dantas, C. T. Rios, L. P. Borges, M. R. Serafini, A. G. Guimarães, and C. I. B. Walker. Zebrafish as a tool for studying inflammation: a systematic review. Reviews in Fisheries Science & Aquaculture, 30(1):101–122, 2022. CODEN ????

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2021.1898538.

Sara:2022:SIA

[610] G. Sarà, M. C. Mangano, M. Berlino, L. Corbari, M. Lucchese, G. Milisenda, S. Terzo, M. S. Azaza, J. M. F. Babarro, R. Bakiu, B. R. Broitman, A. H. Buschmann, R. Christofoletti, A. Deidun, Y. Dong, J. Galdies, B. Glamuzina, O. Luthman, P. Makridis, A. J. A. Nogueira, M. G. Palomo, R. Dineshram, G. Rilov, P. Sanchez-Jerez, H. Sevgili, M. Troell, K. Y. AbouelFadl, M. N. Azra, P. Britz, C. Brugere, E. Carrington, I. Celić, F. Choi, C. Qin, T. Dobroslavić, P. Galli, D. Giannetto, J. Grabowski, M. J. H. Lebata-Ramos, P. T. Lim, Y. Liu, S. M. Llorens, G. Maricchiolo, S. Mirto, M. Pećarević, N. Ragg, E. Ravagnan, D. Saidi, K. Schultz, M. Shaltout, C. Solidoro, S. H. Tan, V. Thiyagarajan, and B. Helmuth. The synergistic impacts of anthropogenic stressors and COVID-19 on aquaculture: a current global perspective. Reviews in Fisheries Science & Aquaculture, 30(1):123–135, 2022. CODEN

???? ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2021.1876633.

Tacon:2022:FFSa

[611] Albert G. J. Tacon, Marc Metian, and Aaron A. McNevin. Future feeds: Suggested guidelines for sustainable development. Reviews in Fisheries Science & Aquaculture, 30(2):135–142, 2022. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/

10.1080/23308249.2020.1860474. See notice of duplicate article publication [671].

Campanati:2022:SIA

[612] Camilla Campanati, David Willer, Jasmin Schubert, and David C. Aldridge. Sustainable intensification of aquaculture through nutrient recycling and circular economies: More fish, less waste, blue growth. Reviews in Fisheries Science & Aquaculture, 30(2):143–169, 2022. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2021.1897520.

Huang:2022:MFB

[613] Bin Kang and Xiaoxia Huang. Mekong fishes: Biogeography, migration, resources, threats, and conservation. Reviews in Fisheries Science & Aquaculture, 30(2):170–194, 2022. CODEN RFSCFJ. ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2021.1906843.

Secor:2022:ASS

[614] D. H. Secor, M. H. P. O’Brien, N. Coleman, A. Horne, I. Park, D. C. Kazyak, D. G. Bruce, and C. Stence. Atlantic sturgeon status and movement ecology in an extremely small spawning habitat: The Nanticoke River–Marshyhope Creek, Chesapeake Bay. Reviews in Fisheries Science & Aquaculture, 30(2):195–214, 2022. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2021.1924617.

Dadswell:2022:DIC

[615] Michael Dadswell, Aaron Spares, Jeffrey Reader, Montana McLean, Tom McDermott, Kurt Samways, and Jessie Lilly. The decline and impending collapse of the Atlantic salmon (Salmo salar) population in the North Atlantic Ocean: a review of possible causes. Reviews in Fisheries Science & Aquaculture, 30(2):215–258, 2022. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2021.1937044.

Love:2022:ORS

[616] David C. Love, Frank Asche, Ruth Young, Elizabeth M. Nussbaumer, James L. Anderson, Robert Botta, Zach Conrad, Halley E. Froehlich, Taryn M. Garlock, Jessica A. Gephart, Andrew Ropicki, Joshua S. Stoll, and Andrew L. Thorne-Lyman. An overview of retail sales of seafood in the USA, 2017–2019. Reviews in Fisheries Science & Aquaculture, 30(2): 259–270, 2022. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2021.1946481.

Tacon:2022:FFSb

[617] Albert G. J. Tacon, Marc Metian, and Aaron A. McNevin. Future feeds: Suggested guidelines for sustainable development. Reviews in Fisheries Science & Aquaculture, 30(2):271–279, 2022. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/

10.1080/23308249.2021.1898539. See notice of duplicate article publication [671].

Shackell:2022:SEA

[618] Nancy L. Shackell, Jonathan A. D. Fisher, Cornelia E. den Heyer, Daniel R. Hennen, Andrew C. Seitz, Arnault Le Bris, Dominique Robert, Michael E. Kersula, Steven X. Cadrin, Richard S. McBride, Christopher H. McGuire, Tony Kess, Krista T. Ransier, Chang Liu, Andrew Czich, and Kenneth T. Frank. Spatial ecology of Atlantic halibut across the Northwest Atlantic: a recovering species in an era of climate change. Reviews in Fisheries Science & Aquaculture, 30(3):281–305, 2022. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2021.1948502.

Naslund:2022:PF

[619] Joacim Näslund and Laith A. Jawad. Pugheadedness in fishes. Reviews in Fisheries Science & Aquaculture, 30(3):306–329, 2022. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2021.1957772.

Avigliano:2022:OMD

[620] Esteban Avigliano. Optimizing the methodological design in fish stock delineation from otolith chemistry: Review of spatio-temporal analysis scales. Reviews in Fisheries Science & Aquaculture, 30(3):330–345, 2022. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2021.1961679.

Stoner:2022:SRR

[621] Allan W. Stoner and Richard S. Appeldoorn. Synthesis of research on the reproductive biology of queen conch (Aliger gigas): Toward the goals of sustainable fisheries and species conservation. Reviews in Fisheries Science & Aquaculture, 30(3):346–390, 2022. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2021.1968789.

Adkison:2022:RSS

[622] Milo Douglas Adkison. A review of salmon spawner-recruitment analysis: The central role of the data and its impact on management strategy. Reviews in Fisheries Science & Aquaculture, 30(3):391–427, 2022. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2021.1972086.

Costa-Pierce:2022:FSP

[623] Barry Antonio Costa-Pierce, Abigail B. Bockus, Bela H. Buck, Sander W. K. van den Burg, Thierry Chopin, Joao G. Ferreira, Nils Goseberg, Kevin G. Heasman, Johan Johansen, Sandra E. Shumway, Neil A. Sims, and Albert G. J. Tacon. A fishy story promoting a false dichotomy to policy-makers: It is not freshwater vs. marine aquaculture. Reviews in Fisheries Science & Aquaculture, 30(4):429–446, 2022. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2021.2014175.

Zhao:2022:AEV

[624] Na Zhao, Qiuxia Deng, Chunhua Zhu, and Bo Zhang. Application of extracellular vesicles in aquatic animals: a review of the latest decade. Reviews in Fisheries Science & Aquaculture, 30(4):447–466, 2022. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2021.1985429.

Ozretich:2022:PAR

[625] Reed W. Ozretich, Chelsea L. Wood, Fiona Allan, Ahou Rachel Koumi, Rachel Norman, Andrew S. Brierley, Giulio A. De Leo, and David C. Little. The potential for aquaculture to reduce poverty and control schistosomiasis in Côte d’Ivoire (Ivory Coast) during an era of climate change: a systematic review. Reviews in Fisheries Science & Aquaculture, 30(4): 467–497, 2022. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2022.2039096.

Frischer:2022:BGM

[626] Marc E. Frischer, Stephen C. Landers, Anna N. Walker, Shirley A. Powell, and Richard F. Lee. Black gill in marine decapod crustaceans: a review. Reviews in Fisheries Science & Aquaculture, 30(4):498–519, 2022. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2022.2047153.

Volvenko:2022:AZE

[627] Igor V. Volvenko. The adaptive zone: From evolutionary biology to ecology and biogeography. review and a case study with examples for fisheries science. Reviews in Fisheries Science & Aquaculture, 30(4):520–541, 2022. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2022.2067741.

Madsen:2022:OHA

[628] Henry Madsen, Hung M. Nguyen, Guy R. Lanza, and Jay R. Stauffer Jr. A one health approach relative to trematode-caused diseases of people and animals associated with aquaculture. Reviews in Fisheries Science & Aquaculture, 30(4):542–566, 2022. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2022.2090830.

Zajicek:2023:RMA

[629] Paul Zajicek, John Corbin, Sebastian Belle, and Robert Rheault. Refuting marine aquaculture myths, unfounded criticisms, and assumptions. Reviews in Fisheries Science & Aquaculture, 31(1):1–28, 2023. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2021.1980767.

OKeefe:2023:ETA

[630] Catherine E. O’Keefe, Steven X. Cadrin, Gildas Glemarec, and Yann Rouxel. Efficacy of time-area fishing restrictions and gear-switching as solutions for reducing seabird bycatch in gillnet fisheries. Reviews in Fisheries Science & Aquaculture, 31(1):29–46, 2023. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2021.1988051.

Dumilag:2023:DES

[631] Richard V. Dumilag, Bea A. Crisostomo, Zae-Zae A. Aguinaldo, Lourie Ann R. Hinaloc, Lawrence M. Liao, Hilly Ann Roa-Quiaoit, Floredel Dangan-Galon, Giuseppe C. Zuccarello, Marie-Laure Guillemin, Juliet Brodie, Elizabeth J. Cottier-Cook, and Michael Y. Roleda. The diversity of eucheumatoid seaweed cultivars in the Philippines. Reviews in Fisheries Science & Aquaculture, 31(1):47–65, 2023. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2022.2060038.

Mrnak:2023:APT

[632] Joseph T. Mrnak, Logan W. Sikora, M. Jake Vander Zanden, and Greg G. Sass. Applying panarchy theory to aquatic invasive species management: a case study on invasive rainbow smelt Osmerus mordax. Reviews in Fisheries Science & Aquaculture, 31(1):66–85, 2023. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2022.2078951.

Rubino:2023:PCM

[633] Michael C. Rubino. Policy considerations for marine aquaculture in the United States. Reviews in Fisheries Science & Aquaculture, 31(1):86– 102, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2022.2083452.

Yadav:2023:CRR

[634] Anamika Yadav, Md Tabish Noori, Abhijit Biswas, and Booki Min. A concise review on the recent developments in the Internet of Things (IoT)-based smart aquaculture practices. Reviews in Fisheries Science & Aquaculture, 31(1):103–118, 2023. CODEN RFSCFJ. ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2022.2090228.

Glencross:2023:DDM

[635] Brett Glencross, Ian Carr, and Ester Santigosa. Distribution, deposition, and modelling of lipid and long-chain polyunsaturated fatty acids in Atlantic salmon fillets. Reviews in Fisheries Science & Aquaculture, 31(1): 119–140, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2022.2090831.

Hansen:2023:TTM

[636] Benni Winding Hansen. Two tropical marine copepods demonstrate physiological properties needed for mass production. Reviews in Fisheries Science & Aquaculture, 31(1):141–159, 2023. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2022.2095198.

Delaney:2023:STA

[637] Alyne Delaney, David G. Reid, Christopher Zimmermann, Marloes Kraan, Nathalie A. Steins, and Michel J. Kaiser. Socio-technical approaches are needed for innovation in fisheries. Reviews in Fisheries Science & Aquaculture, 31(2):161–179, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2022.2047886.

Humphries:2023:AAG

[638] Fran Humphries, Charles Lawson, John A. H. Benzie, and Clare Morrison. African aquaculture: Genetic resource and traditional knowledge access and benefit sharing measures. Reviews in Fisheries Science & Aquaculture, 31(2):180–196, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2022.2027866.

Sarr:2023:DSM

[639] Ousmane Sarr, Richard Kindong, Siquan Tian, Fambaye Ngom Sow, Momar Ka, and Ousmane Niang Seye. Diagnosis of the Senegalese marine fisheries profile during the last two decades: a perspective toward fisheries management. Reviews in Fisheries Science & Aquaculture, 31(2): 197–215, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2022.2057184.

Yu:2023:PDO

[640] Jinkai Yu and Jiatong Liu. Policies in the development of offshore cage aquaculture in China: Evolution, performance, and prospects. Reviews in Fisheries Science & Aquaculture, 31(2):216–232, 2023. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2022.2103644.

Mubango:2023:IBA

[641] Elliot Mubango, Samad Tavakoli, Yueyue Liu, Yanyan Zheng, Xinrou Huang, Soottawat Benjakul, Tan Yuqing, Yongkang Luo, and Hui Hong. Intermuscular bones in Asian carps: Health threats, solutions, and future directions. Reviews in Fisheries Science & Aquaculture, 31(2):233– 258, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2022.2117980.

Love:2023:IOA

[642] David C. Love, Frank Asche, Jessica A. Gephart, Jiafeng Zhu, Taryn Garlock, Joshua S. Stoll, James Anderson, Zach Conrad, Elizabeth M. Nussbaumer, Andrew L. Thorne-Lyman, and Martin W. Bloem. Identifying opportunities for aligning production and consumption in the U.S. fisheries by considering seasonality. Reviews in Fisheries Science & Aquaculture, 31(2):259–273, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2022.2121601.

Tacon:2023:CFS

[643] Albert G. J. Tacon. Contribution of fish and seafood to global food and feed supply: an analysis of the FAO Food Balance Sheet for 2019. Reviews in Fisheries Science & Aquaculture, 31(2):274–283, 2023. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2022.2124364.

Troell:2023:FOS

[644] M. Troell, P. J. G. Henriksson, A. H. Buschmann, T. Chopin, and S. Quahe. Farming the ocean — seaweeds as a quick fix for the climate? Reviews in Fisheries Science & Aquaculture, 31(3):285–295, 2023. CO-

DEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2022.2048792.

Delargy:2023:GRC

[645] Adam J. Delargy, Lynda Blackadder, Isobel Bloor, Carrie McMinn, David B. Rudders, Claire L. Szostek, Helen Dobby, Mervi Kangas, Bryce D. Stewart, James R. Williams, and Kevin D. E. Stokesbury. A global review of catch efficiencies of towed fishing gears targeting scallops. Reviews in Fisheries Science & Aquaculture, 31(3):296–319, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2022.2139170.

Kamaruzzaman:2023:OAE

[646] Yeny Nadira Kamaruzzaman and Muzzneena Ahmad Mustapha. An overview assessment of the effectiveness of satellite images and remote sensing in predicting potential fishing grounds and its applicability for Rastrelliger kanagurta in the Malaysian EEZ off the South China Sea. Reviews in Fisheries Science & Aquaculture, 31(3):320–341, 2023. CO-

DEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2023.2183341.

Wirgin:2023:EHR

[647] Isaac Wirgin, R. Christopher Chambers, John R. Waldman, Nirmal K. Roy, David A. Witting, and Mark T. Mattson. Effects of Hudson River stressors on Atlantic tomcod: Contaminants and a warming environment. Reviews in Fisheries Science & Aquaculture, 31(3):342–371, 2023. CO-

DEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2023.2189483.

Haase:2023:ABS

[648] Kevin Haase, Oliver Reinhardt, Wolf-Christian Lewin, Marc Simon Weltersbach, Harry V. Strehlow, and Adelinde M. Uhrmacher. Agentbased simulation models in fisheries science. Reviews in Fisheries Science & Aquaculture, 31(3):372–395, 2023. CODEN RFSCFJ. ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2023.2201635.

Scarnecchia:2023:ECP

[649] Dennis L. Scarnecchia. The extinction of the Chinese paddlefish Psephurus gladius: Transnationalism, technology transfer, and timescape. Reviews in Fisheries Science & Aquaculture, 31(3):396–419, 2023. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2023.2201636.

Villasante:2023:CSF

[650] Alejandro Villasante, Carolina Ramírez, Elías Figueroa Villalobos, Wellison A. Pereira, Madison S. Powell, Delbert M. Gatlin III, Patricio Dantagnan, and Jaime Romero. Creatine in sustainable fish aquaculture. Reviews in Fisheries Science & Aquaculture, 31(3):420–451, 2023. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2023.2210689.

Tomasso:2023:BRA

[651] Joe Tomasso. Book review: Aquaculture: An Introductory Text — 4th Edition, Robert R. Stickney and Delbert M. Gatlin III. Reviews in Fisheries Science & Aquaculture, 31(3):452, 2023. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2022.2160570.

Peng:2023:ICM

[652] Daomin Peng, Yongtong Mu, Yugui Zhu, Jiansong Chu, and U. Rashid Sumaila. Insights from Chinese mariculture development to support global blue growth. Reviews in Fisheries Science & Aquaculture, 31(4):453– 457, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2023.2167515.

Sun:2023:SMM

[653] Ming Sun, Yunzhou Li, Lisa Suatoni, Alexander Kempf, Marc Taylor, Elizabeth Fulton, Cody Szuwalski, Maria Teresa Spedicato, and Yong Chen. Status and management of mixed fisheries: a global synthesis. Reviews in Fisheries Science & Aquaculture, 31(4):458–482, 2023. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2023.2213769.

Yu:2023:AID

[654] Jinkai Yu and Tingjun Yan. Analyzing industrialization of deep-sea cage mariculture in China: Review and performance. Reviews in Fisheries Science & Aquaculture, 31(4):483–496, 2023. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2023.2214623.

Anonymous:2023:AER

[655] Kim Birnie-Gauvin, David A. Patterson, Steven J. Cooke, Scott G. Hinch, and Erika J. Eliason. Anaerobic exercise and recovery: Roles and implications for mortality in Pacific salmon. Reviews in Fisheries Science & Aquaculture, 31(4):497–522, 2023. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2023.2224902.

Anonymous:2023:GRD

[656] Taryn Garlock, Frank Asche, James L. Anderson, Jana Hilsenroth, Kai Lorenzen, Ruth Beatriz Mezzalira Pincinato, and Ragnar Tveterås. Global and regional determinants of diversity in blue foods. Reviews in Fisheries Science & Aquaculture, 31(4):523–534, 2023. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2023.2225627.

Anonymous:2023:RBQ

[657] Allan W. Stoner, Megan Davis, and Andrew S. Kough. Relationships between queen conch larval biology and recruitment, connectivity, and fishery management. Reviews in Fisheries Science & Aquaculture, 31(4): 535–597, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2023.2228905.

VanLeeuwen:2023:CWT

[658] Travis E. Van Leeuwen, Sarah J. Lehnert, Cindy Breau, Michelle Fitzsimmons, Nicholas I. Kelly, J. Brian Dempson, Victoria M. Neville, Mark Young, Donald Keefe, Tomas J. Bird, and David Cote. Considerations for water temperature-related fishery closures in recreational Atlantic salmon (Salmo salar) catch and release fisheries: a case study from Eastern Canada. Reviews in Fisheries Science & Aquaculture, 31(4):598–619, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2023.2242959.

Karatayev:2023:NEZ

[659] Alexander Y. Karatayev, Daniel P. Molloy, and Lyubov E. Burlakova. Natural enemies of zebra and quagga mussels: Predators, parasites, and ecological competitors. Reviews in Fisheries Science & Aquaculture, 32 (1):1–80, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2023.2235437.

Long:2023:GSN

[660] James M. Long and Lauren Seguy. Global status of non-native largemouth bass (Micropterus salmoides, Centrachidae) and smallmouth bass (Micropterus dolomieu, Centrarchidae): Disparate views as beloved sportfish and feared invader. Reviews in Fisheries Science & Aquaculture, 32 (1):81–98, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2023.2244078.

Encinas-Garcia:2023:RSC

[661] T. Encinas-García, F. Mendoza-Cano, A. Muhlia-Almazán, Marco Porchas-Cornejo, and A. Sánchez-Paz. A review of shrimp cellular receptors for WSSV: Potential targets for antiviral strategies in shrimp aquaculture. Reviews in Fisheries Science & Aquaculture, 32(1):99–126, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2023.2254401.

Diggles:2023:RSA

[662] Benjamin K. Diggles, Robert Arlinghaus, Howard I. Browman, Steven J. Cooke, Robin L. Cooper, Ian G. Cowx, Charles D. Derby, Stuart W. Derbyshire, Paul Jb Hart, Brian Jones, Alexander O. Kasumyan, Brian Key, Julian G. Pepperell, D. Christopher Rogers, James D. Rose, Alex Schwab, Anne B. Skiftesvik, Don Stevens, Jeffrey D. Shields, and Craig Watson. Reasons to be skeptical about sentience and pain in fishes and aquatic invertebrates. Reviews in Fisheries Science & Aquaculture, 32(1): 127–150, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2023.2257802.

Gray:2023:FET

[663] Tim Gray. Fishing Europe’s troubled waters: 50 years of fisheries policy,. Reviews in Fisheries Science & Aquaculture, 32(1):151, 2023. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2023.2225325.

Peng:2023:SGB

[664] Daomin Peng, Yugui Zhu, Sandra E. Shumway, Jiansong Chu, and U. Rashid Sumaila. Supporting global blue economy through sustainable molluscan mariculture with a focus on China. Reviews in Fisheries Science & Aquaculture, 32(1):152–170, 2023. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2023.2260489.

Persson:2023:PBS

[665] Agneta Persson, Barry C. Smith, Jennifer H. Alix, and Gary H. Wikfors. Properties and behavior of sexual life stages underlying dinoflagellate HAB events of cyst-producing species that disrupt fisheries and aquaculture. Reviews in Fisheries Science & Aquaculture, 32(2):171–188, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2023.2268715.

Love:2023:SRA

[666] David C. Love, Lisa M. Weltzien, Andrew L. Thorne-Lyman, Natalie S. Armstrong, Elizabeth Chatpar, Marisa Koontz, Dahiany Zayas-Toro, Lionel Dabbadie, Jonathan Lansley, Felix Marttin, Stefania Vannuccini, Elizabeth M. Nussbaumer, and Florence Poulain. A scoping review of aquatic food systems during the COVID-19 pandemic. Reviews in Fisheries Science & Aquaculture, 32(2):189–210, 2023. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2023.2231096.

Tacon:2023:ABS

[667] Albert G. J. Tacon, Rafael T. I. Coelho, Jéssica Levy, Thaís M. Machado, Cristiane R. P. Neiva, and Daniel Lemos. Annotated bibliography of selected papers dealing with the health benefits and risks of fish and seafood consumption. Reviews in Fisheries Science & Aquaculture, 32 (2):211–305, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2023.2238821.

Boyd:2023:RUP

[668] Claude E. Boyd and Aaron A. McNevin. Resource use and pollution potential in feed-based aquaculture. Reviews in Fisheries Science & Aquaculture, 32(2):306–333, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2023.2258226.

FernandezCunha:2023:MCF

[669] Marta Fernandez Cunha, Ezequiel R. Coscueta, María Emilia Brassesco, Frederico Almada, David Gonçalves, and Manuela Pintado. Methods for the collection of fish mucus: A systematic review. Reviews in Fisheries Science & Aquaculture, 32(2):334–368, 2023. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2023.2289012.

Oliver:2024:LHA

[670] Luke P. Oliver, Christopher A. Myrick, Gary Fornshell, and Kenneth D. Cain. Life history and aquaculture methods for burbot (Lota lota). Reviews in Fisheries Science & Aquaculture, 32(2):369–386, 2024. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2024.2308917.

Anonymous:2023:C

[671] Anonymous. Correction. Reviews in Fisheries Science & Aquaculture, 32(2):387, 2023. CODEN RFSCFJ. ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2023.2165299.

See [611, 617].

Tacon:2024:CNI

[672] Albert G. J. Tacon and Sandra E. Shumway. Critical need to increase aquatic food production and food supply from aquaculture and capture fisheries: Trends and outlook. Reviews in Fisheries Science & Aquaculture, 32(3):389–395, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2024.2324321.

Glencross:2024:SAU

[673] Brett Glencross, Xiaowen Ling, Delbert Gatlin, Sachi Kaushik, Margareth Øverland, Richard Newton, and Luisa M. P. Valente. A SWOT analysis of the use of marine, grain, terrestrial-animal and novel protein ingredients in aquaculture feeds. Reviews in Fisheries Science & Aquaculture, 32(3): 396–434, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2024.2315049.

Quinn:2024:EIB

[674] Thomas P. Quinn, John Epifanio, Peter B. Moyle, Stan Gregory, Daniel J. Hasselman, Erik Merrill, Kenneth Rose, Thomas F. Turner, and Thomas C. Wainwright. Ecological interactions between non-native American shad and Pacific salmon: The Columbia River case study. Reviews in Fisheries Science & Aquaculture, 32(3):435–449, 2024. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2024.2337434.

Midway:2024:SFM

[675] Stephen R. Midway, Frederick S. Scharf, Michael A. Dance, Nancy J. Brown-Peterson, Joseph C. Ballenger, Nicolette S. Beeken, Russell J. Borski, Tanya L. Darden, Kenneth A. Erickson, Troy M. Farmer, Ashley Fincannon, John Godwin, Patrick M. Graham, Jennifer L. Green, Henry Hershey, Dylan Kiene, Laura M. Lee, Michael S. Loeffler, Anne Markwith, Caitlin McGarigal, Sean P. Powers, David R. Smith, Aaron M. Watson, and Shelby B. White. Southern flounder: Major milestones and remaining knowledge gaps in their biology, ecology, and fishery management. Reviews in Fisheries Science & Aquaculture, 32(3):450–478, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.

com/doi/full/10.1080/23308249.2024.2341017.

Dempson:2024:RFP

[676] J. Brian Dempson, Travis E. Van Leeuwen, Ian R. Bradbury, Sarah J. Lehnert, David Coté, Frédéric Cyr, Christina Pretty, and Nicholas I. Kelly. A review of factors potentially contributing to the long-term decline of Atlantic salmon in the Conne River, Newfoundland, Canada. Reviews in Fisheries Science & Aquaculture, 32(3):479–504, 2024. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2024.2341023.

Mahe:2024:APP

[677] Ruiji Mahe, Zhangying Ye, Ming Qi, Wenlong Cai, Yanci Wen, Jinxing Xiao, Ying Liu, Songming Zhu, and Jian Zhao. Application progress and prospect of game theory on aquaculture: From the perspectives of breeding, rearing, and marketing. Reviews in Fisheries Science & Aquaculture, 32(3):505–525, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2024.2347627.

Rose:2023:RFF

[678] James D. Rose. Review of Fish Feel Pain — Scrutiny of a Dogma. Reviews in Fisheries Science & Aquaculture, 32(3):526–527, 2023. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2023.2299648.

Weston:2024:NRH

[679] Julia Cirne Lima Weston and Ingrid Kelling. Navigating responsibility for human rights compliance in the fishing industry. Reviews in Fisheries Science & Aquaculture, 32(4):529–544, 2024. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2024.2329915.

Glencross:2024:ESM

[680] Brett D. Glencross, Enrico Bachis, David Robb, and Richard Newton. The evolution of sustainability metrics for the marine ingredient sector: Moving towards holistic assessments of aquaculture feed. Reviews in Fisheries Science & Aquaculture, 32(4):545–561, 2024. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2024.2337426.

Colt:2024:DSP

[681] John Colt, Joseph Tetreault, and Rachel L. Fogle. Development and standardization of physical, operational, and performance metrics for aquaponics. Reviews in Fisheries Science & Aquaculture, 32(4):562– 578, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2024.2353578.

Farhadi:2024:GIP

[682] Ahmad Farhadi, Arya Vazirzadeh, Andrew G. Jeffs, and Shane D. Lavery. Genetic insights into the population connectivity, biogeography, and management of fisheries-important spiny lobsters (Palinuridae). Reviews in Fisheries Science & Aquaculture, 32(4):579–611, 2024. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2024.2355576.

Yochum:2024:GEA

[683] Noëlle Yochum, Junita D. Karlsen, Jesse F. Senko, John H. Wang, Valentina Melli, Michele Luca Geraci, Anne Christine Utne-Palm, Michael Breen, Elsa Cuende, Shannon M. Bayse, Jasmine Somerville, Andreas Hermann, Alexius Edridge, Emma K. Mackenzie, Peter Ljungberg, Jérôme Chladek, Mattias Van Opstal, Dan Watson, Mark J. M. Lomeli, and Martin Oliver. Guidelines for evaluating artificial light to mitigate unwanted fisheries bycatch. Reviews in Fisheries Science & Aquaculture, 32(4):612–656, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2024.2359417.

Simmons:2024:OKM

[684] Olivia Meredith Simmons, David Aldvén, Torbjørn Forseth, Stephanie Müller, Olle Calles, Patrik Andreasson, and Ana T. Silva. An overview of kelt migration in regulated rivers: Status, knowledge gaps, and future directions toward safe downstream passage at hydropower facilities. Reviews in Fisheries Science & Aquaculture, 32(4):657–678, 2024. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2024.2362221.

Owatari:2024:LTL

[685] Marco Shizuo Owatari, Vinícius Ronzani Cerqueira, Manecas Francisco Baloi, Filipe dos Santos Cipriano, Gilberto Caetano Manzoni, Fabio Carneiro Sterzelecki, and Caio Magnotti. Low-trophic-level species Sardinella brasiliensis in aquaculture: Crosstalk between aquaculture and fisheries. Reviews in Fisheries Science & Aquaculture, 32(4):679– 697, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2024.2362236.

DAbramo:2024:RPA

[686] Louis R. D’Abramo. Realizing the potential of aquaculture: Undertaking the wicked problems of climate change, fed production systems and global food security. Reviews in Fisheries Science & Aquaculture, 33 (1):1–7, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2024.2389537.

Soengas:2024:FIR

[687] José L. Soengas, Sara Comesaña, Ayelén M. Blanco, and Marta Conde-Sieira. Feed intake regulation in fish: Implications for aquaculture. Reviews in Fisheries Science & Aquaculture, 33(1):8–60, 2024. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2024.2374259.

Zielinski:2024:ARF

[688] D. P. Zielinski, J. Kerr, K. M. Bærum, O. M. Simmons, A. T. Silva, and R. A. Goodwin. Advancements in riverine fish movement modeling: Bridging environmental complexity and fish behavior. Reviews in Fisheries Science & Aquaculture, 33(1):61–76, 2024. CODEN RF-

SCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/
10.1080/23308249.2024.2374964.

Mechaly:2024:CEM

[689] Alejandro S. Mechaly, Cynthia Awruch, Elsa Cabrita, Benjamín Costas, Jorge M. O. Fernandes, Victor Gallego, Jorge Hirt-Chabbert, Ioannis Konstantinidis, Catarina Olivera, Sandra Ramos-Júdez, Lourenço Ramos-Pinto, and Elvira Fatsini. Cutting-edge methods in teleost and chondrichthyan reproductive biology. Reviews in Fisheries Science & Aquaculture, 33(1):77–112, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http:

//www.tandfonline.com/doi/full/10.1080/23308249.2024.2377999.

Wirgin:2024:PSB

[690] Isaac Wirgin, Lorraine Maceda, Joseph Stabile, Mike Bednarski, and John Waldman. Population structure and broad-scale movements of Atlantic sturgeon along the North American Atlantic Coast inferred from genetic analysis. Reviews in Fisheries Science & Aquaculture, 33(1): 113–137, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2024.2383470.

Chacon-Guzman:2024:ASF

[691] Jonathan Chacón-Guzmán, Ricardo Jiménez-Montealegre, Wanshu Hong, Enric Gisbert, Sandra Ramos-Júdez, Juan Carlos Pérez-Urbiola, and Neil Duncan. Aquaculture of the Sciaenidae family: Main species cultured worldwide and emerging species in Latin America, offering new opportunities for aquaculture diversification. Reviews in Fisheries Science & Aquaculture, 33(1):138–163, 2024. CODEN RFSCFJ. ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2024.2383849.

Hargreaves:2024:PSA

[692] John A. Hargreaves. Book review: Principles of Sustainable Aquaculture, 2nd edition. Reviews in Fisheries Science & Aquaculture, 33(1): 164–165, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2024.2428904.

Glencross:2024:OFA

[693] Brett D. Glencross, Enrico Bachis, Monica B. Betancor, Philip Calder, Nina Liland, Richard Newton, and Bente Ruyter. Omega-3 futures in aquaculture: Exploring the supply and demands for long-chain omega-3 essential fatty acids by aquaculture species. Reviews in Fisheries Science & Aquaculture, 33(2):167–216, 2024. CODEN RFSCFJ. ISSN

2330-8249. URL http://www.tandfonline.com/doi/full/10.1080/
23308249.2024.2388563.

Caballero-Huertas:2024:SCF

[694] Marta Caballero-Huertas, Marcela Salazar-Moscoso, and Laia Ribas. Sex is a crucial factor in the immune response: an ichthyological perspective. Reviews in Fisheries Science & Aquaculture, 33(2):217–237, 2024. CO-

DEN RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/
doi/full/10.1080/23308249.2024.2390965.

Fricke:2024:ESP

[695] A. Fricke, E. Capuzzo, R. Bermejo, L. C. Hofmann, I. Hernández, R. Pereira, S. W. K. Van den Burg, T. Pereira, A. H. Buschmann, and E. J.

Cottier-Cook. Ecosystem services provided by seaweed cultivation: State of the art, knowledge gaps, constraints and future needs for achieving maximum potential in Europe. Reviews in Fisheries Science & Aquaculture, 33(2):238–256, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://

www.tandfonline.com/doi/full/10.1080/23308249.2024.2399355.

Cowx:2024:UTF

[696] Ian G. Cowx, An V. Vu, Zeb Hogan, Martin Mallen-Cooper, Lee J. Baumgartner, T. Quan Lai, Gunther Grill, and Catherine A. Sayer. Understanding the threats to fish migration: Applying the global swimways concept to the Lower Mekong. Reviews in Fisheries Science & Aquaculture, 33(2): 257–285, 2024. CODEN RFSCFJ. ISSN 2330-8249. URL http://www.

tandfonline.com/doi/full/10.1080/23308249.2024.2401018.

Gebauer:2024:CSO

[697] Tatyana Gebauer, Radek Gebauer, Vlastimil Stejskal, Marieke Verleih, Ewumi Azeez Folorunso, Tomás Korytár, and Alexander Rebl. Current status and opportunities of immunological research on Percid fishes. Reviews in Fisheries Science & Aquaculture, 33(2):286–318, 2024. CODEN

RFSCFJ. ISSN 2330-8249. URL http://www.tandfonline.com/doi/
full/10.1080/23308249.2024.2414225.

Tacon:2024:FHR

[698] Albert G. J. Tacon, Jéssica Levy, Rafael Coelho, Thais M. Machado, Cristiane R. P. Neiva, and Daniel E. L. Lemos. Fish for health: Role of fish in global food and animal protein supply. Reviews in Fisheries Science & Aquaculture, 33(2):319–333, 2024. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2024.2420917.

Kuhn:2024:MLA

[699] Bernhard Kühn, Arjay Cayetano, Jennifer I. Fincham, Hassan Moustahfid, Maria Sokolova, Neda Trifonova, Jordan T. Watson, Jose A. Fernandes-Salvador, and Laura Uusitalo. Machine learning applications for fisheries-at scales from genomics to ecosystems. Reviews in Fisheries Science & Aquaculture, 33(2):334–357, 2024. CODEN RFSCFJ.

ISSN 2330-8249. URL http://www.tandfonline.com/doi/full/10.
1080/23308249.2024.2423189.